### v1

# New Tight Bounds for SGD without Variance Assumption: A Computer-Aided Lyapunov Analysis

Abstract
The analysis of Stochastic Gradient Descent (SGD) often relies on making some assumption on the variance of the stochastic gradients, which is usually not satisfied or difficult to verify in practice. This paper contributes to a recent line of works which attempt to provide guarantees without making any variance assumption, leveraging only the (strong) convexity and smoothness of the loss functions. In this context, we prove new theoretical bounds derived from the monotonicity of a simple Lyapunov energy, improving the current state-of-the-art and extending their validity to larger step-sizes. Our theoretical analysis is backed by a Performance Estimation Problem analysis, which allows us to claim that, empirically, the bias term in our bounds is tight within our framework.
1 Introduction
Given $n\geq 1$ convex and smooth real-valued functions $f_{1},\dots,f_{n}\colon\mathbb{R}^{d}\to\mathbb{R}$, we consider the problem given by
$\min_{x\in\mathbb{R}^{d}}f(x),\quad\text{where }f(x)=\frac{1}{n}\sum_{i=1}^{n}f_{i}(x).$ (Finite-Sum)
We denote $S=\operatorname{argmin}(f)$, and assume it to be nonempty.
The Stochastic Gradient Descent (SGD) algorithm (robbins_stochastic_1951) is one of the most popular approaches for solving Problem (Finite-Sum), and is widely used for large-scale machine learning and stochastic optimization problems. In its standard version, at each iteration, SGD performs a gradient step using one of the functions in the sum, which is chosen at random. More precisely, given the iterate $x_{t}$, SGD picks $i_{t}\in\{1,\dots,n\}$ uniformly at random, and computes $x_{t+1}$ via
$x_{t+1}=x_{t}-\gamma\nabla f_{i_{t}}(x_{t}),$ (SGD)
where $\gamma>0$ is the step-size. Intuitively, SGD is able to progress towards the solutions of the problem because it follows a direction that, on average, is an unbiased estimator for $\nabla f$. The simplicity and efficiency of SGD make it the go-to algorithm for training deep neural networks and other models on massive datasets, where computing the entire gradient $\nabla f$ is intractable or too expensive.
Complexity of SGD. Quantitative results on the behavior of SGD consist in upper bounds for a metric $\Delta(T)$, which is to be interpreted as an optimality gap at iteration $T$. Typical choices include the difference between the expected value of $f$ at a point obtained after $T$ iterations and its minimum value, or the expected square-distance from such a point to a solution. In most known results, these upper bounds are expressed as the sum of two terms, taking the form
$\Delta(T)\leq{\rm Bias}(T)+{\rm Variance}(T),$
where the bias term typically vanishes when $T\to+\infty$, and where the variance term can ideally be made arbitrarily small by taking a small enough step-size. Ideally, we would like the bias term to be equal to the rate of convergence of Gradient Descent (GD).
Assumptions for SGD. Classical studies on SGD typically assume that each function $f_{i}$ is $L$-smooth and either convex, strongly convex or satisfies a Polyak-Lojasiewicz inequality. Some works assume relaxations of these hypotheses, making assumptions under the form of variational inequalities, such as the expected smoothness (gower_sgd_2019), the $(L_{0},L_{1})$-smoothness (zhang_why_2020), the expected residual (gower_stochastic_2021), or the ABC condition (khaled_better_2023). Other works assume that only the function $f$ is smooth or convex. For the sake of readability, we will assume that each of the functions $f_{i}$ is convex and smooth, even though we show in the Appendix A that we simply need an assumption analog to expected smoothness.
Besides the assumptions on the problem itself, it is quite standard to make an assumption on the algorithm, through assumptions on the variance of the stochastic gradients $\nabla f_{i_{t}}(x_{t})$. Many different variance assumptions have been studied in the context of SGD. The oldest and possibly the best-known of such assumptions is the uniform boundedness of the variance (robbins_stochastic_1951; nemirovski_robust_2009; rakhlin_making_2012; schmidt_fast_2013), or a variant, namely:
$\sup_{x\in\mathbb{R}^{d}}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x)-\nabla f(x)\|^{2}}\right]\mathclose{}<+\infty\quad\text{or}\quad\sup_{x\in\mathbb{R}^{d}}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x)\|^{2}}\right]\mathclose{}<+\infty.$
The latter assumption strictly implies the former, and is usually made in non-smooth stochastic optimization (bubeck_convex_2015, Section 6), or in online optimization (hazan_introduction_2016), whereas the former is more common in our convex and smooth setting. These assumptions are however unrealistic in practice (bottou_optimization_2018; nguyen_sgd_2018), and are never verified for strongly convex functions. The bounded variance assumption has been relaxed into weaker assumptions, such as the Blum-Gladyshev assumption (blum_approximation_1954; gladyshev_stochastic_1965; alacaoglu_towards_2025), maximal strong growth (schmidt_fast_2013), strong and weak growth (vaswani_fast_2019), or relaxed growth (bertsekas_gradient_2000; bottou_optimization_2018), to name a few. But the problem remains that those assumptions cannot be verified a priori in practice for smooth problems.
In this paper, we make no boundedness assumptions on the variance, a setting already pioneered in (bach_non-asymptotic_2011; needell_stochastic_2016; gower_sgd_2019).
Our contribution. We establish new and improved quantitative results on the complexity of SGD in the convex and $L$-smooth setting, without making any variance assumption. Instead our bounds will rely on the so-called solution variance constant, which is always finite and defined as
$\sigma_{*}^{2}=\frac{1}{m}\sum_{i=1}^{n}\|\nabla f_{i}(x_{*})\|^{2},$
where $x_{*}$ is any minimizer of $f$. Indeed, the constant does not depend on the choice of $x_{*}$, see (garrigos_handbook_2024, Lemma 4.17). Our results take the form
$\Delta(T)\leq{\rm Bias}(T)\cdot\|x_{0}-x_{*}\|^{2}+{\rm Variance}(T)\cdot\sigma_{*}^{2},$
and will be valid for the whole range of step-sizes $\gamma L\in(0,2)$. We focus on obtaining the smallest bias term possible, and on top of improving it with respect to the existing results, we conducted numerical experiments illustrating the sharpness of our bound. To the best of our knowledge this is the first work where a unified and comprehensive study of SGD, without bounded variance assumptions and in the whole range of step-sizes, has been carried out in the convex and smooth setting.
In the convex case, we establish bounds for the expected function value gap $\Delta(T)=\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}$, where $\bar{x}_{T}$ is the Cesàro average of the iterates generated by SGD up to iteration $T$. Our bounds, with simplified notation for ease of presentation, are:
${\rm Bias}(T)=\begin{cases}\tfrac{1}{2\gamma T}&\text{ if }\gamma L\in(0,1),\\ \tfrac{1}{(2-\varepsilon)\gamma T}&\text{ if }\gamma L=1,\varepsilon>0,\\ \tfrac{1}{2\gamma(2-\gamma L)T}&\text{ if }\gamma L\in(1,2),\end{cases}\quad{\rm Variance}(T)=\begin{cases}\tfrac{\gamma}{2(1-\gamma L)}&\text{ if }\gamma L\in(0,1),\\ \tfrac{\gamma(2+\varepsilon)}{\varepsilon(2-\varepsilon)}&\text{ if }\gamma L=1,\varepsilon>0,\\ \tfrac{\exp(T)}{2-\gamma L}&\text{ if }\gamma L\in(1,2).\end{cases}$
In our setting where no assumptions on the variance are made, the only existing bounds are available for step-sizes limited to $\gamma L\in(0,\tfrac{1}{2})$, and provide worse constants (gower_sgd_2021). Our results are the best so far and the first one valid in the whole interval $(0,1)$. Moreover, it matches the bias term previously obtained under the bounded variance assumption (taylor_stochastic_2019). For $\gamma L=1$ we are able to have a bias term as close as possible from the optimal bias $\tfrac{1}{2\gamma L}$, but cannot have equality with a finite variance term. Numerical results suggest that it is not possible to achieve this optimal bias. For $\gamma L\in(1,2)$, which is a completely uncharted territory for this problem, we find that the variance term grows exponentially with $T$, making it impractical. Again, our numerical experiments suggest that one cannot avoid this growth in the variance term. Note that those problems encountered for large step-sizes $\gamma L\in[1,2)$ are typically avoided when making a uniformly bounded variance assumption, see e.g. taylor_stochastic_2019.
As a by-product, our new bounds for SGD with the step-size $\gamma L=1$ makes it possible to study the stochastic proximal algorithm through the lens of SGD. To the best of our knowledge, we establish the first complexity guarantees for this method in a general convex nonsmooth setting.
In the case where each function $f_{i}$ is $\mu$-strongly convex, we provide bounds over the expected distance to the solution, namely $\Delta(T)=\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}$. Our results show that we can take ${\rm Bias}(T)=\phi^{2T}$ where $\phi$ corresponds to the standard rate of GD:
$\phi=\max\{1-\gamma\mu;\gamma L-1\},$
with the exception of the optimal step-size $\gamma=\tfrac{2}{\mu+L}$, for which the bias term must be relaxed to avoid a blow up in the variance. These results improve on (bach_non-asymptotic_2011; needell_stochastic_2016; gower_sgd_2019) by allowing for the full range of step-sizes $\gamma L\in(0,2)$, and by having a sharp bias term. Our numerical experiments suggest again that we cannot improve those results.
The strategy. For each fixed time horizon $T$, we propose a time-dependent energy random sequence
$E_{t}=a_{t}\cdot\|x_{t}-x_{*}\|^{2}+\rho\cdot\sum_{s=0}^{t-1}\big{(}f(x_{s})-\min f\big{)}-\sum_{s=0}^{t-1}e_{s}\sigma_{*}^{2}\quad\forall t=0,\ldots,T,$ (Lyapunov)
where $\rho$, $(a_{t})$ and $(e_{t})$ are nonnegative. The first ingredient of our Lyapunov energy is the distance to the solution $\|x_{t}-x_{*}\|^{2}$, a classical term which typically decreases for deterministic gradient dynamics. The second term involves the function gap $f(x_{t})-\min f$, which also typically decreases for gradient descent. The standard Lyapunov for gradient descent usually contains a term $t(f(x_{t})-\min f)$, which we have replaced here with the sum of the past function gaps, which is of the same order in time. This choice is critical to obtain bounds without variance assumptions (compare with the Lyapunov energies considered by taylor_stochastic_2019). The last term is a negative cumulated sum, where the $e_{t}$’s play the role of a variance term. It is here to compensate the fluctuations caused by the variance in the SGD algorithm, and will allow the Lyapunov energy to decrease.
A decrease of $\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ along the iterations means that $\mathbb{E}\mathopen{}\left[{E_{T}}\right]\mathclose{}\leq E_{0}$, which translates into a bound for SGD. Our strategy is therefore to find admissible Lyapunov parameters $\rho$, $(a_{t})$ and $(e_{t})$, which are those that make the energy decrease in expectation. Among those admissible parameters, our goal is to find the ones minimizing, whenever possible, the bias term in our bound.
To this end, we rely on the approach followed in taylor_stochastic_2019 to design Lyapunov potentials in a systematic fashion. It is based on the Performance Estimation Problem (PEP) methodology, which originally appeared in drori_performance_2014, and was further developed in taylor_smooth_2017. A similar line of thought was followed in (fercoq_defining_2024). Automated Lyapunov analyses have also been developed in (taylor_lyapunov_2018; upadhyaya_automated_2025) by combining the PEP approach with the Integral Quadratic Constraints framework (lessard_analysis_2016). The PEP framework allowed us to numerically drive our theoretical proofs, and to assess the sharpness of our bounds.
This paper is organized as follows: Section 2 presents our main results in the convex setting, while Section 3 focuses on the strongly convex setting. We discuss the tightness of our results in Section 4, supported by numerical experiments. All technical results and extensions of our results are gathered in the appendices.
2 Convex Setting
In this section, we consider the smooth convex setting, which we formalize below. Note that the assumption could be relaxed, see the discussion in Section 5.1.
Assumption 2.1 (Convexity and smoothness).
Considering Problem (Finite-Sum), we assume that each function $f_{i}$ is convex and $L$-smooth for some $L\in(0,+\infty)$.
As mentioned in the introduction, our analysis is based on Lyapunov energies of the form (Lyapunov). The next lemma illustrates how a Lyapunov decrease leads to upper bounds for (SGD).
Lemma 2.2 (Bound from Lyapunov decrease).
Let Assumption 2.1 hold. Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for every $t=0,\dots,T-1$, that $\rho>0$, and, without loss of generality, that $a_{0}=1$. Then
$\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]\leq\frac{\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\bar{e}\sigma_{*}^{2}}{\rho},$
where $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$ and $\bar{e}=\tfrac{1}{T}\sum_{t=0}^{T-1}e_{t}$.
The next results are consequences of Lemma 2.2, after finding specific values of $\rho$, $(a_{k})$ and $(e_{k})$, aiming at the largest possible value for $\rho$. Most results will be presented with a uniform (and simpler) upper bound on $e_{k}$ for ease of presentation, and the more detailed constants can be found in Appendix C, whereto all proofs are postponed. The presentation of our results is split into three parts based on the normalized step-size $\gamma L$, namely Theorem 2.3 for short step-sizes $\gamma L\in(0,1)$, Theorem 2.4 for the optimal step-size $\gamma L=1$ and Theorem 2.7 for large step-sizes $\gamma L\in(1,2)$.
Theorem 2.3 (Convex case, short step-sizes).
Let Assumption 2.1 hold, and let $(x_{t})$ be generated by (SGD) with $\gamma L\in(0,1)$. Then, for every $T\geq 1$, with $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}\leq\frac{L\|x_{0}-x_{*}\|^{2}}{2\gamma LT+2(1-\gamma L)}+\frac{\gamma\sigma_{*}^{2}}{2(1-\gamma L)}\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\frac{\gamma\sigma_{*}^{2}}{2(1-\gamma L)}.$
The setting $\gamma L\in(0,1)$, usually considered short step-sizes in the GD literature, is the only one for which we have competitors in the literature. In (gower_sgd_2021, Theorem D.6) the authors obtain bounds of the same order as ours, albeit worse, for step-sizes limited to $\gamma L\in\left(0,\frac{1}{2}\right)$. They also obtain results by relaxing the convexity and smoothness assumptions, see more details in Appendix C.4. The authors in (garrigos_handbook_2024, Theorem 5.5) focus on the same assumptions as ours, and also obtain similar but worse bounds, for $\gamma L\in(0,\frac{1}{4}]$. We are not aware of any works for SGD without variance assumption providing bounds for the full range $\gamma L\in(0,1)$.
To the best of our knowledge, making a uniform variance assumption does not allow to obtain a bound with a better bias term than the one we obtained in Theorem 2.3. For instance (taylor_stochastic_2019, Theorem 6) obtain the same bias term as ours for short step-sizes, and liu_revisiting_2024 achieved a bias of $\frac{1}{\gamma L}$, although their results focuses on the last iterate.
Note that our result focuses on the Cesaro averages of the iterates $\bar{x}_{T}$, which is standard for SGD with no variance assumption. This is in contrast with the literature with uniformly bounded variance, for which it is possible to derive bounds on the last iterate $x_{T}$. The first result of this kind in the convex and smooth setting was established by bach_non-asymptotic_2011, with bounds recently improved in (taylor_stochastic_2019, Theorem 5) and (liu_revisiting_2024, Theorem 3.1). They obtain a variance term that scales with $T$ and $\log(T)$, respectively. Both were able to derive complexity bounds, with the latter offering a significantly better one over the former. It remains an open question to know whether or not it is possible to obtain bounds on the last iterate, similar to the one established in Theorem 2.3, without uniformly bounded variance. For the time being, this showcases a discrepancy between the uniformly bounded variance setting and ours.
We now turn to the case $\gamma L=1$, which we call the optimal step-size, with a slight abuse of terminology, in analogy with the optimal step-size for GD, and in view of the numerical results presented in Section 4.
Theorem 2.4 (Convex case, optimal step-size).
Let Assumption 2.1 hold, and let $(x_{t})$ be generated by (SGD) with $\gamma L=1$. Then, for every $T\geq 1$ and $\varepsilon\in(0,2)$, with $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]\leq\frac{\|x_{0}-x_{*}\|^{2}}{(2-\varepsilon)\gamma T}+\frac{\gamma(2+\varepsilon)\sigma_{*}^{2}}{\varepsilon(2-\varepsilon)}.$
The case $\gamma L=1$, which is standard for GD, has surprisingly not yet seen any theoretical guarantees without bounded variance assumptions. As far as we know, this is the first upper bound provided in this setting. We point out the fact that Theorem 2.4 cannot be seen as the limit of Theorem 2.3 when $\gamma L\to 1$, because it would require to set $\varepsilon=0$, which results in an infinite variance term. Although it would seem like a flaw of our analysis that $\varepsilon=0$ is not feasible, we present evidence against this in Section 4.1. This singularity came as a surprise to us, and suggests that having a bias term of the order of $\tfrac{L}{2T}$ is impossible for SGD without any variance assumptions. This is a striking difference with GD, but also with SGD with bounded variance assumption, where no singularity holds at $\gamma L=1$, see for instance (taylor_stochastic_2019, Theorem 6).
Remark 2.5 (Stochastic proximal algorithm).
While the convergence rates for the deterministic proximal algorithm are well-studied, complexity results for its stochastic counterpart are unknown in a general setting. But a classic trick allows to rewrite the stochastic proximal algorithm as a particular instance of SGD, for a problem with no bounded variance, and with a step-size which must be $\gamma L=1$. Thanks to Theorem 2.4 we are now able to handle this optimal step-size and to derive the first complexity bound for this algorithm with no other assumption on the functions than convexity, avoiding, for instance, to impose finiteness and Lipschitzness as in davis_stochastic_2019. For more details, see Appendix H.
Interestingly, when $\sigma_{*}^{2}=0$ the variance term vanishes and allows us to take $\varepsilon\to 0$. This setting is known as the interpolation regime, because it is equivalent to ask that all the functions $f_{i}$ share a common minimizer, see (garrigos_handbook_2024, Lemma 4.17). It is a feature typically shared by overparametrized models.
Corollary 2.6 (Convex case with interpolation, optimal step-size).
Let Assumption 2.1 hold and assume that $\sigma_{*}^{2}=0$. Let $(x_{t})$ be generated by (SGD) with $\gamma L=1$. Then, for every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T},$
where $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$.
The case $\gamma L\in(1,2)$, referred to as large step-sizes, is yet unexplored for SGD without bounded variance assumptions. We present below the first result of this type.
Theorem 2.7 (Convex case, large step-sizes).
Let Assumption 2.1 hold, and let $(x_{t})$ be generated by (SGD) with $\gamma L\in(1,2)$. Then, for every $T\geq 1$, with $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}\leq\frac{\delta\|x_{0}-x_{*}\|^{2}}{2\gamma(2-\gamma L)T}+\frac{\gamma\bar{e}_{T}^{\prime}\sigma_{*}^{2}}{2(2-\gamma L)^{3}},$
where $\delta=1-(1-\gamma L)^{2T}\in(0,1)$, and $\bar{e}_{T}^{\prime}$ is exponential in $T$ (see its definition in Appendix C).
In this result, we obtain a bound where the variance term increases exponentially with $T$. This means that, for large step-sizes, the variance term will quickly dominate the bias term, preventing the algorithm from making significant progress. Of course, one might argue that this comment is not relevant because it is based on an upper bound, and that the proof of this upper bound could be improved. However, we provide empirical evidence in Appendix F which supports the claim that our bound is tight, and that the exponential increase in the variance term cannot be avoided.
This highlights a second discrepancy between the uniformly bounded variance setting and ours. In the large step-size regime, we are unable to provide a uniform bound of the variance term with respect to the time horizon $T$. This contrasts with the result in (taylor_stochastic_2019, Theorem 6), which, under the assumption of uniformly bounded variance, establishes a variance term that remains bounded with respect to $T$, for step-sizes $\gamma L\in(0,\tfrac{1+\sqrt{5}}{2})$.
3 Strongly Convex Setting
We now focus on the strongly convex setting. As in the convex setting, the below assumption may be relaxed, as we discuss further in Section 5.1.
Assumption 3.1 (Strong convexity and smoothness).
Considering Problem (Finite-Sum), we assume that each function $f_{i}$ is $\mu$-strongly convex and $L$-smooth for some $\mu,L\in(0,+\infty)$.
Our analysis is still based on the Lyapunov energy introduced in (Lyapunov). The following lemma explains how a decrease in the energy yields an upper bound for the iterates of (SGD).
Lemma 3.2 (Bound from Lyapunov decrease).
Let Assumption 3.1 hold. Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for all $t=0,\ldots,T-1$, $\rho\geq 0$, and that, without loss of generality, $a_{T}=1$. Then
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq{a_{0}\|x_{0}-x_{*}\|^{2}}+{e^{\text{sum}}\sigma_{*}^{2}},\quad\text{ with }\quad e^{\text{sum}}=\sum_{t=0}^{T-1}e_{t}.$
We now present our main results for the strongly convex case, which are obtained after carefully choosing Lyapunov parameters in Appendix D. The first theorem provides a bound with sharp bias terms for all step-sizes, except for the so-called optimal step-size $\gamma=\frac{2}{L+\mu}$ for which the variance explodes. The second theorem provides a bound valid for every step-size (including the optimal one), with an $\varepsilon$-sub-optimal bias term which allows for a finite variance term.
Theorem 3.3 (Strongly convex case, sharp bias).
Let Assumption 3.1 hold, and let $(x_{t})$ be generated by (SGD) with $\gamma L\in(0,2)$ such that $\gamma\neq\gamma_{opt}=\tfrac{2}{\mu+L}$. Then, for every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\phi^{2T}\cdot\|x_{0}-x_{*}\|^{2}+\frac{1-\phi^{2T}}{1-\phi^{2}}e\sigma_{*}^{2},$
where $\phi=\max\{1-\gamma\mu;\gamma L-1\}\in[0,1)$, and $e=\gamma^{2}\left(1+\tfrac{\gamma(L-\mu)}{|\gamma-\gamma_{opt}|(L+\mu)}\right)$.
Theorem 3.4 (Strongly convex case, sub-optimal bias).
Let Assumption 3.1 hold, and let $(x_{t})$ be generated by (SGD) with $\gamma L\in(0,2)$. Then, for every $T\geq 1$ and $\varepsilon>0$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\phi^{2T}\cdot\|x_{0}-x_{*}\|^{2}+\frac{1-\phi^{2T}}{1-\phi^{2}}e\sigma_{*}^{2},$
where $\phi^{2}=\varepsilon+(\max\{1-\gamma\mu;\gamma L-1\})^{2}\in[0,1)$, and $e=\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)^{2}}{4\varepsilon}\right)$.
As in the convex case, when interpolation holds, namely when $\sigma_{*}^{2}=0$, the division by $\varepsilon$ is no longer a problem, allowing us to take $\varepsilon\to 0$.
Corollary 3.5 (Strongly convex case with interpolation, optimal bias).
Let Assumption 3.1 hold, and assume that $\sigma_{*}^{2}=0$. Let $(x_{t})$ be generated by (SGD) with $\gamma L\in(0,2)$. Then, for every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\phi^{2T}\cdot\|x_{0}-x_{*}\|^{2},$
where $\phi=\max\{1-\gamma\mu;\gamma L-1\}\in[0,1)$.
The first message to take from our results is that it is possible to have bounds for SGD whose bias term corresponds exactly to the standard rate for GD, namely a geometrical rate governed by $\phi^{2}$. A second insight is that it is possible to obtain bounds for the full range of step-sizes $\gamma L\in(0,2)$. This is in contrast with (bach_non-asymptotic_2011, Theorem 1), which employs a similar Lyapunov approach to ours, and obtains a bias term governed by $\phi^{2}+\gamma^{2}(L^{2}-\mu^{2})$ for step-sizes $\gamma L\in(0,\tfrac{\mu}{L})$. We can also compare with (needell_stochastic_2016, Theorem 2.1), which provides an improved bias term governed by $\phi^{2}+2\gamma^{2}\mu(L-\mu)$ for short step-sizes $\gamma L\in(0,1)$, or the more recent (gower_sgd_2019, Theorem 3.1), whose bias term is governed by $\phi$ (with no square) for step-sizes $\gamma L\in(0,\tfrac{1}{2})$.
It is remarkable that being able to work with large step-sizes allows us to observe a singular phenomenon for the optimal step-size, which, to our knowledge, was never observed before. In the deterministic setting, the step-size $\gamma=\tfrac{2}{\mu+L}$ is the only one leading to the optimal geometric rates governed by $\phi^{2}=\tfrac{(L-\mu)^{2}}{(L+\mu)^{2}}$. In our stochastic setting, it appears that achieving such a bias term is incompatible with having a finite variance. It seems reasonable to think that achieving the optimal rate for a deterministic method requires such precision that introducing a small perturbation or variance might break it. Such conjecture seems to be validated by our numerical experiments presented in the next section.
4 Sharpness of our bounds
In this section, we delve into discussing the tightness of our bounds. To this end, we employ a numerical approach based on the Performance Estimation Problem (PEP) methodology, initially introduced by drori_performance_2014, further improved by taylor_smooth_2017, and adapted by taylor_stochastic_2019 in the context of stochastic algorithms, that we will follow closely. We direct the reader to Appendix E for an in-depth explanation of our PEP framework. Roughly speaking, the idea consists in reformulating the problem of finding admissible Lyapunov parameters ensuring the energy decrease into a semi-definite program, which may be solved numerically. For this paper, we focused on finding admissible coefficients yielding minimal bias terms in our bounds.
This framework inspired and drove our proofs, helping us to postulate the best Lyapunov parameters that could be found. We made this clear in our proofs, which can be found in the Appendices C and D. Although most of our theorems in Sections 2 and 3 are stated using simplified values, for an easier presentation, in this section we will make use of the exact values found in our proofs. Moreover, it allowed us to verify the tightness of our bounds, at least numerically. We recognize that theoretical results overlapping with numerical observations do not constitute a proof of optimality. Nevertheless, these validations allow us to formulate conjectures that our results are sharp, some of which we partially prove. This tightness assessment will be the focus of the rest of this section.
All of our numerical experiments [^1] Code provided in supplementary material. We will provide a link to a public repository in the final version. were developed in Python 3.13 using the solvers MOSEK version 11.0.19 (mosek_mosek_2025) and Clarabel version 0.9.0 (goulart_clarabel_2024) to solve the resulting semi-definite programs, implemented using CVXPY version 1.6.5 (diamond_cvxpy_2016), and run on Intel Xeon Platinum 8380 CPUs.
4.1 Convex Setting
The primary focus of our work has been to identify sharp bias terms, which directly translates to maximizing the value of $\rho$ via Lemma 2.2, once we assume, without loss of generality, that $a_{0}=1$. We denote by $\rho_{\text{opt}}$ the supremum of all such possible admissible constants $\rho$. In parallel, we define $\rho_{\text{theory}}$ to be our guess for $\rho_{\text{opt}}$ driven by our results in Theorems 2.3, 2.4 and 2.7:
$\rho_{\text{theory}}\coloneqq\begin{cases}2\gamma+\tfrac{2}{TL}(1-\gamma L)&\text{ if }\gamma L\in(0,1),\\ 2\gamma&\text{ if }\gamma L=1,\\ 2\gamma(2-\gamma L)\left[1-(1-\gamma L)^{2T}\right]^{-1}&\text{ if }\gamma L\in(1,2).\end{cases}$
Figure 1 compares our analytically derived $\rho_{\text{theory}}$ with the value of $\rho_{\text{opt}}$ obtained numerically. The perfect match across the full range of step-sizes validates the sharpness of the bias term in our theoretical bounds for step-sizes $\gamma L\neq 1$. This empirical sharpness is validated by a formal proof, for large step-sizes (see Proposition C.10 for more details).
Proposition 4.1.
For large step-sizes $\gamma L\in(1,2)$, there exists no Lyapunov parameters with $a_{0}=1$ and $\rho>\rho_{\text{theory}}$ making the energy decrease.
For the optimal step-size $\gamma L=1$, we empirically see that we cannot do better than $2\gamma$, but we were not able to achieve it in Theorem 2.4. To understand why, we turn to the study of the variance term.
Figure 1: Theoretical and numerical bias term ($L=1$).
As a second set of experiments, we minimized the constant $\overline{e}$ (which governs the variance term in our bound from Lemma 2.2) under the constraint that $\rho=\rho_{\text{theory}}$. We will denote the infimum of such constants as $\overline{e}_{\text{opt}}$, and will note $\overline{e}_{\text{theory}}$ the constant we obtained in our results. We compare the values of those two constants in Figure 2, for non-optimal step-sizes. We can see that, as $\gamma L\to 1$, either from the left or from the right, the variance term tends to infinity, suggesting that it is not possible to maintain $\rho=\rho_{\text{opt}}$ when $\gamma L=1$. Moreover, the solver fails when trying to compute $\overline{e}_{\text{opt}}$ for this optimal step-size, which again points to the fact that $\rho=2\gamma$ is not a feasible Lyapunov parameter when $\gamma L=1$. This is why in Theorem 2.4 we proposed a bias term with any $\rho_{\varepsilon}<\rho_{\text{opt}}$, in view of maintaining a finite variance. In consequence we conjecture that, within such Lyapunov framework, it is not possible to obtain a bound for SGD with a bias term of the order $\tfrac{1}{2\gamma T}$ when $\gamma L=1$.
A bi-product of the results displayed in Figure 2 is that the variance terms we obtained in Theorems 2.3 and 2.7 overlap with the variance returned by the solver. This suggests that we cannot avoid the issue we had with Theorem 2.7, where the variance term grows exponentially with time.
We note that the problem, specially for large step-sizes, is highly numerically unstable, which is why we do not provide plots for $T>2$. This is further discussed in Appendix F.
Figure 2: Theoretical and numerical average variance term ($L=1$).
4.2 Strongly Convex Setting
In this setting, our focus has also been to identify sharp bias terms which, through Lemma 3.2, amounts to minimizing the value of $a_{0}$, once we assume, without loss of generality, that $a_{T}=1$. We denote by $a_{0}^{\text{opt}}$ the infimum of all such possible admissible constants. Driven by our result in Theorem 3.3, we expect $a_{0}^{\text{opt}}$ to be equal to $\phi^{2T}$, with $\phi=\max\{1-\gamma\mu,\gamma L-1\}$. Figure 3 compares $\phi^{2T}$ with the value of $a_{0}^{\text{opt}}$ obtained numerically, and illustrates again the tightness of our results. Note that the iteration count is smaller here than in the convex setting, due to numerical instability issues, which we detail in Appendix F.
Figure 3: Theoretical and numerical bias term ($L=1$, $\mu=0.25$ and $\tfrac{2}{\mu+L}=1.6$).
For the optimal step-size $\gamma=\frac{2}{L+\mu}$, we empirically see that we cannot do better than $\phi^{2T}$, but we were not able to achieve it in Theorem 3.3. Here again we will understand why by looking at the variance terms. To do so, we implement the minimization of the constant $e^{sum}$ appearing in Lemma 3.2 under the constraint that $a_{0}$ is equal to $\phi^{2T}$. Figure 4 compares the optimal value $e^{sum}_{\text{opt}}$ to the constant we obtained in our proofs. We can see that, as the step-size gets closer to the optimal one, the variance tends to +$\infty$+ ∞. This strongly suggests that it is not possible to obtain a bias term of the order $\left(\tfrac{L-\mu}{L+\mu}\right)^{2T}$ by means of a Lyapunov analysis such as ours.
Figure 4: Theoretical and numerical total variance term ($L=1$, $\mu=0.25$ and $\tfrac{2}{\mu+L}=1.6$).
5 Conclusion and perspectives
In this paper we proved new and improved upper bounds for SGD without any variance assumption, both in the convex and strongly convex setting, for a full range of step-sizes. We moreover illustrated the sharpness of these results numerically using the Performance Estimation Problem methodology. Doing so, we raised questions regarding unexpected singularities appearing for optimal step-sizes.
5.1 Extensions
For the sake of presentation, we formulated our main results under strong but easy to grasp assumptions, and left the more general and technical assumptions to Appendix A. For instance, all of our results remain valid when replacing the Problem (Finite-Sum) with the minimization of an expectation $\mathbb{E}\mathopen{}\left[{f_{i}}\right]\mathclose{}$ of a (possibly infinite) family of functions $f_{i}$. Moreover, Assumptions 2.1 and 3.1 may be replaced by a weaker symmetric expected cocoercivity assumption.
We moreover note that the analysis carried out for (SGD) extends to the case of mini-batching and to non-uniform sampling without much additional effort. This is explained in Appendix G.
5.2 Future Work
We leave the following questions open and lines of research:
  • Are the upper bounds obtained improvable through more general forms of Lyapunov energies? We worked with an energy of the form (Lyapunov), but do not claim this is the only sensible form.
  • Is it possible to obtain upper bounds on (SGD) whose bias term match the best upper bound in the deterministic gradient descent algorithm? Both in the convex and strongly convex setting, we still observe a gap, and we conjecture this is not due to non-tightness of our analysis but rather a flaw in the core analysis technique.
  • The analysis carried out solely focuses on constant step-sizes. Can this type of analysis be extended to cover varying step-sizes?
  • In this paper, we focused on achieving sharp bias terms. However, it is not the only possible approach, and we intend in a future work to investigate bounds providing sharp complexity rates.
Acknowledgments and Disclosure of Funding
This work benefited from the support of the FMJH Program Gaspard Monge for optimization and operations research and their interactions with data science. We also thank the Center for Information Technology of the University of Groningen for their support and for providing access to the Hábrók high performance computing cluster.
Appendix A Problem definition and main assumptions
Throughout the appendix sections, we consider a more general setting than in Sections 2 and 3. Instead of limiting ourselves to the minimization of a finite sum of functions in $\mathbb{R}^{d}$, we consider the minimization of an expectation of functions defined on a real Hilbert space.
Problem A.1.
[…]
  1. 1.
    the problem is well-defined, in the sense that $i\mapsto f_{i}(x)$ is $\mathcal{D}$-measurable, and that $\mathbb{E}\mathopen{}\left[{f_{i}(x)}\right]\mathclose{}$ is finite for every $x\in\mathcal{H}$;
  2. 2.
    the problem is well-posed, in the sense that ${\rm{argmin}}~{}f\neq\emptyset$;
  3. 3.
    the problem is differentiable, in the sense that each $f_{i}$ is differentiable, and so is $f$, with $\nabla f(x)=\mathbb{E}\mathopen{}\left[{\nabla f_{i}(x)}\right]\mathclose{}$.
Within the context of Problem A.1, we study the performance of the SGD algorithm, which generates a sequence $(x_{t})_{t\geq 0}$ according to
$x_{t+1}=x_{t}-\gamma\nabla f_{i_{t}}(x_{t}),$ (SGDE)
where $\gamma>0$ is a fixed step-size, and $i_{t}\in\mathcal{I}$ is sampled i.i.d. from the distribution $\mathcal{D}$.
A.1 Smoothness and convexity
The main assumptions that we present in the core of the paper are smoothness and convexity (or strong convexity) of the functions $f_{i}$. We define a setting combining both convexity and strong convexity, which allows us to unify parts of the analyses. We also define a more relaxed setting, which is all that is needed for our results to hold.
Assumption A.2 (Smoothness and $\mu$-strong convexity).
Considering Problem A.1, we assume that
  • there exists $L\in(0,+\infty)$ such that each function $f_{i}$ is $L$-smooth, in the sense that $\nabla f_{i}\colon\mathcal{H}\to\mathcal{H}$ is $L$-Lipschitz continuous;
  • there exists $\mu\in[0,+\infty)$ such that each function $f_{i}$ is $\mu$-strongly convex. When $\mu=0$, this simply means that $f_{i}$ is convex.
Smoothness and $\mu$-strong convexity can be characterized by the means of a variational inequality:
Lemma A.3 (Expected Cocoercivity).
Let Assumption A.2 hold. Then, for every $(x,y)\in\mathcal{H}^{2}$,
$\ \frac{1}{2}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(y)-\nabla f_{i}(x)\|^{2}}\right]\mathclose{}+\frac{\mu L}{2}\|y-x\|^{2}$
$\leq$ $\ (L-\mu)\left(f(y)-f(x)\right)+\langle\mu\nabla f(y)-L\nabla f(x),y-x\rangle.$ (EC)
Proof.
Because $f_{i}$ is $L$-smooth and $\mu$-convex, we can use (taylor_smooth_2017, Theorem 4) to write, at every $(x,y)$:
$\frac{1}{2}\|\nabla f_{i}(y)-\nabla f_{i}(x)\|^{2}+\frac{\mu L}{2}\|y-x\|^{2}$
$\leq$ $(L-\mu)\left(f_{i}(y)-f_{i}(x)\right)+\langle\mu\nabla f_{i}(y)-L\nabla f_{i}(x),y-x\rangle.$
The conclusion follows after taking expectation. ∎
As will be made clear throughout our proofs, we may relax Assumption A.2 to only assume that the conclusion of Lemma A.3, namely (EC), holds between any point $x$ and any minimizer $x^{*}$. We will decline this assumption into two flavors.
Assumption A.4 (Expected Cocoercivity).
Considering Problem A.1, we assume that there exist $L\in(0,+\infty)$ and $\mu\in[0,L]$ such that
for every $x\in\mathcal{H}$ and every $x_{*}\in{\rm{argmin}}~{}f$, (EC) holds for $(x,x_{*})$. (EC)
Assumption A.5 (Symmetric Expected Cocoercivity).
Considering Problem A.1, we assume that there exist $L\in(0,+\infty)$ and $\mu\in[0,L]$ such that
for every $x\in\mathcal{H}$ and every $x_{*}\in{\rm{argmin}}~{}f$, (EC) holds for $(x,x_{*})$ and $(x_{*},x)$. (SEC)
Remark A.6 (Comparing (SEC) and (EC) with other assumptions in the literature).
From what precedes, it is clear that
each $f_{i}$ is $\mu$-convex and $L$-smooth $\Rightarrow$ (SEC) $\Rightarrow$ (EC).
We also note that (EC) implies the $\mu$-strongly quasi-convexity assumption from gower_sgd_2019. If we focus on the case $\mu=0$, we see that (SEC) holds if, and only if, (EC) and (ES) hold, where the latter is the Expected Smoothness property introduced by gower_sgd_2019, stating that, for all $x\in\mathcal{H}$ and $x_{*}\in{\rm{argmin}}~{}f$,
$\frac{1}{2L}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})-\nabla f_{i}(x)\|^{2}}\right]\mathclose{}\leq f(x)-f(x_{*}).$ (ES)
A.2 Gradient variance at the solution
We also assume that the variance of $\nabla f_{i}(x)$ exists at the solutions. More precisely,
Assumption A.7 (Solution Gradient Variance).
Considering Problem A.1, we assume that the variance at the solution exists, meaning that
$\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]<+\infty$ (GV)
for every $x_{*}\in{\rm{argmin}}~{}f$. We will note
$\sigma_{*}^{2}\coloneqq\sup\limits_{x_{*}\in{\rm{argmin}}~{}f}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}.$
It is important to note that under Assumption A.4 the variance term $\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]$ does not depend on the choice of $x_{*}\in\operatorname{argmin}f$, as was already observed in (garrigos_handbook_2024, Lemma 4.17).
Lemma A.8.
Let Assumption A.4 hold. Then $\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]$ is constant over ${\rm{argmin}}~{}f$.
Proof.
Let $x_{*},x_{*}^{\prime}$ be two minimizers of $f$. From Assumption A.4 we know that (EC) holds true for the couple $(x_{*},x_{*}^{\prime})$, which means that
$\frac{1}{2}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*}^{\prime})-\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$
$\leq$ $(L-\mu)\left(f(x_{*}^{\prime})-f(x_{*})\right)+\langle\mu\nabla f(x_{*}^{\prime})-L\nabla f(x_{*}^{\prime}),y-x\rangle.$
Because $f(x_{*}^{\prime})=f(x_{*})=\inf f$ and $\nabla f(x_{*}^{\prime})=\nabla f(x_{*})=0$, we deduce that
$\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*}^{\prime})-\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}=0.$
Then, almost surely with respect to the distribution $\mathcal{D}$ over the indices $i\in\mathcal{I}$, we have $\|\nabla f_{i}(x_{*}^{\prime})-\nabla f_{i}(x_{*})\|^{2}=0$. This means that $\nabla f_{i}(x_{*}^{\prime})=\nabla f_{i}(x_{*})$ almost surely, from which we deduce that
$\mathbb{E}\left[\|\nabla f_{i}(x_{*}^{\prime})\|^{2}\right]=\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right],$
meaning that in fact $\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]$ is constant over ${\rm{argmin}}~{}f$. ∎
Assuming that the variance at the solution $\sigma_{*}^{2}$ exists is trivially satisfied in the case of the minimization of a finite sum, as presented in Problem (Finite-Sum).
Lemma A.9 (Finite-sum problems have finite solution variance).
If Problem A.1 treats a finite sum of functions, i.e. $\mathcal{I}$ is finite, then Assumption A.7 holds true.
Even for true expectation-minimization problems, where $\mathcal{I}$ is infinite, the Assumption A.7 is very mild. In particular, this assumption is automatically verified in problems where the loss functions $f_{i}$ are nonnegative, which is standard for problems arising in inverse problems and machine learning.
Lemma A.10 (Sufficient condition for finite solution variance).
Consider Problem A.1 and assume that the functions $f_{i}$ are $L$-smooth and are uniformly bounded from below functions: there exists $r\in\mathbb{R}$ such that $f_{i}(x)\geq r$ for every $i\in\mathcal{I}$ and $x\in\mathcal{H}$. Then Assumption A.7 is true, with moreover the information that
$\sigma_{*}^{2}\leq 2L(\min f-r).$
Proof.
Since each $f_{i}$ is $L$-smooth, we can write (garrigos_handbook_2024, Lemma 2.28), for all $x\in\mathcal{H}$,
$\|\nabla f_{i}(x)\|^{2}\leq 2L(f_{i}(x)-\min f_{i})\leq 2L(f_{i}(x)-r).$
Setting $x=x_{*}\in{\rm{argmin}}~{}f$ and taking the expectation of this inequality, we obtain
$\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]\leq 2L(f(x_{*})-r)=2L(\min f-r)<+\infty,$
which concludes since $\sigma_{*}^{2}=\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]$. ∎
A.3 Interpolation
We conclude by connecting the variance at the solution with interpolation. The following facts can be found in garrigos_handbook_2024, which we have extended here from a finite-sum problem to our expectation-minimization Problem A.1.
Definition A.11.
Consider Problem A.1. We say that interpolation holds for the family of functions $(f_{i})_{i\in\mathcal{I}}$ if
$\bigcap\limits_{i\in\mathcal{I}}\ {\rm{argmin}}~{}f_{i}\neq\emptyset,$
where the intersection is to be understood for $\mathcal{D}$-almost every $i\in\mathcal{I}$.
Lemma A.12 (Consequences for interpolation).
Consider Problem A.1. If interpolation holds, then
  1. 1.
    $\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}=\inf f$.
  2. 2.
    $\cap_{i\in\mathcal{I}}\ {\rm{argmin}}~{}f_{i}={\rm{argmin}}~{}f$.
Proof.
Because interpolation holds, we may select $x_{*}\in\cap_{i\in\mathcal{I}}\ {\rm{argmin}}~{}f_{i}$. Let us start by showing that $x_{*}\in{\rm{argmin}}~{}f$. For this, consider any $x\in\mathcal{H}$ and write
$f(x_{*})=\mathbb{E}\mathopen{}\left[{f_{i}(x_{*})}\right]\mathclose{}=\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{f_{i}(x)}\right]\mathclose{}=f(x).$
This proves that $x_{*}\in{\rm{argmin}}~{}f$. Now let us prove the first part, by observing that
$\inf f=f(x_{*})=\mathbb{E}\mathopen{}\left[{f_{i}(x_{*})}\right]\mathclose{}=\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}.$
Secondly, we prove the second point, for which we are only left with the reverse inclusion. Consider some $x\in{\rm{argmin}}~{}f$. Then
$f(x)=\inf f=\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}\ \implies\ \mathbb{E}\mathopen{}\left[{f_{i}(x)-\inf f_{i}}\right]\mathclose{}=0.$
Since $f_{i}(x)-\inf f_{i}\geq 0$ for all $i$, it must hold that $f_{i}(x)-\inf f_{i}=0$ for almost every $i\in\mathcal{I}$, which proves the claim. ∎
Lemma A.13 (Interpolation means $\sigma_{*}^{2}=0$).
Consider Problem A.1.
  1. 1.
    If interpolation holds, then $\sigma_{*}^{2}=0$.
  2. 2.
    The above becomes an equivalence if all the $f_{i}$ are convex.
Proof.
If interpolation holds, then for every $x_{*}\in{\rm{argmin}}~{}f$ we have that $x_{*}\in{\rm{argmin}}~{}f_{i}$ for almost every $i\in\mathcal{I}$. For such $i\in\mathcal{I}$, the optimality condition gives $\nabla f_{i}(x_{*})=0$, which after taking the norm and expectation leads to $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}=0$.
Reciprocally, if we assume that $\sigma_{*}^{2}=0$, it means that there exists $x_{*}\in{\rm{argmin}}~{}f$ such that $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}=0$. This is an expectation of nonnegative terms, such that $\nabla f_{i}(x_{*})=0$ for almost every $i\in\mathcal{I}$. Exploiting our convexity assumption, we deduce that $x_{*}$ is a minimizer of $f_{i}$ for almost every $i\in\mathcal{I}$, which shows that interpolation holds. ∎
Appendix B Reduction to a system of inequalities
B.1 Bounds from a Lyapunov decrease
Let $T\geq 1$ be fixed, let $x_{*}\in{\rm{argmin}}~{}f$, and let $(x_{t})_{t=0}^{T}$ be generated by the SGD algorithm for some fixed step-size $\gamma>0$. Given a set of parameters $\rho,a_{0},\dots,a_{T},e_{0},\dots,e_{T-1}\geq 0$, we define the following Lyapunov energy, for $t=0,\dots,T-1$:
$E_{t}\coloneqq a_{t}\|x_{t}-x_{*}\|^{2}+\rho\sum_{s=0}^{t-1}\left(f(x_{s})-\min f\right)-\sum_{s=0}^{t-1}e_{s}\sigma_{*}^{2},$
where, by convention, the empty sum $\sum_{s=0}^{-1}$ is equal to zero.
The first term of this Lyapunov energy is the distance to the solution $\|x_{t}-x_{*}\|^{2}$, a classical term which typically decreases for deterministic monotone gradient dynamics. The second term involves the function gap $f(x_{t})-\min f$, which also typically decreases for gradient descent. The standard Lyapunov for gradient descent usually contains the term $t(f(x_{t})-\min f)$, which we have replaced with the sum of the past function gaps (observe that both are of the same order in time). The last term for this Lyapunov energy is a negative cumulated sum, where the $e_{t}$’s play the role of a variance term. It is meant to compensate the fluctuations caused by the uncertainty in the SGD algorithm, and will allow the Lyapunov energy to decrease.
The core argument of our analysis is that a decrease of the Lyapunov (in expectation) yields upper bounds for SGD. Let us start by computing the discrete derivative of the energy.
Lemma B.1 (Lyapunov discrete derivative).
With our notations, for every $t=0,\dots,T-1$,
$E_{t+1}-E_{t}=a_{t+1}\|x_{t+1}-x_{*}\|^{2}-a_{t}\|x_{t}-x_{*}\|^{2}+\rho\left(f(x_{t})-\min f\right)-e_{t}\sigma_{*}^{2}.$
Proof.
Immediate from the definition of $E_{t}$. ∎
A first result consists in deriving bounds on the function value gap, provided the Lyapunov parameter $\rho$ is nonzero. This bound will typically be obtained for convex smooth problems.
Lemma B.2 (SGD bound from Lyapunov decrease - Convex).
Consider Problem A.1. Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for every $t=0,\dots,T-1$, and that $\rho>0$. Then
$\min\limits_{t=0,\dots,T-1}\ \mathbb{E}\left[f(x_{t})-\min f\right]\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\bar{e}\sigma_{*}^{2}}{\rho},$
where $\bar{e}=\tfrac{1}{T}\sum_{t=0}^{T-1}e_{t}$. If we further assume that $f$ is convex, then
$\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\bar{e}\sigma_{*}^{2}}{\rho},$
where $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$.
Proof.
Iterating $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for $t=0,\dots,T-1$ yields that $\mathbb{E}\mathopen{}\left[{E_{T}}\right]\mathclose{}\leq E_{0}$, or, equivalently,
$a_{T}\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{0}\|^{2}}\right]\mathclose{}+\rho\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\min f}\right]\mathclose{}-\sum_{t=0}^{T-1}e_{t}\sigma_{*}^{2}\leq a_{0}\|x_{0}-x_{*}\|^{2}.$
Bounding $a_{T}\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{0}\|^{2}}\right]\mathclose{}\geq 0$ and dividing by $\rho T>0$ yields
$\min\limits_{t=0,\dots,T-1}\ \mathbb{E}\left[f(x_{t})-\min f\right]\leq\frac{1}{T}\sum_{t=0}^{T-1}[f(x_{t})-\min f]\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\overline{e}\sigma_{*}^{2}}{\rho}.$
If $f$ is convex, we may use Jensen’s inequality to write $\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]\leq\tfrac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[f(x_{t})-\min f\right]$, which implies the wanted conclusion. ∎
As a second result, a Lyapunov decrease can yield a bound of the square distance of the iterate $x_{t}$ to a minimizer, provided the Lyapunov parameter $a_{T}$ is nonzero. This is standard for strongly convex problems.
Lemma B.3 (SGD bound from Lyapunov decrease - Strongly convex).
Consider Problem A.1, and assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for every $t=0,\dots,T-1$, and that $a_{T}>0$. Then
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{a_{T}}+\frac{e_{T}^{sum}\sigma_{*}^{2}}{a_{T}},$
where $e_{T}^{sum}=\sum_{t=0}^{T-1}e_{t}$.
Proof.
Iterating $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for $t=0,\ldots,T-1$ yields $\mathbb{E}\mathopen{}\left[{E_{T}}\right]\mathclose{}\leq E_{0}$, or in other words that
$a_{T}\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}+\rho\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\min f}\right]\mathclose{}\leq a_{0}\|x_{0}-x_{*}\|^{2}+e_{T}^{sum}\sigma_{*}^{2}.$
Since $f(x_{t})-\min f\geq 0$ for all $t=0,\ldots,T-1$, and $\rho\geq 0$, this yields the wanted inequality after division by $a_{T}>0$. ∎
B.2 Lyapunov decrease from a system of inequalities
Which choices of parameters $\rho,a_{0},\dots,a_{T},e_{0},\dots,e_{T-1}\geq 0$ allow the Lyapunov to decrease? The next theorem provides sufficient conditions which, as we will see later, are seemingly necessary, at least numerically. This is useful, as we will be able to focus on a system of inequalities between real numbers, instead of having to deal with all possible sequences generated by SGD. Finding parameters satisfying these sufficient conditions will be the object of the next sections.
Theorem B.4 (Sufficient conditions for Lyapunov decrease).
Let Assumptions A.4 and A.7 hold true. Assume that there exists parameters $(\alpha_{t},\beta_{t})_{t=0}^{T-1}$ such that, for every $t=0,\dots,T-1$, the following conditions are verified:
  1. 1.
    $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}\geq 0$,
  2. 2.
    $\rho\leq 2(L-\mu)(\alpha_{t}-\beta_{t})$,
  3. 3.
    $a_{t+1}\leq\mu L(\alpha_{t}+\beta_{t})+a_{t}$,
  4. 4.
    $a_{t+1}\gamma^{2}\leq\alpha_{t}+\beta_{t}$,
  5. 5.
    $(a_{t+1}\gamma-\alpha_{t}L-\beta_{t}\mu)^{2}\leq(\mu L(\alpha_{t}+\beta_{t})+a_{t}-a_{t+1})(\alpha_{t}+\beta_{t}-a_{t+1}\gamma^{2})$,
  6. 6.
    $a_{t+1}\gamma^{2}(\alpha_{t}+\beta_{t})\leq(\alpha_{t}+\beta_{t}-a_{t+1}\gamma^{2})e_{t}$.
Then $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for every $t=0,\ldots,T-1$.
Proof.
From Lemma B.1, we know we want to show that
$E_{t+1}-E_{t}=a_{t+1}\mathbb{E}\left[\|x_{t+1}-x_{*}\|^{2}\right]-a_{t}\mathbb{E}\left[\|x_{t}-x_{*}\|^{2}\right]+\rho\mathbb{E}\left[f(x_{t})-\min f\right]-e_{t}\sigma_{*}^{2}\leq 0.$
Using the definition of $x_{t+1}$ allows us to write
$\|x_{t+1}-x_{*}\|^{2}=\|x_{t}-x_{*}\|^{2}+\gamma^{2}\|\nabla f_{i_{t}}(x_{t})\|^{2}-2\gamma\langle\nabla f_{i_{t}}(x_{t}),x_{t}-x_{*}\rangle.$
Therefore,
$E_{t+1}-E_{t}$ $=$ $\rho\left(f(x_{t})-\min f\right)-e_{t}\sigma_{*}^{2}+(a_{t+1}-a_{t})\|x_{t}-x_{*}\|^{2}$
$+a_{t+1}\gamma^{2}\|\nabla f_{i}(x_{t})\|^{2}-2a_{t+1}\gamma\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle.$
To show that this random variable is nonpositive in expectation, we shall make use of an assumption on the loss functions. While it would be natural to exploit the fact that the $f_{i}$ are $\mu$-convex and smooth for all $i\in\mathcal{I}$, our numerical simulations suggest that all we need is (SEC).
In what follows, we denote by $\mathbb{E}_{t}\coloneqq\mathbb{E}\mathopen{}\left[{\cdot\ |\ x_{0},\dots,x_{t}}\right]\mathclose{}$ the conditional expectation with respect to the iterates up to and including $x_{t}$. Property (SEC) reads
$\frac{1}{2}\mathbb{E}_{t}\mathopen{}\left[{\|\nabla f_{i}(x_{*})-\nabla f_{i}(x_{t})\|^{2}}\right]\mathclose{}+\frac{\mu L}{2}\|x_{*}-x_{t}\|^{2}\leq(L-\mu)\left(\min f-f(x_{t})\right)-L\langle\nabla f(x_{t}),x_{*}-x_{t}\rangle$
at the pair $(x_{t},x_{*})$, and
$\frac{1}{2}\mathbb{E}_{t}\mathopen{}\left[{\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}+\frac{\mu L}{2}\|x_{t}-x_{*}\|^{2}\leq(L-\mu)\left(f(x_{t})-\min f\right)+\mu\langle\nabla f(x_{t}),x_{t}-x_{*}\rangle$
at the pair $(x_{*},x_{t})$. Multiplying these inequalities by $2\alpha_{t}\geq 0$ and $2\beta_{t}\geq 0$ respectively, and summing them gives
$({\alpha_{t}+\beta_{t}})\mathbb{E}_{t}\mathopen{}\left[{\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}+{\mu L(\alpha_{t}+\beta_{t})}\|x_{t}-x_{*}\|^{2}$
$\leq$ $2(L-\mu)(\beta_{t}-\alpha_{t})\left(f(x_{t})-\min f\right)+2(\alpha_{t}L+\beta_{t}\mu)\langle\nabla f(x_{t}),x_{t}-x_{*}\rangle.$
Adding $\mathbb{E}_{t}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}-E_{t}$ on both sides and using the expression obtained in (B.2), we get
$\mathbb{E}_{t}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}-E_{t}$$\leq$ $\mathbb{E}_{t}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}-E_{t}$
+ $+\ 2(L-\mu)(\beta_{t}-\alpha_{t})(f(x_{t})-\min f)+2(\alpha_{t}L+\beta_{t}\mu)\langle\nabla f(x_{t}),x_{t}-x_{*}\rangle$
$-\ (\alpha_{t}+\beta_{t})\mathbb{E}_{t}\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{*})\|^{2}-\mu L(\alpha_{t}+\beta_{t})\|x_{t}-x_{*}\|^{2}$
$=$ $\rho\left(f(x_{t})-\min f\right)-e_{t}\sigma_{*}^{2}+(a_{t+1}-a_{t})\|x_{t}-x_{*}\|^{2}$
$+\ a_{t+1}\gamma^{2}\|\nabla f_{i}(x_{t})\|^{2}-2a_{t+1}\gamma\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle$
+ $+\ 2(L-\mu)(\beta_{t}-\alpha_{t})(f(x_{t})-\min f)+2(\alpha_{t}L+\beta_{t}\mu)\langle\nabla f(x_{t}),x_{t}-x_{*}\rangle$
$-\ (\alpha_{t}+\beta_{t})\mathbb{E}_{t}\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{*})\|^{2}-\mu L(\alpha_{t}+\beta_{t})\|x_{t}-x_{*}\|^{2}$
$=$ $(\rho+2(L-\mu)(\beta_{t}-\alpha_{t}))(f(x_{t})-\min f)-e_{t}\sigma_{*}^{2}$
$+(a_{t+1}-a_{t}-\mu L(\alpha_{t}+\beta_{t}))\|x_{t}-x_{*}\|^{2}+\ a_{t+1}\gamma^{2}\|\nabla f_{i}(x_{t})\|^{2}$
$+2(\alpha_{t}L+\beta_{t}\mu-a_{t+1}\gamma)\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle$
$-(\alpha_{t}+\beta_{t})\mathbb{E}_{t}\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{*})\|^{2}.$
Taking the expectation and developing the last square yields
$\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}$$\leq$ $(\rho+2(L-\mu)(\beta_{t}-\alpha_{t}))\mathbb{E}\left[f(x_{t})-\min f\right]$
$+(a_{t+1}-a_{t}-\mu L(\alpha_{t}+\beta_{t}))\mathbb{E}\left[\|x_{t}-x_{*}\|^{2}\right]$
$+\ (a_{t+1}\gamma^{2}-\alpha_{t}-\beta_{t})\mathbb{E}\left[\|\nabla f_{i}(x_{t})\|^{2}\right]$
$-(e_{t}+\alpha_{t}+\beta_{t})\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]$
$+2(\alpha_{t}L+\beta_{t}\mu-a_{t+1}\gamma)\mathbb{E}\left[\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle\right]$
$+2(\alpha_{t}+\beta_{t})\mathbb{E}\left[\langle\nabla f_{i}(x_{t}),\nabla f_{i}(x_{*})\rangle\right].$
For the Lyapunov energy to be decreasing in expectation, it is enough that the right-hand side is nonpositive. For the first term, because $f(x_{t})-\min f\geq 0$, it is enough to assume that $\rho\leq 2(L-\mu)(\alpha_{t}-\beta_{t})$, which corresponds to Condition 2. For the other terms, we enforce that
$A\mathbb{E}\left[\|x_{t}-x_{*}\|^{2}\right]+B\mathbb{E}\left[\|\nabla f_{i}(x_{t})\|^{2}\right]+B^{\prime}\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]$ (2)
$+$ $2C\mathbb{E}\left[\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle\right]+2D\mathbb{E}\left[\langle\nabla f_{i}(x_{t}),\nabla f_{i}(x_{*})\rangle\right]\geq 0,$
where we simplified the expression by introducing the following constants
$\begin{cases}A&=\mu L(\alpha_{t}+\beta_{t})+a_{t}-a_{t+1},\\ B&=\alpha_{t}+\beta_{t}-a_{t+1}\gamma^{2},\\ B^{\prime}&=e_{t}+\alpha_{t}+\beta_{t},\\ C&=a_{t+1}\gamma-\alpha_{t}L-\beta_{t}\mu,\\ D&=-(\alpha_{t}+\beta_{t}).\end{cases}$
The quantity in (2) is simply the expectation of a quadratic polynomial in $X=x_{t}-x_{*}$, $Y=\nabla f_{i}(x_{t})$ and $Y^{\prime}=\nabla f_{i}(x_{*})$. Using elementary linear algebra, it can be shown that for the expression in (2) to be nonnegative, it is enough to require
$A,B,B^{\prime}\geq 0,\quad C^{2}\leq AB,\quad\text{and}\quad D^{2}\leq BB^{\prime}.$
We postpone the proof of this claim to Lemma B.6.
A simple calculation shows that these conditions correspond exactly to the remaining wanted. In fact, $A\geq 0$ corresponds to Condition 3 and $B\geq 0$ to Condition 4. The inequality $B^{\prime}\geq 0$ is trivially satisfied because $e_{t},\alpha_{t},\beta_{t}\geq 0$ by Condition 1. Moreover, $C^{2}\leq AB$ may be developed as
$(a_{t+1}\gamma-\alpha_{t}L-\beta_{t}\mu)^{2}\leq(\mu L(\alpha_{t}+\beta_{t})a_{t}-a_{t+1})(\alpha_{t}+\beta_{t}-a_{t+1}\gamma^{2}),$
which corresponds to Condition 5. Finally, $D^{2}\leq BB^{\prime}$ expands to
$(\alpha_{t}+\beta_{t})^{2}\leq(\alpha_{t}+\beta_{t}-a_{t+1}\gamma^{2})(e_{t}+\alpha_{t}+\beta_{t}),$
which can be simplified to
$a_{t+1}\gamma^{2}(\alpha_{t}+\beta_{t})\leq(\alpha_{t}+\beta_{t}-a_{t+1}\gamma^{2})e_{t},$
which is exactly Condition 6. ∎
Remark B.5 (Relaxing (SEC) when $\beta_{t}=0$).
Looking at how we used (SEC) in the above proof, it is clear that if we impose $\beta_{t}=0$, then one of the two inequalities in (SEC) is useless, and that all we need in that case is (EC). This will be, for instance, the case when proving our bounds for short step-sizes, see later in Remark C.3.
We end this section by proving the technical lemma at the core of the previous proof.
Lemma B.6 (Nonnegative quadratic polynomial).
Let $A,B,B^{\prime},C,D\in\mathbb{R}$ be such that
$A,B,B^{\prime}\geq 0,\quad C^{2}\leq AB,\quad\text{and}\quad D^{2}\leq BB^{\prime}.$ (3)
Let $X,Y,$ and $Y^{\prime}$ be three random variables over $\mathcal{H}$ such that $\mathbb{E}\left[Y^{\prime}\right]=0$. Then
$\mathbb{E}\mathopen{}\left[{A\|X\|^{2}+B\|Y\|^{2}+B^{\prime}\|Y^{\prime}\|^{2}+2C\langle Y,X\rangle-2D\langle Y,Y^{\prime}\rangle}\right]\mathclose{}\geq 0.$
Proof.
Let us note $P$ to be the quantity that we want to see nonnegative. We start by exploiting the fact that $\mathbb{E}\mathopen{}\left[{Y^{\prime}}\right]\mathclose{}=0$ to write that $P=P+\theta\mathbb{E}\mathopen{}\left[{\langle X,Y^{\prime}\rangle}\right]\mathclose{}$, for every $\theta\in\mathbb{R}$. Now, see that for $P\geq 0$ to be true, it is enough for
$\hat{P}_{\theta}(x,y,y^{\prime})=A\|x\|^{2}+B\|y\|^{2}+B^{\prime}\|y^{\prime}\|^{2}+2C\langle y,x\rangle-2D\langle y,y^{\prime}\rangle+2\theta\langle x,y^{\prime}\rangle$
to be nonnegative for every $x,y,y^{\prime}\in\mathcal{H}$ and for some $\theta\in\mathbb{R}$. Our polynomial $\hat{P}_{\theta}$ is equal to $\langle M_{\theta}z,z\rangle$ where $z=(x,y,y^{\prime})$ and
$M_{\theta}=\begin{pmatrix}A&C&\theta\\ C&B&-D\\ \theta&-D&B^{\prime}\end{pmatrix}.$
As such, $\hat{P}_{\theta}\geq 0$ if, and only if, $M_{\theta}\succeq 0$, which, by Sylvester’s criterion, is equivalent to
$A,B,B^{\prime}\geq 0,$ (4a)
$C^{2}\leq AB,\quad\theta^{2}\leq AB^{\prime},\quad D^{2}\leq BB^{\prime},$ (4b)
$A(BB^{\prime}-D^{2})\geq B^{\prime}C^{2}+B\theta^{2}+2CD\theta.$ (4c)
So $P\geq 0$ if there exists $\theta\in\mathbb{R}$ such that Inequalities (4) are verified. We can already see Equations (4a)-(4b) imply Equations (3). Let us prove that in fact the existence of a $\theta\in\mathbb{R}$ satisfying Equations (4) is equivalent to Equations (3).
To prove equivalence, we assume Equations (3) and show that there exists a $\theta\in\mathbb{R}$ such that
$AB^{\prime}\geq\theta^{2}\quad\text{ and }\quad A(BB^{\prime}-D^{2})\geq B^{\prime}C^{2}+B\theta^{2}+2CD\theta.$ (5)
The last inequality is quadratic in $\theta$, and has solutions if, and only if, its discriminant $\Delta$ is nonnegative. A straightforward computation shows that
$\Delta=4(BB^{\prime}-D^{2})(AB-C^{2}),$
which guarantees that $\Delta\geq 0$ given Equations (3). The quadratic may hence be rewritten as
$\frac{-2CD-\sqrt{\Delta}}{2B}\leq\theta\leq\frac{-2CD+\sqrt{\Delta}}{2B}.$
For (5) to have a solution, the above must hold simultaneously with $-\sqrt{AB^{\prime}}\leq\theta\leq\sqrt{AB^{\prime}}$. This is equivalent to the intersection of two intervals to be nonempty, which is the case if
$\begin{cases}\frac{-2CD+\sqrt{\Delta}}{2B}\geq-\sqrt{AB^{\prime}}\quad&\text{if $CD\geq 0$},\\ \frac{-2CD-\sqrt{\Delta}}{2B}\leq\sqrt{AB^{\prime}}\quad&\text{if $CD\leq 0$}.\end{cases}$
or equivalently,
$2|CD|\leq 2B\sqrt{AB^{\prime}}+\sqrt{\Delta}.$
As all terms are nonnegative, this inequality is equivalent to
$4C^{2}D^{2}\leq 4B^{2}AB^{\prime}+4B\sqrt{AB^{\prime}\Delta}+\Delta,$
which holds true since $D^{2}\leq BB^{\prime}$, $C^{2}\leq AB$ and $\Delta\geq 0$. ∎
Appendix C Proofs in the smooth convex setting
In this section, we derive parameters that satisfy the sufficient conditions of Theorem B.4 in order to derive upper bounds for SGD in the convex setting. The propositions in this section can be verified symbolically, yet we include the proofs to provide intuition on how these parameters were derived.
C.1 Bounds for short step-sizes: $\gamma L<1$
Proposition C.1 (Lyapunov parameters. Convex case, short step-sizes).
Let $\gamma L\in(0,1)$, and consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\rho=2\gamma+\tfrac{2(1-\gamma L)}{LT}$,
  • $a_{t}=\tfrac{T-t}{T}\tfrac{1+\gamma L(T-1)}{1+\gamma L(T-t-1)}$,
  • $e_{t}=\frac{a_{t+1}\gamma^{2}\alpha_{t}}{\alpha_{t}-a_{t+1}\gamma^{2}}$,
  • $\alpha_{t}\equiv\alpha=\tfrac{\rho}{2L}$,
  • $\beta_{t}\equiv\beta=0$.
These parameters satisfy the sufficient Lyapunov conditions of Theorem B.4. Moreover, it holds that
$e_{t}=\frac{\gamma^{2}}{1-\gamma L}\frac{T-t-1}{T}\frac{(1-\gamma L)+\gamma LT}{(1-\gamma L)+\gamma L(T-t)}\leq\frac{\gamma^{2}}{1-\gamma L}.$
Proof.
Notice that the sufficient conditions from Theorem B.4 are homogeneous. Because our goal is to obtain bounds whose bias term is the smallest possible, we can without loss of generality impose that $a_{0}=1$, and try to maximize $\rho$. Through our numerical analysis, we empirically observe that we can take $\alpha_{t}\equiv\alpha>0$ and $\beta_{t}\equiv\beta=0$, thus justifying these choices. We dedicate the remainder of the proof to justify our remaining choices.
We start by focusing on Condition 5 from Theorem B.4, which reads
$\left(a_{t+1}\gamma-L{\alpha}\right)^{2}\leq({\alpha}-a_{t+1}\gamma^{2})(a_{t}-a_{t+1}).$
Rearranging this inequality yields
$\iff$ $\left(a_{t+1}\gamma-L{\alpha}\right)^{2}\leq(1-\gamma L)\alpha(a_{t}-a_{t+1})+(\gamma L\alpha-a_{t+1}\gamma^{2})(a_{t}-a_{t+1})$
$\iff$ $\left(a_{t+1}\gamma-L{\alpha}\right)^{2}-\gamma(L\alpha-a_{t+1}\gamma)(a_{t}-a_{t+1})\leq(1-\gamma L)\alpha(a_{t}-a_{t+1})$
$\iff$ $\left(L\alpha-a_{t+1}\gamma\right)(L\alpha-a_{t+1}\gamma-\gamma(a_{t}-a_{t+1}))\leq(1-\gamma L)\alpha(a_{t}-a_{t+1})$
$\iff$ $\left(L\alpha-a_{t+1}\gamma\right)(L\alpha-a_{t}\gamma)\leq(1-\gamma L)\alpha(a_{t}-a_{t+1}).$
Condition 3 boils down to $a_{t}\geq a_{t+1}$. By additionally imposing $a_{t}>a_{t+1}$, we may divide by $a_{t}-a_{t+1}$ to obtain
$\iff$ $\frac{\left(L\alpha-a_{t+1}\gamma\right)(L\alpha-a_{t}\gamma)}{a_{t}-a_{t+1}}\leq(1-\gamma L)\alpha$
$\iff$ $\frac{\left(L\alpha-a_{t+1}\gamma\right)(L\alpha-a_{t}\gamma)}{(L\alpha-a_{t+1}\gamma)-(L\alpha-a_{t}\gamma)}\leq\frac{(1-\gamma L)\alpha}{\gamma}.$
As $(a_{t})$ is decreasing, with $a_{0}=1$, it holds true that $a_{t}\leq 1$ for all $t=0,\ldots,T-1$. Specifically, if $L\alpha>\gamma$ (which we shall assume), then both $L\alpha-a_{t+1}\gamma>0$ and $L\alpha-a_{t}\gamma>0$, thus allowing us to write
$\iff$ $\frac{(L\alpha-a_{t+1}\gamma)-(L\alpha-a_{t}\gamma)}{\left(L\alpha-a_{t+1}\gamma\right)(L\alpha-a_{t}\gamma)}\geq\frac{\gamma}{(1-\gamma L)\alpha}$
$\iff$ $\frac{1}{L\alpha-a_{t}\gamma}-\frac{1}{L\alpha-a_{t+1}\gamma}\geq\frac{\gamma}{(1-\gamma L)\alpha}$
$\iff$ $\frac{1}{L\alpha-a_{t+1}\gamma}\leq\frac{1}{L\alpha-a_{t}\gamma}-\frac{\gamma}{(1-\gamma L)\alpha}.$
Introducing the temporary variables $u_{t}\coloneqq\tfrac{1}{L\alpha-\gamma a_{t}}$ and $c=\tfrac{\gamma}{(1-\gamma L)\alpha}$, we see that we obtained an arithmetic inequality $u_{t+1}\leq u_{t}-c$, such that $u_{t}\leq u_{0}-ct$. We know that $u_{0}=\tfrac{1}{L\alpha-\gamma}$, so we can write
$u_{t}\leq\frac{1}{L\alpha-\gamma}-\frac{t\gamma}{(1-\gamma L)\alpha}=\frac{(1-\gamma L)\alpha-t\gamma(L\alpha-\gamma)}{(1-\gamma L)\alpha(L\alpha-\gamma)}.$
Returning to the definition of $u_{t}$ we deduce the following bound for $a_{t}$:
$a_{t}$$\leq$ $\frac{1}{\gamma}\left(L\alpha-\frac{(1-\gamma L)\alpha(L\alpha-\gamma)}{(1-\gamma L)\alpha-t\gamma(L\alpha-\gamma)}\right)$
$=$ $\frac{1}{\gamma}\frac{L\alpha^{2}(1-\gamma L)-L\alpha t\gamma(L\alpha-\gamma)-(1-\gamma L)\alpha(L\alpha-\gamma)}{(1-\gamma L)\alpha-t\gamma(L\alpha-\gamma)}$
$=$ $\frac{1}{\gamma}\frac{-L\alpha t\gamma(L\alpha-\gamma)+\alpha\gamma(1-\gamma L)}{(1-\gamma L)\alpha-t\gamma(L\alpha-\gamma)}$
$=$ $\alpha\frac{(1-\gamma L)-tL(L\alpha-\gamma)}{(1-\gamma L)\alpha-t\gamma(L\alpha-\gamma)}.$
We note that our Condition 1 requires that $a_{t}\geq 0$ for all $t=0,\ldots,T$, which is equivalent to having $a_{T}\geq 0$ as we assumed $(a_{t})$ to be decreasing. Specifically, we must have
$(1-\gamma L)-TL(L\alpha-\gamma)\geq 0\quad\iff\quad\alpha\leq\frac{1-\gamma L+\gamma TL}{TL^{2}}.$
As we want $\rho$ as large as possible, and due to Condition 2 which reads $\rho\leq 2L(\alpha-\beta)$ with $\beta=0$, we want $\alpha$ to be as large as possible. Therefore we fix
$\alpha=\frac{1-\gamma L+T\gamma L}{TL^{2}}\quad\text{ and }\quad\rho=2L\alpha=\frac{2(1-\gamma L)+2T\gamma L}{TL}.$
This choice respects the assumption $L\alpha>\gamma$ we made earlier, since $01-\gamma L>$ 0.
To get an expression for $a_{t}$, we simply set the inequalities to equalities, and replace $\alpha$ by its chosen value. We then obtain
$L\alpha-\gamma=\frac{(1-\gamma L)+\gamma LT-\gamma LT}{TL}=\frac{1-\gamma L}{TL}\Rightarrow u_{0}=\frac{TL}{1-\gamma L}.$
So
$u_{t}$ $=$ $u_{0}-tc=\frac{TL}{1-\gamma L}-\frac{t\gamma}{\alpha(1-\gamma L)}=\frac{\alpha TL-t\gamma}{\alpha(1-\gamma L)}$
$=$ $\frac{\tfrac{1}{L}(1-\gamma L)+\gamma T-t\gamma}{\alpha(1-\gamma L)}=\gamma\frac{T-t-1+\tfrac{1}{\gamma L}}{\alpha(1-\gamma L)}.$
We now use the fact that $\gamma a_{t}=L\alpha-u_{t}^{-1}$ to write
$a_{t}$ $=$ $\frac{L\alpha}{\gamma}-\frac{1}{\gamma u_{t}}=\frac{\alpha L}{\gamma}-\frac{1}{L\gamma}\frac{(1-\gamma L)\tfrac{\alpha L}{\gamma}}{T-t-1+\tfrac{1}{\gamma L}}$
$=$ $\frac{\alpha L}{\gamma}\left(1-\frac{1}{\gamma L}\frac{(1-\gamma L)}{T-t-1+\tfrac{1}{\gamma L}}\right)=\frac{\alpha L}{\gamma}\left(1-\frac{(1-\gamma L)}{\gamma L(T-t-1)+1}\right)$
$=$ $\frac{\alpha L}{\gamma}\frac{\gamma L(T-t)}{\gamma L(T-t-1)+1}=\frac{(1-\gamma L)+\gamma LT}{T\gamma L}\frac{\gamma L(T-t)}{\gamma L(T-t-1)+1}$
$=$ $\frac{(T-t)}{T}\frac{1+\gamma L(T-1)}{\gamma L(T-t-1)+1}.$
To conclude the proof, we need to check all the sufficient Lyapunov conditions of Theorem B.4. Conditions 1, 2, 3 and 5 are already satisfied, by construction. It remains to check Conditions 4 and 6.
Condition 4 is equivalent to
$a_{t+1}\gamma^{2}\leq\alpha$ $\iff$ $\gamma^{2}\frac{(T-t-1)}{T}\frac{1+\gamma L(T-1)}{1+\gamma L(T-t-2)}\leq\frac{(1-\gamma L)+\gamma LT}{TL^{2}}$
$\iff$ $\frac{\gamma^{2}L^{2}(T-t-1)}{1+\gamma L(T-t-2)}(1+\gamma L(T-1))\leq 1+\gamma L(T-1)$
$\iff$ $\gamma^{2}L^{2}(T-t-1)\leq 1+\gamma L(T-t-2)=1+\gamma L(T-t-1)-\gamma L$
$\iff$ $0\leq 1-\gamma L+\gamma L(T-t-1)(1-\gamma L),$
which is true for every $t=0,\dots,T-1$.
Condition 6 is a lower bound on $e_{t}$, of which the equality case coincides with the chosen value of $e_{t}$. More specifically, this gives
$e_{t}=\frac{a_{t+1}\gamma^{2}\alpha}{\alpha-a_{t+1}\gamma^{2}}=\frac{\gamma^{2}}{\frac{1}{a_{t+1}}-\frac{\gamma^{2}}{\alpha}},$
where
$\frac{1}{a_{t+1}}-\frac{\gamma^{2}}{\alpha}$ $=$ $\frac{T}{T-t-1}\frac{(1-\gamma L)+\gamma L(T-t-1)}{(1-\gamma L)+\gamma LT}-\frac{T\gamma^{2}L^{2}}{(1-\gamma L)+\gamma LT}$
$=$ $\frac{T}{T-t-1}\frac{(1-\gamma L)+\gamma L(T-t-1)}{(1-\gamma L)+\gamma LT}-\frac{T\gamma^{2}L^{2}}{(1-\gamma L)+\gamma LT}$
$=$ $T\frac{(1-\gamma L)+\gamma L(T-t-1)-\gamma^{2}L^{2}(T-t-1)}{(T-t-1)((1-\gamma L)+\gamma LT)}$
$=$ $T\frac{(1-\gamma L)((1-\gamma L)+\gamma L(T-t))}{(T-t-1)((1-\gamma L)+\gamma LT)}.$
We thus get
$e_{t}$ $=$ $\frac{\gamma^{2}}{1-\gamma L}\frac{T-t-1}{T}\frac{(1-\gamma L)+\gamma LT}{(1-\gamma L)+\gamma L(T-t)}.$
We readily verify that $e_{t}$ is nonincreasing with respect to $t$. Indeed, its derivative with respect to $t$ has the same sign as
$(T-1)((1-\gamma L)+\gamma LT)\gamma L-((1-\gamma L)+\gamma LT)^{2},$
which is always nonpositive. Therefore $e_{t}\leq e_{0}$, where
$e_{0}=\frac{\gamma^{2}}{1-\gamma L}\frac{T-1}{T}\leq\frac{\gamma^{2}}{1-\gamma L},$
thus concluding the proof. ∎
Theorem C.2 (Bounds for SGD. Convex case, short step-sizes).
Let Assumptions A.4 and A.7 hold, with $\mu=0$. Let $x_{t}$ be generated by SGD, with $\gamma L\in(0,1)$. Then for every $T\geq 1$
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\inf f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\gamma^{2}\sigma_{*}^{2}}{\rho(1-\gamma L)}\bar{e}_{T}^{\prime},$
where $\rho=2\gamma+\tfrac{2(1-\gamma L)}{LT}$ and
$\bar{e}_{T}^{\prime}=\frac{1}{T}\sum_{t=0}^{T-1}e_{t}^{\prime},\quad\text{with$e_{t}^{\prime}=\frac{T-t-1}{T}\frac{(1-\gamma L)+\gamma LT}{(1-\gamma L)+\gamma L(T-t)}$}.$
Moreover, we have the simpler bounds
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\inf f}\right]\mathclose{}\leq\frac{L\|x_{0}-x_{*}\|^{2}}{2\gamma LT+2(1-\gamma L)}+\frac{\gamma\sigma_{*}^{2}}{2(1-\gamma L)}\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\frac{\gamma\sigma_{*}^{2}}{2(1-\gamma L)}.$
If moreover $f$ is convex, the above bounds hold true for $\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]$, where $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$.
Proof.
We simply combine Proposition C.1 with Theorem B.4 and Lemma B.2. We moreover exploit the fact that $\beta_{t}=0$ in Proposition C.1, which allows us to replace Assumption A.5 with Assumption A.4 in Theorem B.4, see Remark B.5 for more details. In the statement, we introduced the notation $e_{t}^{\prime}$ such that $e_{t}=\tfrac{\gamma^{2}}{1-\gamma L}e_{t}^{\prime}$. ∎
Remark C.3.
We note that the parameters selected in Proposition C.1 include $\beta_{t}\equiv 0$, which, in line with Remark B.5, justify that we only require Assumption A.4 instead of Assumption A.5
C.2 Bounds for the optimal step-size: $\gamma L=1$
Considering the result of Theorem C.2 and letting $\gamma L\to 1$, we note that the bias term tends to $\tfrac{1}{2\gamma T}$ while the variance term diverges since its denominator is a multiple of $1-\gamma L$. This suggests there might be a complication to obtain tight bounds in the case $\gamma L=1$. As we will see later in Section F, we have empirical results showcasing that it is not possible to obtain a bound for SGD where the bias term is $\tfrac{1}{2\gamma T}$ for $\gamma L=1$ whilst preserving a finite variance term. Therefore, in this section, we will prove bounds where the bias term is given by $\frac{1}{\rho T}$, with $\rho<2\gamma$.
Proposition C.4 (Lyapunov parameters. Convex case, optimal step-size).
Let $\gamma L=1$ and $\varepsilon\in(0,2)$, and consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\rho=(2-\varepsilon)\gamma$,
  • $a_{t}\equiv a=1$,
  • $\alpha_{t}=\gamma^{2}$,
  • $\beta_{t}=\tfrac{\gamma^{2}\varepsilon}{2}$,
  • $e_{t}=\gamma^{2}\tfrac{2+\varepsilon}{\varepsilon}$.
These parameters satisfy the sufficient Lyapunov conditions of Theorem B.4.
Proof.
First, let us impose that $a_{0}=1$ without loss of generality. Second, we set $\rho=(2-\varepsilon)\gamma$ for $\varepsilon>0$ since numerical experiments suggest that $\rho=2\gamma$ is impossible (see Section F.2). Third, taking inspiration from our numerical findings, we impose that the sequence $a_{t}$ is constantly equal to $1$, that is $a_{t}\equiv a=1$.
If we want to satisfy the sufficient Lyapunov conditions of Theorem B.4, we need in particular to verify Condition 5, which reduces to
$\gamma=L\alpha_{t},$
because $a_{t}-a_{t+1}=0$. This means that $\alpha_{t}\equiv\alpha=\gamma^{2}$, since $\gamma L=1$. Condition 2 implies that $\beta_{t}$ must satisfy
$\beta_{t}\leq\alpha-\frac{\rho}{2L}=\frac{\gamma^{2}\varepsilon}{2}.$
Condition 4 is trivially satisfied since it requires
$\gamma^{2}\leq\alpha_{t}+\beta_{t}=\gamma^{2}+\beta_{t},$
that is $\beta_{t}\geq 0$. It remains to verify Condition 6, which requires that
$e_{t}\geq\frac{\gamma^{2}(\alpha+\beta_{t})}{\alpha+\beta_{t}-\gamma^{2}}.$
First, observe that this is the only constraint on $e_{t}$, which we want to see as small as possible. So we shall set the inequality to be an equality. Second, the expression is decreasing with respect to $\beta_{t}$ (its derivative has the same sign as $-\gamma^{2}$). Therefore, $e_{t}$ will be minimal if $\beta_{t}$ is maximal, meaning we shall fix $\beta_{t}$ to constantly be equal to its earlier derived upper bound, that is $\beta_{t}\equiv\beta=\tfrac{\gamma^{2}\varepsilon}{2}$. Finally, we compute
$e_{t}=\gamma^{2}\frac{\gamma^{2}+\tfrac{\gamma^{2}\varepsilon}{2}}{\gamma^{2}+\tfrac{\gamma^{2}\varepsilon}{2}-\gamma^{2}}=\gamma^{2}\frac{2+{\varepsilon}}{{\varepsilon}}.$
Note in particular that $e_{t}$ is constant here. ∎
Theorem C.5 (Bounds for SGD. Convex case, optimal step-size).
Let Assumptions A.5 and A.7 hold, with $\mu=0$. Let $x_{t}$ be generated by SGD, with $\gamma L=1$. Then, for every $T\geq 1$ and $\varepsilon\in(0,2)$,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\inf f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}}{(2-\varepsilon)\gamma T}+\frac{(2+\varepsilon)\gamma\sigma_{*}^{2}}{\varepsilon(2-\varepsilon)}.$
If moreover $f$ is convex, the above bound holds true for $\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]$, where $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$.
Proof.
We simply combine Proposition C.4 with Theorem B.4 and Lemma B.2. ∎
In Section 4 we explained that our numerical experiments suggest that a bias with $\rho=2\gamma$, where $\gamma L=1$, is not achievable whilst keeping the variance finite. Theorem B.4 provides sufficient conditions on the parameters $\left(\rho,(a_{t}),(e_{t})\right)$ for the Lyapunov decrease to happen. Those sufficient conditions are the ones we use throughout the paper, and allow us to theoretically match the bounds obtained numerically. We therefore have good reasons to believe that those sufficient conditions are very close to be necessary. In what follows, we show that the system of inequalities in Theorem B.4 is not feasible for the optimal value $\rho=2\gamma$ when $\gamma L=1$.
Proposition C.6.
Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for all $t=0,\dots,T-1$. If $\gamma L=1$, $a_{0}=1$ and $\rho=2\gamma$, then it must hold true that $a_{t}\geq 1$ for all $t=0,\ldots,T$.
Proof.
We consider $d=1$ and $m=2$, and let $f\in\mathcal{F}_{0,L}(\mathbb{R})$ be defined through
$f(x)=\frac{1}{2}\left(f_{+}(x)+f_{-}(x)\right)=\frac{1}{2}\left(\frac{L}{2}(x+\delta)^{2}+\frac{L}{2}(x-\delta)^{2}\right).$
One can easily compute that
$f(x)=\frac{L\left(x^{2}+\delta^{2}\right)}{2},\quad\min f=\frac{L\delta^{2}}{2},\quad f(x)-\min f=\frac{Lx^{2}}{2},\quad x_{*}=0$
$\nabla f(x)=Lx,\quad\nabla f_{\pm}(x)=L(x\pm\delta),\quad\nabla f_{\pm}(x_{*})=\pm L\delta,$
We know that $\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}\leq 0$ for all $t=0,\ldots,T-1$ and for all $x_{0}\in\mathbb{R}$. Expanding the Lyapunov decrease and rearranging yields
$a_{t+1}\mathbb{E}_{t}\mathopen{}\left[{|x_{t+1}|^{2}}\right]\mathclose{}-a_{t}|x_{t}|^{2}+\rho\left(f(x_{t})-\min f\right)\leq e_{t}\sigma_{*}^{2}$
$\Longleftrightarrow$ $~{}a_{t+1}\left((1-\gamma L)^{2}|x_{t}|^{2}+(\gamma L)^{2}\delta^{2}\right)-a_{t}|x_{t}|^{2}+\rho L\left(\frac{|x_{t}|^{2}+\delta^{2}}{2}-\frac{\delta^{2}}{2}\right)\leq 2e_{t}\delta^{2}$
$\Longleftrightarrow$ $\left(a_{t+1}(1-\gamma L)^{2}-a_{t}+\frac{\rho L}{2}\right)|x_{t}|^{2}\leq\left(2e_{t}-a_{t+1}(\gamma L)^{2}\right)\delta^{2}.$
This inequality is to be satisfied for all $\delta$ and $x_{t}$. Specifically, the inequality must hold for $(x_{t},\delta)=(1,0)$, showing that $a_{t+1}(1-\gamma L)^{2}-a_{t}+\rho L/2\leq 0$. Setting $\rho=2\gamma$, and $\gamma L=1$ yields $a_{t}\geq 1$. ∎
Corollary C.7.
The sufficient conditions of Theorem B.4 cannot be verified with $\rho=2\gamma$, $a_{0}=1$ and $\gamma L=1$.
Proof.
Let $\rho,(e_{t},a_{t})_{t=0}^{T}$ be parameters, normalized with $a_{0}=1$, and suppose $\gamma L=1$ and $\rho=2\gamma$. Suppose that there exists $(\alpha_{t},\beta_{t})$ such that the system of inequality in Theorem B.4 holds. From Proposition C.6, we know that $a_{t}\geq 1$ for all $t=0,\ldots,T$. Moreover, from Condition 3 and a simple induction argument, we deduce that $a_{t}\leq a_{0}=1$. Hence, $a_{t}\equiv 1$. This together with Condition 5 yields
$\left(\frac{1}{L}-\alpha_{t}L\right)^{2}\leq 0\iff\alpha_{t}=\frac{1}{L^{2}}.$
Plugging this value of $\alpha_{t}$ in Condition 2 gives
$\frac{2}{L}\leq 2L\left(\frac{1}{L^{2}}-\beta_{t}\right)\Longrightarrow\beta_{t}\leq 0\Longrightarrow\beta_{t}=0.$
Finally by injecting the values of $\alpha_{t}$ and $\beta_{t}$ into Condition 6, we obtain
$\frac{1}{L^{4}}\leq 0.$
This contradicts $L>0$, thus concluding. ∎
C.3 Bounds for large step-sizes $\gamma L>1$
Proposition C.8 (Lyapunov parameters. Convex case, large step-sizes).
Let $21<\gamma L<$ 2, and consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\rho=\tfrac{2\gamma(2-\gamma L)}{1-\theta^{T}}$ where $\theta=(1-\gamma L)^{2}$,
  • $a_{t}=\tfrac{1-\theta^{T-t}}{1-\theta^{T}}$,
  • $\beta_{t}=\tfrac{\gamma(\gamma L-1)(1-\theta^{T-t-1})}{L(1-\theta^{T})}$,
  • $\alpha_{t}=\gamma\tfrac{(2-\gamma L)+(\gamma L-1)(1-\theta^{T-t-1})}{L(1-\theta^{T})}$,
  • $e_{t}=\tfrac{\gamma^{2}}{2-\gamma L}\tfrac{1-\theta^{T-t-1}}{1-\theta^{T}}\left(\tfrac{\gamma L}{\theta^{T-t-1}}-2(\gamma L-1)\right)$.
These parameters satisfy the sufficient Lyapunov conditions of Theorem B.4. Moreover, the averaged variance $\bar{e}_{T}\coloneqq\tfrac{1}{T}\sum_{t=0}^{T-1}e_{t}$ has the following asymptotic equivalence as $T\to+\infty$:
$\bar{e}_{T}\sim\frac{\gamma^{2}}{(2-\gamma L)^{2}}\frac{1}{T(1-\gamma L)^{2T-2}}\to+\infty.$
Proof.
To determine the parameters, we will rely on insights from our numerical experiments. Specifically, we were able to guess that
$\beta_{t}=\frac{a_{t+1}\gamma(\gamma L-1)}{L},$
and that Condition 2 is satisfied with equality, meaning that
$\rho=2L(\alpha_{t}+\beta_{t})\iff\alpha_{t}=\frac{\rho}{2L}+\beta_{t},$
Considering Condition 5, we have
$(a_{t+1}\gamma-L\alpha_{t})^{2}\leq(\alpha_{t}+\beta_{t}-\gamma^{2}a_{t+1})(a_{t}-a_{t+1}).$
The latter will give us an expression of $a_{t}$ in terms of $a_{t+1}$, from which we will deduce a general expression for $a_{t}$ by induction. By injecting the values of $\alpha_{t}$ and $\beta_{t}$, we obtain
$\frac{1}{L^{2}}\left(\frac{\rho L}{2}-a_{t+1}\gamma L(2-\gamma L)\right)^{2}\leq\frac{1}{L^{2}}\left(\frac{\rho L}{2}-a_{t+1}\gamma L(2-\gamma L)\right)(a_{t}-a_{t+1}).$
This is satisfied when $\rho>2\gamma L(2-\gamma L)$ and
$\frac{\rho L}{2}-a_{t+1}\gamma L(2-\gamma L)\leq a_{t}-a_{t+1}\iff a_{t}\geq\frac{\rho L}{2}+a_{t+1}(\gamma L-1)^{2}.$
We observe an arithmetic-geometric relation given by, where we set $\theta=(\gamma L-1)^{2}$,
$a_{t}\geq\frac{\rho L}{2}+a_{t+1}\theta\iff a_{t+1}\leq\theta^{-1}a_{t}-\frac{\rho L}{2\theta}.$
A simple induction argument leads to
$a_{t}\leq\theta^{-t}a_{0}-\frac{\rho L}{2\theta}\frac{\theta^{-t}-1}{\theta^{-1}-1}.$
As done previously, without loss of generality, we impose $a_{0}=1$. We impose the sequence to be nonincreasing and nonnegative. The latter leads to
$a_{T}\geq 0\iff\frac{\rho L}{2\theta}\frac{\theta^{-T}-1}{\theta^{-1}-1}\leq\theta^{-T}\iff\frac{\rho L}{2}\leq\theta^{-T}\frac{1-\theta}{\theta^{-T}-1}=\frac{1-\theta}{1-\theta^{T}}.$
As we aim to maximize $\rho$, we impose
$\frac{\rho L}{2}=\frac{1-\theta}{1-\theta^{T}}\iff\rho=\frac{2(1-\theta)}{L(1-\theta^{T})}=\frac{2\gamma(2-\gamma L)}{1-\theta^{T}}.$
In particular, we obtain
$a_{t}\leq\theta^{-t}-\frac{\rho L}{2}\frac{\theta^{-t}-1}{1-\theta}=\theta^{-t}-\frac{1-\theta}{1-\theta^{T}}\frac{\theta^{-t}-1}{1-\theta}=\theta^{-t}-\frac{\theta^{-t}-1}{1-\theta^{T}}=\frac{1-\theta^{T-t}}{1-\theta^{T}}.$
To enforce that $(a_{t})$ is nonincreasing, we set the above to be an equality, and since $a_{T}\geq 0$, it is also nonnegative, as wanted.
We now verify the remaining conditions from Theorem B.4. The nonnegativity of $a_{t}$ implies the nonnegativity of $\alpha_{t},\beta_{t},$. Conditions 1, 3, 5 are verified by construction. It remains to check that Conditions 4 and 6 hold true.
In the following, we introduce $k\coloneqq T-t-1$ to simplify notation. To check that Condition 4 is satisfied, we compute $\alpha_{t}+\beta_{t}$:
$\alpha_{t}+\beta_{t}=\frac{\rho}{2L}+2\beta_{t}=\frac{\gamma(2-\gamma L)+2\gamma(\gamma L-1)(1-\theta^{k})}{L(1-\theta^{T})}.$
Therefore, Condition 4, which states $\alpha_{t}+\beta_{t}-\gamma^{2}a_{t+1}\geq 0$, becomes
00 0 $\leq$ $\frac{\gamma(2-\gamma L)+2\gamma(\gamma L-1)(1-\theta^{k})}{L(1-\theta^{T})}-\gamma^{2}\frac{1-\theta^{k}}{1-\theta^{T}}$
$=$ $\frac{\gamma(2-\gamma L)+2\gamma(\gamma L-1)(1-\theta^{k})-\gamma^{2}L(1-\theta^{k})}{L(1-\theta^{T})}$
$=$ $\frac{\gamma(2-\gamma L)-(1-\theta^{k})\gamma(2-\gamma L)}{L(1-\theta^{T})}$
$=$ $\frac{\gamma(2-\gamma L)\theta^{k}}{L(1-\theta^{T})},$
which is true because $\gamma L\in(1,2)$. It remains to study Condition 6, which is satisfied by setting
$e_{t}$ $=$ $\gamma^{2}\frac{a_{t+1}(\alpha_{t}+\beta_{t})}{\alpha_{t}+\beta_{t}-\gamma^{2}a_{t+1}}$
$=$ $\gamma^{2}\frac{1-\theta^{k}}{1-\theta^{T}}\frac{\gamma(2-\gamma L)+2\gamma(\gamma L-1)(1-\theta^{k})}{L(1-\theta^{T})}\frac{L(1-\theta^{T})}{\gamma(2-\gamma L)\theta^{k}}$
$=$ $\gamma^{2}\frac{1-\theta^{k}}{1-\theta^{T}}\frac{(2-\gamma L)+2(\gamma L-1)(1-\theta^{k})}{(2-\gamma L)\theta^{k}}$
$=$ $\gamma^{2}\frac{1-\theta^{k}}{1-\theta^{T}}\frac{\gamma L-2(\gamma L-1)\theta^{k}}{(2-\gamma L)\theta^{k}}$
$=$ $\frac{\gamma^{2}}{2-\gamma L}\frac{1-\theta^{k}}{1-\theta^{T}}\left(\frac{\gamma L}{\theta^{k}}-2(\gamma L-1)\right).$
To conclude the proof, it remains to analyze the average of the $e_{t}$’s. Denote $e_{t}^{\prime\prime}=({1-\theta^{k}})\left(\frac{\gamma L}{\theta^{k}}-2(\gamma L-1)\right)$, such that
$\sum_{t=0}^{T-1}e_{t}^{\prime\prime}$ $=$ $\gamma L\sum_{t=0}^{T-1}\frac{1-\theta^{k}}{\theta^{k}}-2(\gamma L-1)\sum_{t=0}^{T-1}(1-\theta^{k})$
$=$ $\gamma L\sum_{t=0}^{T-1}(\theta^{-k}-1)-2(\gamma L-1)\sum_{t=0}^{T-1}(1-\theta^{k})$
$=$ $\gamma L\left(\sum_{t=0}^{T-1}\theta^{-k}\right)-T\gamma L-2(\gamma L-1)T+2(\gamma L-1)\left(\sum_{t=0}^{T-1}\theta^{k}\right)$
$=$ $\gamma L\left(\sum_{t=0}^{T-1}\theta^{-t}\right)-T(3\gamma L-2)+2(\gamma L-1)\left(\sum_{t=0}^{T-1}\theta^{t}\right)$
$=$ $\gamma L\frac{1-\theta^{-T}}{1-\theta^{-1}}-T(3\gamma L-2)+2(\gamma L-1)\frac{1-\theta^{T}}{1-\theta}$
$=$ $\frac{\theta}{2-\gamma L}(\theta^{-T}-1)-T(3\gamma L-2)+\frac{2(\gamma L-1)}{\gamma L(2-\gamma L)}(1-\theta^{T}),$
which is asymptotically equivalent to $\tfrac{\theta}{2-\gamma L}\theta^{-T}$. Therefore we can write
$\bar{e}_{T}=\frac{\gamma^{2}}{2-\gamma L}\frac{1}{1-\theta^{T}}\frac{1}{2-\gamma L}\bar{e}_{T}^{\prime},$
where $e_{t}^{\prime}=(2-\gamma L)e_{t}^{\prime\prime}$ and
$\bar{e}_{T}^{\prime}=\frac{1}{T}\sum_{t=0}^{T-1}e_{t}^{\prime}=\frac{\theta}{T}(\theta^{-T}-1)-(2-\gamma L)(3\gamma L-2)+\frac{2(\gamma L-1)}{\gamma LT}(1-\theta^{T}).$ (6)
In particular, one sees that $\bar{e}_{T}^{\prime}\sim\frac{1}{T\theta^{T-1}}$. ∎
Theorem C.9 (Bounds for SGD. Convex case, large step-sizes).
Let Assumptions A.5 and A.7 hold, with $\mu=0$. Let $x_{t}$ be generated by SGD, with $\gamma L\in(1,2)$. Then, for every $T\geq 1$,
$\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\inf f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}\delta}{2\gamma(2-\gamma L)T}+\frac{\gamma\sigma_{*}^{2}\bar{e}_{T}^{\prime}}{2(2-\gamma L)^{3}},$
where $\delta=1-(1-\gamma L)^{2T}\in(0,1)$, and $\bar{e}_{T}^{\prime}$ is defined in (6) and grows exponentially with $T$. If moreover $f$ is convex, the above bound holds true for $\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]$, where $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$.
Proof.
We simply combine Proposition C.8 with Theorem B.4 and Lemma B.2. ∎
We may moreover show that the bias term in Theorem C.9 is tight, in the sense that no Lyapunov proof using energies of our form may improve upon the bias term.
Proposition C.10.
Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for all $t=0,\dots,T-1$. If $\gamma L>1$ and $a_{0}=1$, then necessarily
$\rho\leq\frac{2\gamma(2-\gamma L)}{1-(1-\gamma L)^{2T}}.$
Proof.
We consider $d=1$ and $m=2$, and let $f\colon\mathbb{R}\to\mathbb{R}$ be defined through
$f(x)=\frac{1}{2}\left(f_{+}(x)+f_{-}(x)\right)=\frac{1}{2}\left(\frac{L}{2}(x+\delta)^{2}+\frac{L}{2}(x-\delta)^{2}\right).$
One can easily compute that
$f(x)=\frac{L\left(x^{2}+\delta^{2}\right)}{2},\quad\min f=\frac{L\delta^{2}}{2},\quad f(x)-\min f=\frac{Lx^{2}}{2},\quad x_{*}=0$
$\nabla f(x)=Lx,\quad\nabla f_{\pm}(x)=L(x\pm\delta),\quad\nabla f_{\pm}(x_{*})=\pm L\delta,$
We know that $\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}\leq 0$ for all $t=0,\ldots,T-1$ and for all $x_{0}\in\mathbb{R}$. We select $x_{0}$ sufficiently large such that $x_{t}$ is sufficiently large (to be defined) for all $t=0,\ldots,T-1$. Expanding the Lyapunov decrease and rearranging yields
$a_{t+1}\mathbb{E}_{t}\mathopen{}\left[{|x_{t+1}|^{2}}\right]\mathclose{}-a_{t}|x_{t}|^{2}+\rho\left(f(x_{t})-\min f\right)\leq e_{t}\sigma_{*}^{2}$
$\Longleftrightarrow$ $~{}a_{t+1}\left((1-\gamma L)^{2}|x_{t}|^{2}+(\gamma L)^{2}\delta^{2}\right)-a_{t}|x_{t}|^{2}+\rho L\left(\frac{|x_{t}|^{2}+\delta^{2}}{2}-\frac{\delta^{2}}{2}\right)\leq 2e_{t}\delta^{2}$
$\Longleftrightarrow$ $\left(a_{t+1}(1-\gamma L)^{2}-a_{t}+\frac{\rho L}{2}\right)|x_{t}|^{2}\leq\left(2e_{t}-a_{t+1}(\gamma L)^{2}\right)\delta^{2}.$
As $x_{t}$ is sufficiently large, and the right-hand side is constant, $a_{t+1}(1-\gamma L)^{2}-a_{t}+\frac{\rho L}{2}$ must be nonnegative, or equivalently
$a_{t+1}\leq\frac{1}{(1-\gamma L)^{2}}a_{t}-\frac{L}{2(1-\gamma L)^{2}}\rho.$
To simplify the notations, we note $\tau=(1-\gamma L)^{2}$, such that $a_{t+1}\leq\tau^{-1}a_{t}-\tau^{-1}\tfrac{\rho L}{2}$. By induction, we deduce that $a_{T}\leq\tau^{-T}a_{0}-\tfrac{\rho L}{2}\sum_{t=1}^{T}\tau^{-t}$. Because $a_{T}\geq 0$, we obtain that
$1=a_{0}\geq\frac{\rho L}{2}\sum\limits_{t=1}^{T}\tau^{T-t}=\frac{\rho L}{2}\sum\limits_{t=0}^{T-1}\tau^{T-t-1}\iff\rho\leq\frac{2a_{0}}{L\sum_{t=0}^{T-1}(1-\gamma L)^{2t}}=\frac{2\gamma(2-\gamma L)}{1-(1-\gamma L)^{2T}},$
concluding the proof. ∎
C.4 Detailed comparison to previous works
For comparative purposes, the results cited from previous works below have been adapted to align with our setting, such as by considering constant step-sizes where non-constant ones were originally used, or by adjusting the more general settings to match ours.
To the best of our knowledge, the analysis of SGD for convex functions with constant step-sizes $\gamma L\geq 1$ has not been studied previously. In the subsections below, we compare our Theorem 2.3, which addresses the case $\gamma L<1$, to the existing literature.
Comparison to gower_sgd_2021.
In (gower_sgd_2021, Theorem D.6), the authors analyzed the standard SGD algorithm under assumptions comparable to ours, while also accommodating for variable step-sizes. When adapted to our constant step-size setting, their result for $\gamma\in(0,\frac{1}{2L})$ boils down to
$\min_{k=0,\ldots,T-1}\mathbb{E}[f(x_{t})-\min f]\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma(1-2\gamma L)T}+\frac{\gamma}{1-2\gamma L}\sigma_{*}^{2}.$
One sees that this concurrent bound diverges when $\gamma$ gets close to $\tfrac{1}{2L}$, on both the bias and variance terms. This comparison must of course take into account the fact that the authors are able to provide bounds under weaker assumptions than our convexity and smoothness assumptions. They indeed assume $f$ to be smooth (instead of each $f_{i}$) and relax the functions $f_{i}$ to be quasar-convex, which is a consequence of our expected cocoercivity assumption, see Remark B.5.
Comparison to garrigos_handbook_2024.
Recently, in (garrigos_handbook_2024, Theorem 5.5), the finite-sum minimization problem for $L$-smooth convex functions under assumptions identical to ours was considered. For $\gamma\in(0,\frac{1}{4L}]$, they derived the bound
$\mathbb{E}[f(\overline{x}_{T})-\min f]\leq\frac{\|x_{0}-x_{*}\|^{2}}{\gamma T}+2\gamma\cdot\sigma_{*}^{2}.$
This result is encompassed as a consequence of Theorem 2.3.
Furthermore, their analysis can easily be extended. For any $\varepsilon>0$ and $\gamma\in(0,\frac{1}{(1+\varepsilon)L})$, it holds that
$\mathbb{E}[f(\overline{x}_{T})-\min f]\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T(1-(1+\varepsilon)\gamma L)}+\frac{(1+\varepsilon^{-1})\gamma}{2(1-(1+\varepsilon)\gamma L)}\cdot\sigma_{*}^{2}.$
This generalized bound is also recovered by Theorem 2.3 for any $\varepsilon>0$. In the interpolation setting, allowing $\varepsilon\to 0$, Theorem 2.3 remains strictly stronger.
Comparison to taylor_stochastic_2019.
In (taylor_stochastic_2019, Theorem 15), the authors investigate a primal averaging variant of SGD, originally with variable step-sizes. With constant step-sizes, their algorithm generates an auxiliary sequence $(y_{k})$ and follows the update rule
$y_{k+1}=\frac{k}{k+1}y_{k}+\frac{1}{k+1}x_{k},\quad x_{k+1}=x_{k}-\gamma\nabla f_{i_{k}}(y_{k+1}),$
where $i_{t}\in\mathcal{I}$ is sampled i.i.d. from the distribution $\mathcal{D}$.
For $\gamma\in(0,\frac{1}{L})$, they establish the bound
$\mathbb{E}[f(y_{T})-\min f]\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\frac{\gamma}{2(1-\gamma L)}\cdot\sigma_{*}^{2}.$
We observe that Theorem 2.3 provides a tighter bound, achieved without the use of the primal averaging scheme. It is worth noting that asymptotically, as $N\to\infty$, the two bounds are equivalent.
The authors also studied the interpolation setting, when $\sigma_{*}^{2}=0$, for their primal averaging scheme over a wider range of step-sizes (taylor_stochastic_2019, Theorem 8). For $\gamma\in(0,\tfrac{2}{L})$, they obtained
$\mathbb{E}\mathopen{}\left[{f(y_{T})-\min f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T(1-[\gamma L-1]_{+})},$
which is also weaker than Theorem 2.3, Corollary 2.6 and Theorem 2.7.
Regarding the primal averaging scheme, the comparison above suggests that for constant step-sizes in this setting, it does not yield faster convergence compared to standard SGD. This observation contrasts with the recent comparison of primal averaging schemes with acceleration schemes (defazio_momentum_2021).
We note that the authors also obtain rates on the vanilla SGD algorithm, where they however assume a uniformly bounded gradient variance $\sup_{x}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x)-\nabla f(x)\|^{2}}\right]\mathclose{}\leq\sigma^{2}$. Specifically, (taylor_stochastic_2019, Theorem 5) obtains a last iterate convergence rate for $\gamma L\in(0,1]$, given by
$\mathbb{E}\mathopen{}\left[{f(x_{T})-\min f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\gamma v_{T}\sigma^{2}\text{ with }v_{T}\simeq T.$
They additionally obtain in (taylor_stochastic_2019, Theorem 6) a bound on the Cesaro average for $\gamma L\in(0,\tfrac{1+\sqrt{5}}{2}]$, given by
$\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\gamma\frac{1+L}{2}\sigma^{2}.$
Appendix D Proofs in the smooth strongly convex setting
As in the previous section, we derive parameters that satisfy the sufficient conditions of Theorem B.4 in order to derive upper bounds for SGD, this time in the strongly convex setting. We again note that all propositions could simply be verified symbolically, yet we include the proofs to provide intuition on how these parameters were derived.
D.1 Bounds for any step-size
Because it seems impossible to obtain a bound with tight bias term for optimal step-sizes (see Sections F.3 and F.4), we provide here a more relaxed bound, where the bias term is $\varepsilon$ away from being tight, and subsequently the variance term remains bounded for every step-size $\gamma L\in(0,2)$.
Proposition D.1 (Lyapunov parameters. Strongly convex case, any step-sizes, $L>\mu$).
Let $\gamma L\in(0,2)$, with $L>\mu$. Let $\varepsilon\geq 0$ and $\gamma_{opt}=\tfrac{2}{\mu+L}$. Assume that either $\varepsilon>0$ or $\gamma\neq\gamma_{opt}$. Consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\rho=0$,
  • $a_{t}=\phi^{2(T-t)}$, where $\phi^{2}=\phi_{opt}^{2}+\varepsilon$, with
    $\phi_{opt}=\max\{1-\gamma\mu;\gamma L-1\}\quad\text{ and }\quad\varepsilon\in[0,1-\phi^{2}_{opt}),$
  • $\alpha_{t}=\beta_{t}=\alpha a_{t+1}$, where $\alpha=\omega+\omega_{\varepsilon}$ and
    $\omega=\frac{\gamma\phi_{opt}}{L-\mu}\quad\text{ and }\quad\omega_{\varepsilon}=\frac{\varepsilon+\sqrt{\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta}}{(L-\mu)^{2}}\quad\text{ with }\quad\delta=|\gamma-\gamma_{opt}|,$
  • $e_{t}=ea_{t+1}$, where
    $e=\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)}{2\omega_{\varepsilon}(L-\mu)+({L+\mu})\gamma\delta}\right).$
Then the sufficient Lyapunov conditions of Theorem B.4 are verified. Moreover, we have
$e_{T}^{sum}\coloneqq\sum_{t=0}^{T-1}e_{t}=e\frac{1-\phi^{2T}}{1-\phi^{2}}\leq\frac{e}{1-\phi^{2}},$
and
$e\leq\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)^{2}}{4\varepsilon+(L+\mu)(L-\mu)\gamma\delta}\right).$
Proof.
We start the proof by exploiting some insights from our numerical analysis. First, we observe that we can impose $\rho=0$ and $\alpha_{t}=\beta_{t}$ without changing anything to the empirical best bias term. We moreover infer that ${a_{t}}/{a_{t+1}}\in(0,1)$ is constant, so let us denote this ratio by $\phi^{2}$. Since $\phi_{\text{opt}}^{2}$ is the optimal rate for the deterministic gradient descent in the strongly convex smooth case, we anticipate that $\phi^{2}\geq\phi_{\text{opt}}^{2}$. Without loss of generality, we impose $a_{T}=1$, from which we deduce that $a_{t}=\phi^{2(T-t)}$. We were also able to deduce that $e_{t}=ea_{t+1}$ and that $\alpha_{t}=\alpha a_{t+1}$ for some constants $e,\alpha>0$, yet to be determined.
It remains to study the conditions of Theorem B.4 to find feasible values for $\alpha$ and $e$.
Condition 6 yields
$a_{t+1}\gamma^{2}(2\alpha a_{t+1})\leq(2\alpha a_{t+1}-a_{t+1}\gamma^{2})ea_{t+1},$
which holds true if
$2\gamma^{2}\alpha\leq(2\alpha-\gamma^{2})e.$
As we want to minimize $e$, we set
$e=\frac{2\gamma^{2}\alpha}{2\alpha-\gamma^{2}},$
and impose that the denominator $2\alpha-\gamma^{2}$ is positive. Now it remains to find $\alpha$. The above expression is decreasing with $\alpha$, so we aim to find the largest $\alpha$ possible such that all conditions are satisfied.
Turning to Condition 5 will determine $\alpha$. With our current notations and restrictions on the parameters, it can be written as
$(a_{t+1}\gamma-\alpha a_{t+1}(L+\mu))^{2}\leq(2\mu L\alpha a_{t+1}+a_{t}-a_{t+1})(2\alpha a_{t+1}-a_{t+1}\gamma^{2}),$
which is true if
$(\gamma-\alpha(L+\mu))^{2}-(2\mu L\alpha-(1-\phi^{2}))(2\alpha-\gamma^{2})\leq 0.$
This is a polynomial of degree at most $2$ in $\alpha$, which must be nonpositive. Expanding the terms to find the coefficients of this polynomial $P(\alpha)$, we obtain
$P(\alpha)$ $=$ $\gamma^{2}+\alpha^{2}(L+\mu)^{2}-2\alpha(L+\mu)\gamma-4\mu L\alpha^{2}+2\mu L\alpha\gamma^{2}+2\alpha(1-\phi^{2})-\gamma^{2}(1-\phi^{2})$
$=$ $\alpha^{2}\left[(L+\mu)^{2}-4\mu L\right]-\alpha\left[2\gamma(L+\mu)-2\mu L\gamma^{2}-2(1-\phi^{2})\right]+\left[\gamma^{2}-\gamma^{2}(1-\phi^{2})\right]$
$=$ $a\alpha^{2}-2b\alpha+c,$
where
$\begin{cases}a&=(L+\mu)^{2}-4\mu L=(L-\mu)^{2},\\ b&=\gamma(L+\mu)-\mu L\gamma^{2}-(1-\phi^{2}),\\ c&=\gamma^{2}\phi^{2}=\gamma^{2}(\phi_{opt}^{2}+\varepsilon).\end{cases}$ (7)
It is a simple exercise to verify that $\phi_{opt}=1-\mu\gamma$ if $\gamma\leq\gamma_{opt}$ and $\phi_{opt}=\gamma L-1$ if $\gamma\geq\gamma_{opt}$, which leads us to compute the value of $b$ through a case distinction:
  • if $\gamma\leq\gamma_{opt}$ then $1-\phi_{opt}^{2}=\gamma\mu(2-\gamma\mu)$ and
    $b$ $=$ $\gamma(L+\mu)-\mu L\gamma^{2}-\gamma\mu(2-\gamma\mu)+\varepsilon$
    $=$ $\gamma\left(L+\mu-\mu L\gamma-2\mu+\gamma\mu^{2}\right)+\varepsilon$
    $=$ $\gamma(1-\gamma\mu)(L-\mu)+\varepsilon$
    $=$ $\gamma\phi_{opt}(L-\mu)+\varepsilon.$
  • if $\gamma\geq\gamma_{opt}$ then $1-\phi_{opt}^{2}=\gamma L(2-\gamma L)$ and
    $b$ $=$ $\gamma(L+\mu)-\mu L\gamma^{2}-\gamma L(2-\gamma L)+\varepsilon$
    $=$ $\gamma\left(L+\mu-\mu L\gamma-2L+\gamma L^{2}\right)+\varepsilon$
    $=$ $\gamma(\gamma L-1)(L-\mu)+\varepsilon$
    $=$ $\gamma\phi_{opt}(L-\mu)+\varepsilon.$
In both cases we obtain $b=\gamma\phi_{opt}(L-\mu)+\varepsilon\geq 0$.
Since $L>\mu$, we have that $a>0$, so the largest feasible $\alpha$ is the largest root of $P$, if existent. The discriminant $\Delta$ of $P$ is given by
$\Delta$ $=$ $(2b)^{2}-4ac$
$=$ $4\left(\gamma\phi_{opt}(L-\mu)+\varepsilon\right)^{2}-4(L-\mu)^{2}\gamma^{2}(\phi^{2}_{opt}+\varepsilon)$
$=$ $4\gamma^{2}\phi_{opt}^{2}(L-\mu)^{2}+4\varepsilon^{2}+8\varepsilon\gamma\phi_{opt}(L-\mu)-4(L-\mu)^{2}\gamma^{2}\phi^{2}_{opt}-4(L-\mu)^{2}\gamma^{2}\varepsilon$
$=$ $4\left[\varepsilon^{2}+\varepsilon\left(2\gamma\phi_{opt}(L-\mu)-(L-\mu)^{2}\gamma^{2}\right)\right]$
$=$ $4\left[\varepsilon^{2}+\varepsilon\gamma(L-\mu)\left(2\phi_{opt}-(L-\mu)\gamma\right)\right].$
A simple calculation shows that
$2\phi_{opt}-(L-\mu)\gamma=\begin{cases}2-\gamma(L+\mu)&\text{ if }\gamma\leq\gamma_{opt}\\ -2+\gamma(L+\mu)&\text{ if }\gamma\geq\gamma_{opt}\end{cases}=|2-\gamma(L+\mu)|=(L+\mu)|\gamma-\gamma_{opt}|,$
such that, with $\delta=|\gamma-\gamma_{opt}|$,
$\Delta=4\left[\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta\right]\geq 0.$
As such, $P$ has real roots. We thus define
$\alpha$ $=$ $\frac{2b+\sqrt{\Delta}}{2a}$
$=$ $\frac{\gamma\phi_{opt}(L-\mu)+\varepsilon+\sqrt{\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta}}{(L-\mu)^{2}}$
$=$ $\omega+\omega_{\varepsilon},$
where
$\omega=\frac{\gamma\phi_{opt}}{L-\mu}\quad\text{ and }\quad\omega_{\varepsilon}=\frac{\varepsilon+\sqrt{\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta}}{(L-\mu)^{2}}.$
To conclude the proof, we are now going to verify that this choice of $\alpha$ is feasible, in the sense that it satisfies all the conditions from Theorem B.4. The nonnegativity Condition 1 is satisfied by construction. Condition 2 is satisfied as we set $\rho=0$ and $\alpha_{t}=\beta_{t}$. Condition 3 is equivalent to
$1\leq 2\alpha\mu L+\frac{a_{t}}{a_{t+1}}=2\alpha\mu L+\phi^{2}.$
We have $\alpha=\omega+\omega_{\varepsilon}\geq\omega$ and $\phi^{2}\geq\phi^{2}_{opt}$, and the above is thus true if
$1\leq\mu L2\omega+\phi^{2}_{opt}$
$\iff$ $1\leq\frac{2\mu L\gamma\phi_{opt}}{L-\mu}+\phi^{2}_{opt}$
$\iff$ $0\leq 2\mu L\gamma\phi_{opt}-(1-\phi^{2}_{opt})(L-\mu).$
To assess whether this inequality is true, we are going to consider two cases:
  • if $\gamma\leq\gamma_{opt}$, then $\phi_{opt}=1-\gamma\mu$, so the inequality becomes
    00 0 $\leq$ $2\mu L\gamma(1-\gamma\mu)-\gamma\mu(2-\gamma\mu)(L-\mu)$
    $=$ $\mu\gamma(-\mu\gamma L-\gamma\mu^{2}+2\mu)$
    $=$ $\mu^{2}\gamma(2-\gamma(L+\mu))$
    $=$ $\mu^{2}\gamma(\mu+L)\delta.$
  • if $\gamma\geq\gamma_{opt}$, then $\phi_{opt}=\gamma L-1$, so the inequality becomes
    00 0 $\leq$ $2\mu L\gamma(\gamma L-1)-\gamma L(2-\gamma L)(L-\mu)$
    $=$ $L\gamma(\mu\gamma L+\gamma L^{2}-2L)$
    $=$ $-L^{2}\gamma(2-\gamma(L+\mu))$
    $=$ $L^{2}\gamma(\mu+L)\delta.$
In both cases the inequality is verified, therefore Condition 3 holds true. Let us now turn to Condition 4, which, in our context, is equivalent to
$0\leq 2\alpha-\gamma^{2},$
which holds true as
$2\alpha-\gamma^{2}$ $=$ $2\omega_{\varepsilon}+2\omega-\gamma^{2}=2\omega_{\varepsilon}+\frac{\gamma}{L-\mu}\left(2\phi_{opt}-\gamma(L-\mu)\right)$
$=$ $2\omega_{\varepsilon}+\frac{\gamma}{L-\mu}(L+\mu)\delta.$
Condition 5 is true by the choice of $\alpha$. Finally, Condition 6 holds true provided $02\alpha-\gamma^{2}>$ 0, which is verified through the above computation provided $\varepsilon>0$ or $\delta\neq 0$. In this case,
$e$ $=$ $\gamma^{2}\frac{2\omega+2\omega_{\varepsilon}}{2\omega_{\varepsilon}+\frac{L+\mu}{L-\mu}\gamma\delta}$
$=$ $\gamma^{2}\frac{2\gamma\phi_{opt}+2(L-\mu)\omega_{\varepsilon}}{2\omega_{\varepsilon}(L-\mu)+({L+\mu})\gamma\delta}$
$=$ $\gamma^{2}\frac{\gamma((L+\mu)\delta+\gamma(L-\mu))+2(L-\mu)\omega_{\varepsilon}}{2\omega_{\varepsilon}(L-\mu)+({L+\mu})\gamma\delta}$
$=$ $\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)}{2\omega_{\varepsilon}(L-\mu)+({L+\mu})\gamma\delta}\right).$
The bound on $e_{T}^{sum}$ is a direct consequence of the fact that $a_{t}$ is a geometric sequence. The upper bound for $e$ comes from the lower bound $\omega_{\varepsilon}\geq\frac{2\varepsilon}{(L-\mu)^{2}}$. ∎
We note the above proposition only holds true for $\mu<L$, thus not covering the case $L=\mu$. We cover this case in a separate proposition.
Proposition D.2 (Lyapunov parameters. Strongly convex case, any step-sizes, $L=\mu$).
Let $\gamma L\in(0,2)$, with $L=\mu$. Let $\varepsilon>0$. Consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\rho=0$,
  • $a_{t}=\phi^{2(T-t)}$, where $\phi^{2}=\phi_{opt}^{2}+\varepsilon$, with
    $\phi_{opt}=\max\{1-\gamma\mu;\gamma L-1\}\quad\text{ and }\quad\varepsilon\in[0,1-\phi^{2}_{opt}),$
  • $\alpha_{t}=\beta_{t}$ and $\alpha_{t}=\alpha a_{t+1}$ where
    $\alpha\geq\max\left\{\frac{\gamma^{2}}{2}\left(1+\frac{\phi_{opt}^{2}}{\varepsilon}\right),\frac{1-\phi^{2}_{opt}-\varepsilon}{2L^{2}}\right\},$
  • $e_{t}=ea_{t+1}$ where
    $e=\frac{2\gamma^{2}\alpha}{2\alpha-\gamma^{2}}.$
Then the sufficient Lyapunov conditions of Theorem B.4. Moreover, we have
$\phi\xrightarrow{\varepsilon\to 0}\phi_{opt},\quad\alpha\xrightarrow{\varepsilon\to 0}+\infty,\quad e\xrightarrow{\varepsilon\to 0}\gamma^{2}.$
Proof.
The arguments in this proof are mostly recycled from Proposition D.1. In particular, we shall use the same values for $\rho$ and $a_{t}$. Regarding the value of $\alpha$, we impose $a\alpha^{2}-2b\alpha+c\leq 0$ where $a,b,c$ are defined in Equation (7), which we recall for convenience:
$\begin{cases}a&=(L-\mu)^{2},\\ b&=\gamma(L+\mu)-\mu L\gamma^{2}-(1-\phi^{2}),\\ c&=\gamma^{2}(\phi_{opt}^{2}+\varepsilon).\end{cases}$
Because $L=\mu$, it holds that $a=0$, and that
$b=2\gamma L-L^{2}\gamma^{2}-(1-\phi_{opt}^{2})+\varepsilon=2\gamma L-L^{2}\gamma^{2}-\gamma L(2-\gamma L)+\varepsilon=\varepsilon.$
Therefore, $\alpha$ must satisfy $c\leq 2b\alpha$, that is
$\alpha\geq\frac{c}{2b}=\frac{\gamma^{2}(\phi_{opt}^{2}+\varepsilon)}{2\varepsilon}=\frac{\gamma^{2}}{2}\left(1+\frac{\phi_{opt}^{2}}{\varepsilon}\right).$ (8)
It then remains to verify the conditions of Theorem B.4. Conditions 1 and 2 are readily satisfied. Condition 3 can be rewritten as
$\alpha\geq\frac{1-\phi^{2}_{opt}-\varepsilon}{2L^{2}},$
thus justified the bound on $\alpha$. Condition 4 requires that $\alpha\geq\tfrac{\gamma^{2}}{2}$, which is already true because of Equation (8). Condition 5 is satisfied under Equation (8). Finally, Condition 6 is satisfied with
$e=\frac{2\gamma^{2}\alpha}{2\alpha-\gamma^{2}},$
where the denominator is strictly positive because of (8). ∎
Theorem D.3 (Bound for SGD. Strongly convex case, general step-sizes).
Let Assumptions A.5 and A.7 hold, with $\mu>0$. Let $\varepsilon\geq 0$ and $\gamma_{opt}=\tfrac{2}{\mu+L}$. Let $x_{t}$ be generated by SGD, with $\gamma L\in(0,2)$. Assume that either $\varepsilon>0$ or $\gamma\neq\gamma_{opt}$. Then, for every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\phi^{2T}\cdot\|x_{0}-x_{*}\|^{2}+\frac{1-\phi^{2T}}{1-\phi^{2}}e\sigma_{*}^{2},$
where $\phi^{2}=\varepsilon+(\max\{1-\gamma\mu;\gamma L-1\})^{2}\in[0,1)$, and $e$ is defined in Proposition D.1 and verifies
$e\leq\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)^{2}}{4\varepsilon+(L+\mu)(L-\mu)\gamma|\gamma-\gamma_{opt}|}\right).$
Proof.
We separate the proof into two cases, namely the case $L>\mu$ and the case $L=\mu$.
  • Consider $L>\mu$. Proposition D.1, combined with Theorem B.4, provides parameters ensuring that the Lyapunov energy decreases with, in particular, $a_{T}=1$, $a_{0}=\phi^{2T}$. We deduce the wanted bound from Lemma B.3.
  • Consider $L=\mu$. Let us consider first $\varepsilon>0$. Combining Lemma B.3, Theorem B.4 and Proposition D.2 we obtain
    $\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\phi^{2T}\cdot\|x_{0}-x_{*}\|^{2}+\frac{1-\phi^{2T}}{1-\phi^{2}}e_{\varepsilon}\sigma_{*}^{2},$
    where $e_{\varepsilon}$ converges to $\gamma^{2}$ and $\phi\to\phi_{\text{opt}}$ when $\varepsilon\to 0$, thus yielding the wanted bound when $\varepsilon=0$. If we wish to prove the claim for $\varepsilon>0$, it is enough to realize that $\phi_{opt}\leq\phi$, together with the fact that $\tfrac{1-\phi^{2T}}{1-\phi^{2}}$ is nondecreasing with $\phi$.
The above propositions allow us to formulate bounds for SGD. As in the convex case, we consider optimal and non-optimal step-sizes as separate cases.
Theorem D.4 (Tight bound for SGD. Strongly convex case, non-optimal step-sizes).
Let Assumptions A.5 and A.7 hold, with $\mu>0$. Let $x_{t}$ be generated by SGD, with $\gamma L\in(0,2)$ such that $\gamma\neq\gamma_{opt}=\tfrac{2}{\mu+L}$. Then, for every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\phi^{2T}\cdot\|x_{0}-x_{*}\|^{2}+\frac{1-\phi^{2T}}{1-\phi^{2}}e\sigma_{*}^{2},$
where $\phi=\max\{1-\gamma\mu;\gamma L-1\}\in[0,1)$, and
$e=\frac{2\gamma^{2}\phi}{|2-\gamma(L+\mu)|}=\gamma^{2}\left(1+\frac{\gamma(L-\mu)}{|\gamma-\gamma_{opt}|(L+\mu)}\right).$
Proof.
This is an immediate consequence of Theorem D.3 with $\varepsilon=0$. ∎
If $\gamma L\leq 1$, the above result can be simplified, allowing for an easier comparison with the literature.
Corollary D.5 (Tight bound for SGD. Strongly convex case, short step-size $\gamma L\leq 1$).
Let Assumptions A.5 and A.7 hold, with $\mu>0$. Let $x_{t}$ be generated by SGD, with $\gamma L\in(0,1]$. Then, for every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq(1-\gamma\mu)^{2T}\cdot\|x_{0}-x_{*}\|^{2}+\frac{2\gamma}{\mu(2-\frac{\mu}{L})}\,\sigma_{*}^{2}.$
Proof.
If $\gamma L\leq 1$, then $\phi=1-\gamma\mu$. Moreover,
$\frac{1-\phi^{2T}}{1-\phi^{2}}e\leq\frac{e}{1-\phi^{2}}=\frac{2\gamma^{2}(1-\gamma\mu)}{\gamma\mu(2-\gamma\mu)\big{(}2-\gamma(L+\mu)\big{)}}=\frac{2\gamma}{\mu}\left[\frac{(1-\gamma\mu)}{(2-\gamma\mu)\big{(}2-\gamma(L+\mu)\big{)}}\right].$
The rational function between the large brackets increases with $\gamma$ on $(0,\frac{1}{L}]$ (on $(0,\frac{2}{L+\mu})$, actually), so it can be bounded from above there by its value on $\gamma=\frac{1}{L}$, which is $\frac{1}{2-\frac{\mu}{L}}$. ∎
Finally, we provide a bound which remains valid for every steps-size, with the counterpart that the bias term gets sub-optimal.
Theorem D.6 (Subtight bound for SGD. Strongly convex case, general step-sizes).
Let Assumptions A.5 and A.7 hold, with $\mu>0$. Let $\varepsilon>0$, and let $x_{t}$ be generated by SGD, with $\gamma L\in(0,2)$. Then, for every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\phi^{2T}\cdot\|x_{0}-x_{*}\|^{2}+\frac{1-\phi^{2T}}{1-\phi^{2}}e\sigma_{*}^{2},$
where $\phi^{2}=\varepsilon+(\max\{1-\gamma\mu;\gamma L-1\})^{2}\in[0,1)$, and $e$ was defined in Proposition D.1 and verifies
$e\leq\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)^{2}}{4\varepsilon}\right).$
Proof.
This is an immediate consequence of Theorem D.3, where we used $\varepsilon>0$ to lower bound the denominator of $e$. ∎
D.2 Detailed comparison to previous works
Comparison to bach_non-asymptotic_2011.
In (bach_non-asymptotic_2011, Theorem 1), the authors analyze the convergence of SGD for strongly convex objectives without variance assumptions, aligning with the setting of our Theorem 3.3. They present the following bound:
$\mathbb{E}\mathopen{}\left[{\|x_{N}-x_{*}\|^{2}}\right]\mathclose{}\leq 2\exp(4L^{2}\gamma^{2}(N-1))\cdot\exp(-\mu\gamma N/4)\cdot(\|x_{0}-x_{*}\|^{2}+\sigma_{*}^{2}/L^{2})+4\gamma\sigma_{*}^{2}/\mu.$
Their proof is based on a Lyapunov energy function similar to the one employed in our analysis, though with different coefficients. The exponential form is presented for clarity in the context of their use of decreasing step-sizes. For the constant step-size scenario, their analysis yields
$\mathbb{E}\mathopen{}\left[{\|x_{N}-x_{*}\|^{2}}\right]\mathclose{}\leq\varphi^{N}\cdot\|x_{0}-x_{*}\|^{2}+2\gamma^{2}\cdot\frac{1-\varphi^{N}}{1-\varphi}\cdot\sigma_{*}^{2},$
where $\varphi=1-2\mu\gamma+2L^{2}\gamma^{2}$. It may be verified that $\phi^{2}\leq\varphi$, where $\phi$ is defined as in Theorem D.4, which connects their result to ours. In specific, our analysis explicitly establishes convergence over a broader range of step-sizes $\gamma$ compared to the range considered in their work, and provides a sharper rate.
Comparison to needell_stochastic_2016.
In (needell_stochastic_2016, Theorem 2.1), the authors obtain, under the same assumptions and for $\gamma<\frac{1}{L}$, the bound
$\mathbb{E}\mathopen{}\left[{\|x_{k}-x_{*}\|^{2}}\right]\mathclose{}\leq\left(1-2\gamma\mu(1-\gamma L)\right)^{k}\cdot\|x_{0}-x_{*}\|^{2}+\frac{\gamma}{\mu(1-\gamma L)}\cdot\sigma_{*}^{2}.$
We note that the bound is weaker than the one obtained in our results, and is restricted to short stepsizes.
Comparison to gower_sgd_2019.
The authors of (gower_sgd_2019, Theorem 3.1) study the quasi-strongly convex setting under the assumption of expected smoothness, which is tightly linked to our assumption (EC), as put forward in Remark A.6
For a constant step-size $\gamma\in(0,1/(2L)]$, they derive the bound
$\mathbb{E}\mathopen{}\left[{\|x_{N}-x_{*}\|^{2}}\right]\mathclose{}\leq(1-\gamma\mu)^{k}\|x_{0}-x_{*}\|^{2}+\frac{2\gamma}{\mu}\cdot\sigma_{*}^{2}.$
Our analysis extends the convergence result to a broader range of step-sizes. Furthermore, within this specific range of step-sizes, our derived convergence rate incorporates a squared term related to $(1-\gamma\mu)$, differing from the linear term presented in their bound. This is an effect often seen in simpler proofs.
Appendix E Performance Estimation Problem methodology
In this section, we describe the Performance Estimation Problem methodology in more detail, and explain how we made use of it in this scenario. Most ideas are inspired by taylor_stochastic_2019.
Throughout this section, we define $\mathcal{F}_{\mu,L}(\mathbb{R}^{d})$ to be the set of $L$-smooth $\mu$-strongly convex functions. We shall often drop the dependency in $\mathbb{R}^{d}$ when the space is clear from context.
E.1 Problem reformulation
As we aim to obtain a computationally tractable problem, we do restrict ourselves to the finite-sum setting in this section. We do however note that the number of functions $m$ does not influence the final result, and can hence theoretically be taken to be infinite, thus recovering the results of Sections C and D.
As explained by Lemmas B.2 and B.3, our goal is to derive nonnegative parameters $\rho,(a_{t},e_{t})$ such that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for all $t=0,\ldots,T-1$, where
$E_{t}=a_{t}\|x_{t}-x_{*}\|^{2}+\rho\cdot\sum_{s=0}^{t-1}[f(x_{s})-\min f]-\sum_{s=0}^{t-1}e_{s}\cdot\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}\quad\text{for $t=0,\ldots,T-1$},$
and $x_{*}\in\operatorname{argmin}f$. We replaced $\sigma_{*}^{2}$ by its equivalent value $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$ for simplicity, to avoid the additional variable $\sigma_{*}^{2}$.
We call parameters $\rho,(a_{t}),(e_{t})$ Lyapunov parameters if $\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}\leq 0$ holds for all $t=0,\ldots,T-1$, for all $d\geq 1$, for all functions $f_{1},\dots,f_{m}\in\mathcal{F}_{\mu,L}(\mathbb{R}^{d})$, all $x_{0}\in\mathbb{R}^{d}$, and all $x_{*}\in\operatorname{argmin}f$. Note that being Lyapunov parameters depends on the number of functions $m$ and the time horizon $T$. We denote the set of all Lyapunov parameters by $\mathcal{V}_{T}$, where we omit the dependence on $m$, as our results happen to be independent of $m$.
To unify the presentation for the convex and the strongly convex setting, we introduce some notation.
  • In the convex setting, Lemma B.2 tells us that we are interested in maximizing $\rho$, or equivalently minimizing $\tfrac{1}{\rho}$, where we fix $a_{0}=1$. In this setting we denote the bias and normalization by
    $\text{Bias}=\frac{1}{\rho}\quad\text{and}\quad\mathcal{N}_{T}=\{\rho,(a_{t},e_{t})\in\mathcal{V}_{T}\colon a_{0}=1\}.$
  • In the strongly convex setting, Lemma B.3 tells us we are interested in minimizing $a_{0}$, where we fix $a_{T}=1$. In this setting we denote the bias and normalization by
    $\text{Bias}=a_{0}\quad\text{and}\quad\mathcal{N}_{T}=\{\rho,(a_{t},e_{t})\in\mathcal{V}_{T}\colon a_{T}=1\}.$
With this notation in mind, we may formally state our problem as
$_{\text{opt}}=\text{Bias}_{\text{opt}}(T)\coloneq\inf_{\rho,(a_{t},e_{t})\in\mathcal{N}_{T}}\left\{\text{Bias}\right\}.$ Bias } .
The definition of being a Lyapunov parameter is such that it generates a decrease in energy at each step. Specifically, it holds true that
$\text{Bias}_{\text{opt}}$ $=\inf_{\mathcal{N}_{T}}\left\{\text{Bias}\colon\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}\leq 0~{}\forall t=0,\ldots,T-1,\forall x_{0}\in\mathbb{R}^{d},\forall f_{i}\in\mathcal{F}_{\mu,L},\forall d\in\mathbb{Z}_{\geq 1}\right\}.$
By permuting the quantifiers, this may then be reformulated as a bilevel program, namely:
$\text{Bias}_{\text{opt}}$ $=\inf_{\mathcal{N}_{T}}\left\{\text{Bias}\colon B_{t}\leq 0\quad\forall t=0,\ldots,T-1\right\},$
where
$B_{t}=B_{t}(\rho,(a_{t},e_{t}))\coloneq\quad\quad\sup_{d\in\mathbb{Z}_{\geq 1}}\quad$ $\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}$
[…]
$\nabla f(x_{*})=0,\quad x_{0},x_{*}\in\mathbb{R}^{d},\quad f_{i}\in\mathcal{F}_{\mu,L}.$
Note that $x_{0}$ is a variable of the follower problem, and hence acts as a slack variable for the constraint that $x_{t}$ is generated from $x_{0}$. Using this, and recalling the definition of $E_{t}$, we may write that
$B_{t}=\quad\quad\sup_{d\in\mathbb{Z}_{\geq 1}}\quad$ $a_{t+1}\mathbb{E}\mathopen{}\left[{\|x_{t+1}-x_{*}\|^{2}}\right]\mathclose{}-a_{t}\|x_{t}-x_{*}\|^{2}+\rho(f(x_{t})-f(x_{*}))-e_{t}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$
subject to $\sum_{i=1}^{m}\nabla f_{i}(x_{*})=0,\quad x_{t},x_{*}\in\mathbb{R}^{d},\quad f_{i}\in\mathcal{F}_{\mu,L}.$
By introducing the notation $f_{j}^{(i)}=f_{i}(x_{j})$ and $g_{j}^{(i)}=\nabla f_{i}(x_{j})$ for $j\in\{t,*\}$ and $i=1,\ldots,m$, and by rewriting the expectations as finite sums, we may rewrite the problem equivalently as
$B_{t}=\quad\quad\sup_{d\in\mathbb{Z}_{\geq 1}}\quad$ $\frac{a_{t+1}}{m}\sum_{i=1}^{m}{\|x_{t}-\gamma g_{t}^{(i)}-x_{*}\|^{2}}-a_{t}\|x_{t}-x_{*}\|^{2}+\frac{\rho}{m}\sum_{i=1}^{m}(f_{t}^{(i)}-f_{*}^{(i)})-\frac{e_{t}}{m}\sum_{i=1}^{m}\|g_{*}^{(i)}\|^{2}$
subject to $\sum_{i=1}^{m}g_{*}^{(i)}=0,$
there exist $f_{i}\in\mathcal{F}_{\mu,L}$ such that $f_{i}(x_{j})=f_{j}^{(i)}$ and $\nabla f_{i}(x_{j})=g_{j}^{(i)}$
for j{t,} and $\qquad\qquad\qquad\qquad\qquad\qquad\text{for $j\in\{t,*\}$ and $i=1,\ldots,m$},$
$x_{t},x_{*}\in\mathbb{R}^{d},\quad g_{t}^{(i)},g_{*}^{(i)}\in\mathbb{R}^{d},\quad f_{t}^{(i)},f_{*}^{(i)}\in\mathbb{R}\quad\text{for all }i=1,\ldots,m.$
The existence condition may seem like a complicated reformulation, but it allows us to use (taylor_smooth_2017, Theorem 4), a cornerstone of the performance estimation framework:
Theorem E.1.
Let $L\geq\mu\geq 0$ and let $\left\{(x_{i},g_{i},f_{i})_{i\in\mathcal{I}}\right\}\subset\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}$ be a finite set of triplets. There exists a function $f\in\mathcal{F}_{\mu,L}$ such that
$f(x_{i})=f_{i}\quad\text{and}\quad\nabla f(x_{i})=g_{i}\quad\text{for all $i\in\mathcal{I}$},$
if, and only if, for every pair of indices $(i,j)\in\mathcal{I}^{2}$, it holds that
$f_{i}-f_{j}-\langle g_{j},x_{i}-x_{j}\rangle\geq\frac{1}{2(1-\frac{\mu}{L})}\left(\frac{1}{L}\|g_{i}-g_{j}\|^{2}+\mu\|x_{i}-x_{j}\|^{2}-2\frac{\mu}{L}\langle g_{i}-g_{j},x_{i}-x_{j}\rangle\right).$ (9)
Using this theorem, we may rewrite our program equivalently as
$B_{t}=\quad\quad\sup_{d\in\mathbb{Z}_{\geq 1}}\quad$ $\frac{a_{t+1}}{m}\sum_{i=1}^{m}{\|x_{t}-\gamma g_{t}^{(i)}-x_{*}\|^{2}}-a_{t}\|x_{t}-x_{*}\|^{2}+\frac{\rho}{m}\sum_{i=1}^{m}(f_{t}^{(i)}-f_{*}^{(i)})-\frac{e_{t}}{m}\sum_{i=1}^{m}\|g_{*}^{(i)}\|^{2}$
subject to $\left\|\sum_{i=1}^{m}g_{*}^{(i)}\right\|^{2}=0,$
[…]
$x_{t},x_{*}\in\mathbb{R}^{d},\quad g_{t}^{(i)},g_{*}^{(i)}\in\mathbb{R}^{d},\quad f_{t}^{(i)},f_{*}^{(i)}\in\mathbb{R}\quad\text{for all }i=1,\ldots,m.$
The variable $g_{*}^{(m)}$ is redundant, as it is given by $g_{*}^{(m)}=-\sum_{i=1}^{m-1}g_{*}^{(i)}$. To simplify notation, we introduce
$P_{t}$ $=\left(g_{t}^{(1)},\ldots,g_{t}^{(m)},g_{*}^{(1)},\ldots,g_{*}^{(m-1)},x_{t}-x_{*}\right)\in\mathbb{R}^{d\times 2m},$ (10a)
$F_{t}$ $=\left(f_{t}^{(1)},\ldots,f_{t}^{(m)},f_{*}^{(1)},\ldots,f_{*}^{(m)}\right)\in\mathbb{R}^{2m},$ (10b)
and define the vectors $\bm{p}_{i}\in\mathbb{R}^{2m}$ for $i=1,\ldots,2m+1$, such that
$\begin{cases}P_{t}\bm{p}_{i}=g_{t}^{(i)}&\text{for $i=1,\ldots,m$},\\ P_{t}\bm{p}_{m+i}=g_{*}^{(i)}&\text{for $i=1,\ldots,m-1$},\\ P_{t}\bm{p}_{2m}=-\sum_{i=1}^{m-1}g_{*}^{(i)},&\\ P_{t}\bm{p}_{2m+1}=x_{t}-x_{*},\end{cases}$
and $\bm{f}_{i}\in\mathbb{R}^{2m}$ for $i=1,\ldots,2m$ such that, for $i=1,\ldots,m$,
$F_{t}\bm{f}_{i}=f_{t}^{(i)},\quad F_{t}\bm{f}_{m+i}=f_{*}^{(i)}.$
With this, we may write our program as
$B_{t}=\quad\quad\quad\sup_{d\in\mathbb{Z}_{\geq 1}}$ $P_{t}^{T}\Delta_{t}P_{t}+F_{t}\tilde{\Delta}$
subject to $F_{t}\bm{f}_{m+i}-F_{t}\bm{f}_{i}+P_{t}^{T}A_{t,*}^{(i)}P_{t}\leq 0\text{ for all $i=1,\ldots,m$},$
$F_{t}\bm{f}_{i}-F_{t}\bm{f}_{m+i}+P_{t}^{T}A_{*,t}^{(i)}P_{t}\leq 0\text{ for all $i=1,\ldots,m$},$
$P_{t}\in\mathbb{R}^{d\times 2m},\quad F_{t}\in\mathbb{R}^{2m},$
where
$\begin{dcases}\Delta_{t}&=\frac{a_{t+1}}{m}\sum_{i=1}^{m}(\bm{p}_{2m+1}-\gamma\bm{p}_{m+i})(\bm{p}_{2m+1}-\gamma\bm{p}_{m+i})^{T}-a_{t}\bm{p}_{2m+1}\bm{p}_{2m+1}^{T}\\ &\qquad\qquad\qquad-\frac{e_{t}}{m}\sum_{i=1}^{m}\bm{p}_{m+i}\bm{p}_{m+i}^{T},\\ \tilde{\Delta}&=\frac{\rho}{m}\sum_{i=1}^{m}\bm{f}_{i}-\bm{f}_{m+i},\\ A_{t,*}^{(i)}&=\bm{p}_{m+i}\bm{p}_{2m+1}^{T}+\frac{1}{2(L-\mu)}(\bm{p}_{i}-\bm{p}_{m+i})(\bm{p}_{i}-\bm{p}_{m+i})^{T}+\frac{\mu L}{2(L-\mu)}\bm{p}_{2m+1}\bm{p}_{2m+1}^{T}\\ &\qquad\qquad\qquad-\frac{\mu}{L-\mu}(\bm{p}_{i}-\bm{p}_{m+i})\bm{p}_{2m+1}^{T},\\ A_{*,t}^{(i)}&=-\bm{p}_{i}\bm{p}_{2m+1}^{T}+\frac{1}{2(L-\mu)}(\bm{p}_{i}-\bm{p}_{m+i})(\bm{p}_{i}-\bm{p}_{m+i})^{T}+\frac{\mu L}{2(L-\mu)}\bm{p}_{2m+1}\bm{p}_{2m+1}^{T}\\ &\qquad\qquad\qquad-\frac{\mu}{L-\mu}(\bm{p}_{i}-\bm{p}_{m+i})\bm{p}_{2m+1}^{T}.\\ \end{dcases}$
This formulation is a non-convex quadratic program. A standard technique to render this solvable is to cast it into a semi-definite program. We do so by introducing $G_{t}=P_{t}^{T}P_{t}\succeq 0$, and obtain
$B_{t}=\sup_{d\in\mathbb{Z}_{\geq 1}}\quad\quad\quad\sup_{G_{t}\in\mathcal{S}^{2m},F_{t}\in\mathbb{R}^{2m}}\quad$ $\operatorname{Tr}(\Delta_{t}G_{t})+F_{t}\tilde{\Delta}$
subject to $F_{t}\bm{f}_{m+i}-F_{t}\bm{f}_{i}+\operatorname{Tr}(A_{t,*}^{(i)}G_{t})\leq 0\text{ for all $i=1,\ldots,m$}$
$F_{t}\bm{f}_{i}-F_{t}\bm{f}_{m+i}+\operatorname{Tr}(A_{*,t}^{(i)}G_{t})\leq 0\text{ for all $i=1,\ldots,m$}$
$\text{rank}(G_{t})\leq d,$
where $\mathcal{S}^{k}$ denotes the set of symmetric positive definite matrices of dimension $k\times k$. This formulation is equivalent to the previous one due to the rank condition: given any matrix $G_{t}\in\mathcal{S}^{k}$ of rank lesser than $d$, a Cholesky factorization $G_{t}=P_{t}^{T}P_{t}$ would yield a $P_{t}\in\mathbb{R}^{d\times 2m}$, which is feasible to the previous formulation. Now, observe that the non-convex non-continuous rank constraint may be removed due to the exterior supremum over $d\geq 1$. Once this is done, we may as well remove the supremum over $d$ because it does not appear anymore in the inner supremum. This allows us to write
$B_{t}=\quad\quad\quad\sup_{G_{t}\in\mathcal{S}^{2m},F_{t}\in\mathbb{R}^{2m}}\quad$ $\operatorname{Tr}(\Delta_{t}G_{t})+F_{t}\tilde{\Delta}$ (Primal)
subject to $F_{t}\bm{f}_{m+i}-F_{t}\bm{f}_{i}+\operatorname{Tr}(A_{t,*}^{(i)}G_{t})\leq 0\text{ for all $i=1,\ldots,m$},$
$F_{t}\bm{f}_{i}-F_{t}\bm{f}_{m+i}+\operatorname{Tr}(A_{*,t}^{(i)}G_{t})\leq 0\text{ for all $i=1,\ldots,m$}.$
We have now expressed $B_{t}$ as the optimal value of a semi-definite program. It is a simple exercise to compute the dual of this problem, which happens to be a feasibility problem:
$\tilde{B}_{t}=\quad\quad\quad\sup_{\Lambda_{t},\lambda_{t,*}^{(i)},\lambda_{*,t}}\quad$ 00 0 (Dual)
subject to $-\Delta_{t}+\sum_{i=1}^{m}\left(\lambda_{t,*}^{(i)}A_{t,*}^{(i)}+\lambda_{*,t}^{(i)}A_{*,t}^{(i)}\right)=\Lambda_{t},$
$-\tilde{\Delta}+\sum_{i=1}^{m}\lambda_{t,*}^{(i)}(\bm{f}_{m+i}-\bm{f}_{i})+\sum_{i=1}^{m}\lambda_{*,t}^{(i)}(\bm{f}_{i}-\bm{f}_{m+i})=0,$
$\Lambda_{t}\in\mathcal{S}^{2m},\quad\lambda_{t,*}^{(i)},\lambda_{*,t}^{(i)}\in\mathbb{R}_{\geq 0}\quad\text{for all $i=1,\ldots,m$}.$
We prove in Section E.3 that this is indeed the dual problem of Problem (Primal). We further prove that strong duality holds, meaning that $\tilde{B}_{t}=-B_{t}$. As a consequence,
[…]
where the last equivalence follows from the structure of Problem (Dual). In conclusion, we can rewrite the problem of minimizing Bias over the Lyapunov parameters into
$\text{Bias}_{\text{opt}}=\quad\quad\inf_{\mathcal{N}_{T}}\quad$ Bias
subject to $-\Delta_{t}+\sum_{i=1}^{m}\left(\lambda_{t,*}^{(i)}A_{t,*}^{(i)}+\lambda_{*,t}^{(i)}A_{*,t}^{(i)}\right)=\Lambda_{t},$
$-\tilde{\Delta}+\sum_{i=1}^{m}\lambda_{t,*}^{(i)}(\bm{f}_{m+i}-\bm{f}_{i})+\sum_{i=1}^{m}\lambda_{*,t}^{(i)}(\bm{f}_{i}-\bm{f}_{m+i})=0,$
$\Lambda_{t}\in\mathcal{S}^{2m},\quad\lambda_{t,*}^{(i)},\lambda_{*,t}^{(i)}\in\mathbb{R}_{\geq 0}\quad\text{for all $i=1,\ldots,m$ and $t=0,\ldots,T-1$}.$
This formulation is a finite-dimensional semi-definite program that can be solved numerically.
E.2 Converting PEP solutions into mathematical proofs
The Lagrangian attached to Problem (Primal) is given by
$L(G_{t},F_{t},\Lambda_{t},\lambda_{t,*}^{(i)},\lambda_{*,t}^{(i)})=$ $-\operatorname{Tr}(\Delta_{t}G_{t})-F_{t}\tilde{\Delta}-\operatorname{Tr}(\Lambda_{t}G_{t})$
$+\sum_{i=1}^{m}\lambda_{t,*}^{(i)}\left(F_{t}\bm{f}_{m+i}-F_{t}\bm{f}_{i}+\operatorname{Tr}(A_{t,*}^{(i)}G_{t})\right)$
$+\sum_{i=1}^{m}\lambda_{*,t}^{(i)}\left(F_{t}\bm{f}_{i}-F_{t}\bm{f}_{m+i}+\operatorname{Tr}(A_{*,t}^{(i)}G_{t})\right).$
If we have a primal-dual optimal solution $(\overline{G}_{t},\overline{F}_{t},\overline{\Lambda}_{t},(\overline{\lambda}_{t,*}^{(i)}),(\overline{\lambda}_{*,t}^{(i)}))$, it holds that, for all $G_{t}\in\mathcal{S}^{2m}$ and $F_{t}\in\mathbb{R}^{2m}$,
$L(\overline{G}_{t},\overline{F}_{t},\overline{\Lambda}_{t},(\overline{\lambda}_{t,*}^{(i)}),(\overline{\lambda}_{*,t}^{(i)}))\leq L(G_{t},F_{t},\overline{\Lambda}_{t},(\overline{\lambda}_{t,*}^{(i)}),(\overline{\lambda}_{*,t}^{(i)})).$
Due to complementary slackness, it holds that,
$L(\overline{G}_{t},\overline{F}_{t},\overline{\Lambda}_{t},(\overline{\lambda}_{t,*}^{(i)}),(\overline{\lambda}_{*,t}^{(i)}))=-\operatorname{Tr}(\Delta_{t}\overline{G}_{t})-\overline{F}_{t}\tilde{\Delta}=B_{t}.$
By strong duality (see Section E.3), we know that the existence of a primal-dual solution implies that $B_{t}=0$. As such, it holds that, for all $G_{t}\in\mathcal{S}^{2m}$ and $F_{t}\in\mathbb{R}^{2m}$,
$0\leq L(G_{t},F_{t},\overline{\Lambda}_{t},(\overline{\lambda}_{t,*}^{(i)}),(\overline{\lambda}_{*,t}^{(i)})),$
or in other terms that
$\operatorname{Tr}(\Delta_{t}G_{t})+F_{t}\tilde{\Delta}+\operatorname{Tr}(\overline{\Lambda}_{t}G_{t})$ $\leq\sum_{i=1}^{m}\overline{\lambda}_{t,*}^{(i)}\left(F_{t}\bm{f}_{m+i}-F_{t}\bm{f}_{i}+\operatorname{Tr}(A_{t,*}^{(i)}G_{t})\right)$ (11)
$+\sum_{i=1}^{m}\overline{\lambda}_{*,t}^{(i)}\left(F_{t}\bm{f}_{i}-F_{t}\bm{f}_{m+i}+\operatorname{Tr}(A_{*,t}^{(i)}G_{t})\right).$
As $(\overline{G}_{t},\overline{F}_{t},\overline{\Lambda}_{t},(\overline{\lambda}_{t,*}^{(i)}),(\overline{\lambda}_{*,t}^{(i)}))$ is primal-dual optimal, in particular it is primal-dual feasible, and hence $\overline{\Lambda}_{t}\succeq 0$ and $\overline{\lambda}_{t,*}^{(i)},\overline{\lambda}_{*,t}^{(i)}\geq 0$ for all $i=1,\ldots,m$. In particular, if $G_{t}$ and $F_{t}$ are generated by functions $f_{i}\in\mathcal{F}_{\mu,L}$ through Equations (10), it holds that the right-hand side of Equation (11) is nonpositive by Theorem E.1. By recalling that $E_{t+1}-E_{t}=\operatorname{Tr}(\Delta_{t}G_{t})+F_{t}\tilde{\Delta}$,
$E_{t+1}-E_{t}\leq-\operatorname{Tr}(\overline{\Lambda}_{t}G_{t})\leq 0,$
where the latter follows since $\overline{\Lambda}_{t},G_{t}\succeq 0$. Specifically, one can prove that $E_{t+1}-E_{t}\leq 0$ by rewriting it in the form of (11), using the obtained dual optimal solution, and writing $\operatorname{Tr}(\overline{\Lambda}_{t}G_{t})$ as an appropriate sum of squares based on the Cholesky factorization of $\overline{\Lambda}_{t}$.
E.3 Strong duality
Theorem E.2.
Problem (Dual) is the dual of Problem (Primal). Moreover, strong duality holds.
Proof.
To see that Problem (Dual) is in fact the dual of Problem (Primal), we consider the Lagrangian, which is given by
$L(G_{t},F_{t},\Lambda_{t},\lambda_{t,*}^{(i)},\lambda_{*,t}^{(i)})=$ $\operatorname{Tr}\left(\left(-\Delta_{t}+\sum_{i=1}^{m}\lambda_{t,*}^{i}A_{t,*}^{(i)}+\sum_{i=1}^{m}\lambda_{*,t}^{i}A_{*,t}^{(i)}-\Lambda_{t}\right)G_{t}\right)$
$+\left(\tilde{\Delta}+\sum_{i=1}^{m}\lambda_{*,t}^{i}(\bm{f}_{m+i}-\bm{f}_{i})+\sum_{i=1}^{m}\lambda_{*,t}^{i}(\bm{f}_{i}-\bm{f}_{m+i})\right)F_{t},$
from which we readily recover Problem (Dual) as the dual. In order to show that strong duality holds, we employ a standard Slater argument (boyd_convex_2023) and construct a feasible point $(G_{t},F_{t})$ to Problem (Primal) such that $G_{t}\succ 0$.
To do so, we follow the same ideas as in (taylor_smooth_2017, Theorem 6). Specifically, leveraging the discussion in Section E.1, we shall construct $m$ functions $f_{i}\in\mathcal{F}_{\mu,L}$ and points $x_{t},x_{*}\in\mathbb{R}^{d}$ such that the matrix $P_{t}$ given by Equation (10a) is upper triangular with positive entries on its diagonal. As such, assuming $d\geq 2m$, it will hold that $G_{t}=P_{t}^{T}P_{t}\succ 0$. We assume without loss of generality that $x_{*}=0$, and set $d=2m$.
We shall start by proving the result for $\mu=1$ and $L=6$. This is without loss of generality, as will be shown later.
We introduce the following functions $f_{i}$ and their gradients $\nabla f_{i}$ for $i=1,\ldots,m$,
$f_{i}\colon\mathbb{R}^{d}\to\mathbb{R},x\mapsto\frac{1}{2}x^{T}Q_{i}x+b_{i}x\quad\text{and}\quad\nabla f_{i}\colon\mathbb{R}^{d}\to\mathbb{R}^{d},x\mapsto Q_{i}x+b_{i}.$
We first define $f_{i}$ for $i=1,\ldots,m-1$. We let $Q_{i}$ be a block diagonal matrix of the form
$Q_{i}=\begin{pmatrix}Q_{i}^{1}&0\\ 0&Q\end{pmatrix}-\bm{e}_{m+i}\bm{e}_{m+i}^{T},$
where $Q_{i}^{1}\in\mathbb{R}^{(m-1)\times(m-1)}$ is given by
$Q_{i}^{1}=\text{diag}(2)+3\bm{e}_{i}\bm{e}_{i}^{T},$
and $Q\in\mathbb{R}^{(m+1)\times(m+1)}$ is a tridiagonal matrix whose diagonal elements are $4$ and whose off diagonal elements are $-1$. We note that $\sigma(Q_{i}^{1})\subset[2,5]$, and that, by the Greshgorian Theorem (quarteroni_numerical_2007), $\sigma(Q-\bm{e}_{i+1}\bm{e}_{i+1}^{T})\subset(1,6)$. Specifically, $\sigma(Q_{i})\subset(1,6)$ for all $i=1,\ldots,m-1$.
We define $x_{t}=[x^{1},x^{2}]$ with $x^{1}\in\mathbb{R}^{m-1}$ and $x^{2}\in\mathbb{R}^{m+1}$. Then
$\nabla f_{i}(x_{t})=g_{t}^{(i)}=2\sum_{j=1,j\neq i}^{m-1}\bm{e}_{j}x^{1}_{j}+5\bm{e}_{i}x^{1}_{i}+Qx^{2}-\bm{e}_{m+i}x^{2}_{i+1}+b_{i}.$
As such, we define $b_{i}$ as
$b_{i}=-2\sum_{j=1,j\neq i}^{m-1}\bm{e}_{j}x^{1}_{j}+\bm{e}_{m+i}x^{2}_{i+1}-\bm{e}_{m},$
and we impose that $x^{2}>0$ and that $Qx^{2}=[1,0,\ldots,0]\in\mathbb{R}^{m+1}$ (may be shown possible by a simple induction argument). This thus gives us that
$g_{t}^{(i)}=5\bm{e}_{i}x^{1}_{i}.$
Moreover, as $x_{*}=0$, we have that $g_{*}^{(i)}=b_{i}$. We additionally impose $x^{1}>0$. Note that $g_{t}^{(i)}$ has a positive entry in its $i$th component and 00s everywhere else. Moreover, $g_{*}^{(i)}$ has a positive component in its $m+i$th component and 00 in all subsequent components. The vector $x_{t}-x_{*}$ also has a positive entry in component $2m$. In order to show that $P_{t}$ is upper triangular with positive diagonal entries, we thus only need to construct $g_{t}^{(m)}$.
By construction we require $\nabla f(x_{*})=0$, namely that
$b_{m}=-\sum_{i=1}^{m-1}b_{i}=\sum_{i=1}^{m-1}(2m-1)\bm{e}_{i}x_{i}^{1}-\sum_{i=1}^{m-1}\bm{e}_{m+i}x_{i+1}^{2}+(m-1)\bm{e}_{m}.$
As such, we may define $Q_{m}$ to be the identity matrix (which satisfies $\sigma(Q_{m})=\{1\}$), such that
$\nabla f_{m}(x_{t})=g_{t}^{(m)}=\sum_{i=1}^{m-1}2m\bm{e}_{i}x_{i}^{1}+m\bm{e}_{m}.$
As such, $g_{t}^{(m)}$ has a positive entry in its $m$th component and zero entries in all subsequent components. As such, $P_{t}$ is upper triangular with positive diagonal entries, as wanted.
As such, if $(\mu,L)=(1,6)$, the result holds. For arbitrary $(\mu,L)$, we consider the following transformed functions
$\tilde{f}_{i}(x)=(f_{i}(x)-\tfrac{1}{2}\|x\|^{2})\frac{L-\mu}{5}+\mu\tfrac{1}{2}\|x\|^{2}.$
As $f_{i}\in\mathcal{F}_{1,6}$, we have $\tilde{f}_{i}\in\mathcal{F}_{\mu,L}$. Moreover, this transformation preserves the property on $P_{t}$, hence constructing a Slater point for all $(\mu,L)$. ∎
Appendix F More details on the numerical experiments
This section complements the numerics included in Section 4. Specifically, we compare the theoretical bounds obtained in Sections C and D with the numerical bounds obtained through the Performance Estimation Problem (PEP) methodology, as explained in Section E.
All our experiments were run with $L=1$, which can always be assumed to hold, up to rescaling the functions defining our problem.
F.1 Convex case, non-optimal step-size
Bias term
Figure 5 showcases the tightness of our theoretical bound on the bias term in the convex setting with non-optimal step-size by comparing it to the worst-case performance computed using the PEP methodology. Across iteration horizons $T=2,20,100$, the theoretical bound closely matches the numerical results obtained via PEP, indicating that the bound is not only valid but also tight. This agreement confirms that our theoretical analysis accurately captures the conditional worst-case behavior of SGD within the range of analyses permitted by our Lyapunov energies, and that the bound cannot be significantly improved under the given assumptions.
Figure 5: Theoretical and numerical bias term in the convex setting for the full range of step-sizes.
Variance term under tight bias
As a bi-product, we are able to exhibit upper bounds on the variance term, once the optimal bias has been established. In the short step-size range, depicted in Figure 6, we observe that our theoretical bounds closely align with those obtained through the PEP methodology, strongly suggesting that the bound cannot be further improved within the constraints of our Lyapunov analysis. In the large step-size range, shown in Figure 7, we observe similar results, although our results widely suffer from numerical instabilities.
Figure 6: Theoretical and numerical average variance in the convex setting for short step-sizes.
Figure 7: Theoretical and numerical average variance in the convex setting for large step-sizes.
F.2 Convex case, optimal step-size
We note that the results in the figure below are not tight due to our restriction that $a_{t}\equiv 1$, which is not the case for $\varepsilon>0$. Nevertheless, as $T\to\infty$, the PEP bound seems to converge to our (asymptotic) bound. The plot is generated by fixing a rate of $\rho=2\gamma-\varepsilon$, and minimizing the average variance.
Figure 8: Numerical and theoretical variance for optimal step-size and sub-optimal bias term.
Combining Figures 6 and 8, we note that it seems impossible to obtain a finite variance for the optimal bias term at the optimal step-size $\gamma L=1$. This confirms our discussion in Section C.2.
F.3 Strongly convex case, non-optimal step-size
Bias term
As in the convex setting, we hint towards the tightness of our theoretical bound on the bias in the strongly convex setting for non-optimal step-sizes in Figure 9. The overlap of the theoretical curve and the numerically generated one once again confirms our tightness.
Figure 9: Theoretical and numerical bias term in the strongly convex setting with $\mu=0.25$.
Variance term under tight bias
We moreover assess the tightness of the variance term, by first fixing the bias term to be the observed optimal, and then minimizing the total variance. The results are shown in Figure 10, where the numerics confirm the theoretical claims that we cannot obtain a finite variance term for the optimal bias term in the optimal step-size.
Figure 10: Theoretical and numerical variance in the strongly convex setting with $\mu=0.25$.
We define the relative error $\tau$ between the theoretical and numerical average variance as
$\tau=\frac{\overline{e}_{\text{opt}}-\overline{e}_{\text{theory}}}{\overline{e}_{\text{theory}}},$
where $\overline{e}_{\text{theory}}$ is the theoretical average variance term obtained in Theorem 3.3. Figure 11 shows this relative error for various values of $T$ over a range of values of step-sizes $\gamma L$ and strong convexity parameters $\mu$.
Figure 11: Relative errors $\tau$ in optimal average variance
F.4 Strongly convex case, optimal step-size
In Figure 12 we plot the numerical and theoretical variance term upon fixing an $\varepsilon$-tight bias term. Our bounds might not appear tight, but we suspect this to be due to numerical instabilities. To emphasize this, we include the same results using a different solver in Figure 13, whose results are widely different. We leave this for further investigation.
Figure 12: Theoretical and numerical variance in the strongly convex setting with $\mu=0.25$, in the optimal step-size (using MOSEK).
Figure 13: Theoretical and numerical variance in the strongly convex setting with $\mu=0.25$, in the optimal step-size (using CLARABEL).
F.5 Details on numerical instability
We note that the problems solved numerically are high-dimensional and grow with the number of iterations $T$. As such, for $T$ large, we suffer from numerical instability. Figure 14 shows that a slight change in tolerance provided to the solver might drastically modify the obtained solution. We shall not delve further into this.
Figure 14: Example of numerical instability for $(L,\mu)=(1,0.5)$.
Appendix G SGD with non-uniform sampling and mini-batching
All the main results in our paper can be applied to variants of (SGD), such as mini-batching or nonuniform sampling. This is because these variants can be seen as instances of (SGD) applied to different but equivalent problems. Our results can be applied as a black-box, and the results follow by computing the main constants of the equivalent problem, namely $L,\mu,\sigma_{*}^{2}$. This strategy was described in gower_sgd_2019, which contains all the formulas needed to compute those constants. For the sake of completeness, we include some details for the two aforementioned instances below.
G.1 SGD with non-uniform sampling
Consider a finite family of functions $f_{1},\dots,f_{n}$ which we assume to be $\mu_{i}$-strongly convex and $L_{i}$-smooth, for $\mu_{i}\geq 0$ and $L_{i}>0$. We want to study the non-uniform SGD algorithm, which computes
$x_{t+1}=x_{t}-\frac{\gamma}{np_{i_{t}}}\nabla f_{i_{t}}(x_{t}),$ (SGDp)
where $i_{t}\in\{1,\dots,n\}$ is sampled according to $\mathbb{P}(i_{t}=i)=p_{i}$, with $p_{i}>0$ and $\sum_{i=1}^{n}p_{i}=1$. Note that we are using here a specific renormalization of the step-size, following the ideas from gower_sgd_2019. It is immediate to see that this algorithm is exactly (SGD) applied to the problem
$\min_{x}\ f(x)=\mathbb{E}_{{\mathcal{P}}}\mathopen{}\left[{\hat{f}_{i}(x)}\right]\mathclose{},\text{ where }\hat{f}_{i}(x)\coloneqq\frac{1}{np_{i}}f_{i}(x),$
where the expectation is taken with respect $\mathcal{P}$, the distribution whose density is given by the probabilities $p_{i}$. Note that the function $f$ remains unchanged, meaning that the set of minimizers is the same. We can then compute (gower_sgd_2019, Propositions 3.7 & 3.8)) the following quantities:
  • $L$, which is the maximum among the constants $\operatorname{Lip}(\nabla\hat{f}_{i})$, so that we can say that each $\hat{f}_{i}$ is $L$-smooth. Clearly,
    $L=\max\limits_{i=1,\dots,n}\ \frac{L_{i}}{np_{i}}.$
  • $\mu$, which, for the same reasons as above, can be computed as
    $\mu=\min\limits_{i=1,\dots,n}\ \frac{\mu_{i}}{np_{i}}.$
  • $\sigma_{*}^{2}$, which is given by
    $\mathbb{E}_{{\mathcal{P}}}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}=\sum_{i=1}^{n}p_{i}\|\nabla\hat{f}_{i}(x_{*})\|^{2}=\frac{1}{n}\sum_{i=1}^{n}\frac{1}{np_{i}}\|\nabla\hat{f}_{i}(x_{*})\|^{2}.$
If we now impose that $\gamma L\in(0,2)$, we can apply all our results and obtain bounds on the iterates of (SGDp) depending on the above constants.
G.2 SGD with mini-batching
Consider a finite family of functions $f_{1},\dots,f_{n}$ which we assume to be $\mu_{i}$-strongly convex and $L_{i}$-smooth, for $\mu_{i}\geq 0$ and $L_{i}>0$. We want to study the mini-batch SGD algorithm, which computes at each iteration a mini-batch of fixed size $1\leq b\leq n$, that is
$x_{t+1}=x_{t}-\frac{\gamma}{b}\sum_{i\in B_{t}}\nabla f_{i}(x_{t}),$ (SGDb)
where $B_{t}$ is sampled i.i.d. and uniformly among the subsets of size $b$ of $\{1,\dots,n\}$. It is a simple exercise to see that this algorithm is exactly (SGD) applied to the problem
$\min_{x}\ f(x)=\mathbb{E}_{{\mathcal{B}}}\mathopen{}\left[{\hat{f}_{B}(x)}\right]\mathclose{},\text{ where }\hat{f}_{B}(x)\coloneqq\frac{1}{b}\sum_{i\in B}\nabla f_{i}(x),$
and where the expectation is taken with respect to $\mathcal{B}$ which is the uniform law over
$\text{batch}_{b}\coloneqq\{B\subset\{1,\dots,n\}\ :\ |B|=b\}.$
To be able to apply our results, we need to verify that this new problem verifies our assumptions. More precisely, all we need is for it to verify (SEC) for some $L$ and some $\mu$. This can be calculated using the results in gower_sgd_2019.
Lemma G.1.
Consider Problem A.1 with $\mathcal{I}$ finite. Assume that each function $f_{i}$ is $\mu_{i}$-convex and $L_{i}$-smooth. Then the family of functions $(f_{B})_{B\in\text{batch}_{b}}$ verifies (SEC) with constants
  • $L=(1-r)\operatorname{Lip}(\nabla f)+r\max_{i}L_{i}$, with $r=\tfrac{n(b-1)}{b(n-1)}\in[0,1]$,
  • $\mu=\min_{i}\mu_{i}$.
Moreover, we have that
$\sigma_{*}^{2}=\mathbb{E}_{{\mathcal{B}}}\mathopen{}\left[{\|\nabla f_{B}(x_{*})\|^{2}}\right]\mathclose{}=\frac{n-b}{nb(n-1)}\sum_{i=1}^{n}\|\nabla f_{i}(x_{*})\|^{2}.$
Proof.
Let $\mu\coloneqq\min_{i}\mu_{i}$, such that each $f_{i}$ is $\mu$-convex. Similarly, because each $\hat{f}_{B}$ is an average of $f_{i}$’s, we know that each $\hat{f}_{B}$ is $\mu$-convex. We can therefore define
$g=f-\tfrac{\mu}{2}\|\cdot\|^{2},\quad g_{i}=f_{i}-\tfrac{\mu}{2}\|\cdot\|^{2}\quad\text{ and }\quad\hat{g}_{B}=\hat{f}_{B}-\tfrac{\mu}{2}\|\cdot\|^{2}$
all of which are convex functions. By definition, we know that each $\hat{g}_{B}$ is an average of functions $g_{i}$ which are each $(L_{i}-\mu)$-smooth and convex. Invoking (gower_sgd_2019, Section G) shows that the family $(\hat{g}_{B})_{B\in\text{batch}_{b}}$ verifies (EC) for constants $(\mathcal{L},0)$ where
$\mathcal{L}=(1-r)\operatorname{Lip}(\nabla g)+r\max_{i}\operatorname{Lip}(\nabla g_{i}),\text{ with }r=\tfrac{n(b-1)}{b(n-1)}\in[0,1].$
Using Lemma G.2 which can be found below, we see that this is equivalent to the family $(\hat{f}_{B})_{B\in\text{batch}_{b}}$ verifying (EC) for constants $(\mathcal{L}+\mu,\mu)$. Because $\operatorname{Lip}(\nabla g)=\operatorname{Lip}(\nabla f)-\mu$ and $\operatorname{Lip}(\nabla g_{i})=\operatorname{Lip}(\nabla f_{i})-\mu$, we conclude by setting $L=\mathcal{L}+\mu$. ∎
Lemma G.2.
Consider Problem A.1. Let $L\geq\mu\geq 0$, and let $(x,y)\in\mathcal{H}$ be fixed. Then the following are equivalent:
  1. 1.
    the family of functions $(f_{i})_{i\in\mathcal{I}}$ verify (EC) with constants $(L,\mu)$;
  2. 2.
    the family of functions $(g_{i})_{i\in\mathcal{I}}$ verify (EC) with constants $(L-\mu,0)$, with $g_{i}=f_{i}-\tfrac{\mu}{2}\|\cdot\|^{2}$.
Proof.
It suffices to show that, if $\hat{f},\hat{g}\colon\mathbb{R}^{d}\to\mathbb{R}$ are such that $\hat{g}=\hat{f}-\tfrac{\mu}{2}\|\cdot\|^{2}$, then
$\tfrac{1}{2}\|\nabla\hat{f}(y)$ $-\nabla\hat{f}(x)\|^{2}+\tfrac{\mu L}{2}\|y-x\|^{2}-(L-\mu)\big{[}\hat{f}(y)-\hat{f}(x)\big{]}-\langle\mu\nabla\hat{f}(y)-L\nabla\hat{f}(x),y-x\rangle$
$=\tfrac{1}{2}\|\nabla\hat{g}(y)-\nabla\hat{g}(x)\|^{2}-(L-\mu)\big{[}\hat{g}(y)-\hat{g}(x)\big{]}+(L-\mu)\langle\nabla\hat{g}(x),y-x\rangle.$ (12)
We start from the right-hand side, substituting $\hat{g}(z)=\hat{f}(z)-\tfrac{\mu}{2}\|z\|^{2}$ and $\nabla\hat{g}(z)=\nabla\hat{f}(z)-\mu z$ for $z=x$ and $z=y$, to obtain
$\tfrac{1}{2}\|\nabla\hat{g}(y)$ $-\nabla\hat{g}(x)\|^{2}-(L-\mu)\big{[}\hat{g}(y)-\hat{g}(x)\big{]}+(L-\mu)\langle\nabla\hat{g}(x),y-x\rangle$
=\ = $\tfrac{1}{2}\|\nabla\hat{f}(y)-\nabla\hat{f}(x)\|^{2}-\mu\langle\nabla\hat{f}(y)-\nabla\hat{f}(x),y-x\rangle+\tfrac{\mu^{2}}{2}\|y-x\|^{2}$
$-(L-\mu)\big{[}\hat{f}(y)-\hat{f}(x)\big{]}+\tfrac{\mu}{2}(L-\mu)\big{[}\|y\|^{2}-\|x\|^{2}\big{]}$
$+(L-\mu)\langle\nabla\hat{f}(x),y-x\rangle-\mu(L-\mu)\langle x,y-x\rangle$
=\ = $\tfrac{1}{2}\|\nabla\hat{f}(y)-\nabla\hat{f}(x)\|^{2}-(L-\mu)\big{[}\hat{f}(y)-\hat{f}(x)\big{]}-\langle\mu\nabla\hat{f}(y)-L\nabla\hat{f}(x),y-x\rangle$
$+\tfrac{\mu}{2}\left(\mu\|y-x\|^{2}+(L-\mu)\big{[}\|y\|^{2}-\|x\|^{2}\big{]}-2(L-\mu)\langle x,y-x\rangle\right)$
=\ = $\tfrac{1}{2}\|\nabla\hat{f}(y)-\nabla\hat{f}(x)\|^{2}-(L-\mu)\big{[}\hat{f}(y)-\hat{f}(x)\big{]}-\langle\mu\nabla\hat{f}(y)-L\nabla\hat{f}(x),y-x\rangle$
$+\tfrac{\mu^{2}}{2}\big{[}\|y-x\|^{2}-\|y\|^{2}-\|x\|^{2}+2\langle x,y\rangle\big{]}+\tfrac{\mu L}{2}\big{[}\|y\|^{2}+\|x\|^{2}-2\langle x,y\rangle\big{]}.$
This is precisely the left-hand side of (12). ∎
Appendix H Stochastic Proximal algorithm
H.1 Problem formulation
In this section we consider an extension of Problem A.1, where we drop the assumption that the $f_{i}$ are differentiable or take finite values. We do impose though that the functions are convex.
Problem H.1.
[…]
  1. 1.
    the problem is well-defined, in the sense that $i\mapsto f_{i}(x)$ is $\mathcal{D}$-measurable;
  2. 2.
    the problem is well-posed, in the sense that ${\rm{argmin}}~{}f\neq\emptyset$;
  3. 3.
    the problem is convex, in the sense that each $f_{i}$ is convex, lower semi-continuous and proper.
An algorithm of choice to solve Problem H.1 is the Stochastic Proximal algorithm. It is defined as
$x_{t+1}=\operatorname{prox}_{\gamma f_{i_{t}}}(x_{t}),$ (SProx)
where $i_{t}\in\mathcal{I}$ is sampled i.i.d. from the distribution $\mathcal{D}$, and where $\operatorname{prox}_{\gamma f_{i}}$ is the proximal operator of $f_{i}$ with step-size $\gamma$, defined by
$\operatorname{prox}_{\gamma f_{i}}(x)=\underset{y\in\mathcal{H}}{\rm{argmin}}~{}f_{i}(y)+\frac{1}{2\gamma}\|y-x\|^{2}.$
Our assumption that $f_{i}$ is convex and lower semi-continuous ensures that $\operatorname{prox}_{\gamma f_{i}}$ is well-defined.
Literature review.
Although the proximal algorithm is widely studied and well understood in the deterministic case, its complexity seems to be poorly understood in the stochastic setting. To the best of our knowledge, the only complexity results available for (SProx) can be divided in a few distinct categories:
  1. 1.
    Complexity results for the stochastic projection algorithm, where each function $f_{i}$ is the indicator function of some nonempty closed convex set $C_{i}$. In this case, the problem becomes a feasibility problem, which consists in finding $x_{*}\in C\coloneqq\cap_{i}C_{i}$. Here the proximal operator of $f_{i}$ becomes the projection onto $C_{i}$, and (SProx) performs random projections onto the sets $C_{i}$. Results typically show that
    $\mathbb{E}\mathopen{}\left[{\operatorname{dist}(\bar{x}_{T};C_{i})^{2}}\right]\mathclose{}\leq\frac{\operatorname{dist}(x_{0};C)^{2}}{T},$
    see for instance (nedic_random_2010, Proposition 6). Other authors might additionally assume that the intersection between the sets $C_{i}$ is regular to obtain better bounds, which is standard when studying deterministic feasibility problems lewis_local_2009. For instance, necoara_randomized_2019 considers the well-known linearly regular intersection assumption
    $(\exists\kappa>0)(\forall x\in\mathcal{H})\quad\operatorname{dist}(x,C)^{2}\leq\kappa\mathbb{E}\mathopen{}\left[{\operatorname{dist}(x;C_{i})^{2}}\right]\mathclose{},$
    and is able to exploit it to derive a stronger linear rate.
  2. 2.
    Complexity results under the assumption that each $f_{i}$ takes finite values, and is $G$-Lipschitz continuous. This setting was explored by bertsekas_incremental_2011, and more recently by patrascu_nonasymptotic_2018, davis_stochastic_2019 and asi_stochastic_2019. Under this assumption, the (sub)gradients are $G$-bounded, and it is possible to show that
    $f(\bar{x}_{T})-\inf f\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\frac{\gamma G^{2}}{2}.$ (13)
  3. 3.
    Complexity results under the assumption that each $f_{i}$ takes finite values and is $L$-smooth. The following bound was obtained in (traore_variance_2024, Theorem 4.3)
    $\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\inf f}\right]\mathclose{}\leq\frac{\|x_{0}-x_{*}\|^{2}}{\gamma T}+2\gamma\sigma_{*}^{2}.$
    Similar results can also be obtained by replacing the smoothness assumption with some weaker version, such as the $(L_{0},L_{1})$ assumption (tovmasyan_revisiting_2025).
  4. 4.
    Complexity results under the assumption that the problem is $\mu$-strongly convex with $\mu>0$. This setting was explored in (asi_stochastic_2019, Proposition 5) and in (richtarik_unified_2024, Theorem 5.3). The latter was able to obtain the following complexity bound, with the additional assumption that the functions are differentiable:
    $\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\left(\frac{1}{1+\gamma\mu}\right)^{2T}\|x_{0}-x_{*}\|^{2}+\frac{\gamma\sigma_{*}^{2}}{\mu(2+\gamma\mu)}.$ (14)
We stress the fact that the gap between the settings considered in the literature is huge. Indeed, we have on the one side functions which are intrinsically taking +$\infty$+ ∞ values, and on the other side functions taking finite values, with various levels of imposed regularity. As far as we know, only the results in the strongly convex case avoid making any regularity assumption on the functions $f_{i}$. In this case however we are discarding the feasibility problem, as it is not reasonable to assume that indicator functions are strongly convex, unless they are singletons, which would trivialize the problem. It is therefore clear that no unified analysis for (SProx) is available.
Studying the complexity of (SProx) is not trivial.
Before trying to obtain bounds on (SProx), one must determine what metric to estimate.
In the deterministic setting, a standard result shows that the function value gap $f(x_{t})-\min f$ is upper bounded by $O(\tfrac{1}{t})$. For SGD, it is standard as well to upper bound the expected averaged function value gap $\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]$. The problem with the stochastic proximal algorithm is that we cannot hope to ever bound a function value gap. Indeed, nothing guarantees that the iterates $x_{t}$ (or their average) remain in the domain of $f$. We of course know that $x_{t+1}\in dom(f_{i_{t}})$, but there is no reason for it to belong in the intersection of all domains. This is even clearer if one considers the stochastic projection algorithm, where $f$ is the indicator function of the intersection $\cap_{i}C_{i}$. For $f(x_{t})$ to be finite, we need $x_{t}$ to already be a solution. As such, we cannot expect bounds based on the function value gap.
One could think about providing a bound for the distance between the current iterate and a solution of the problem (or the distance to the set of solutions). However, this must also be discarded, as it is known in the deterministic case that such distance can decrease arbitrarily slowly (garrigos_convergence_2023). This is why, for convex problems, we usually do not have bounds on terms $\|x_{t}-x_{*}\|$, contrary to strongly convex problems.
In this section, we provide bounds for a different function gap, which is always well-defined, and which in some cases can be connected to well-known quantities, allowing to recover existing results.
H.2 Stochastic Proximal algorithm as SGD
The proximal algorithm is a particular case of the gradient descent algorithm. This can be stated formally, by introducing the notion of Moreau regularization for a convex function. Given a proper convex lower semi-continuous function $f\colon\mathcal{H}\to\mathbb{R}\cup\{+\infty\}$, given $\gamma>0$, we define its $\gamma$-Moreau regularization as the function $f^{\gamma}$ by
$f^{\gamma}(x)=\inf\limits_{y\in H}f(y)+\frac{1}{2\gamma}\|y-x\|^{2}.$
The Moreau regularization enjoys a few properties (bauschke_convex_2017, Proposition 12.30), namely
  1. 1.
    $f^{\gamma}\colon\mathcal{H}\to\mathbb{R}$ takes finite values, and is differentiable;
  2. 2.
    $\nabla f^{\gamma}$ is $\tfrac{1}{\gamma}$-Lipschitz continuous;
  3. 3.
    $\nabla f^{\gamma}(x)=\frac{x-\operatorname{prox}_{\gamma f}(x)}{\gamma}$.
In particular, we see that computing $\operatorname{prox}_{\gamma f}(x)$ is the same as computing one step of gradient descent for the function $f^{\gamma}$, satisfying $\operatorname{Lip}(\nabla f^{\gamma})=\frac{1}{\gamma}$, with step-size $\gamma$:
$\operatorname{prox}_{\gamma f}(x)=x-\gamma\nabla f^{\gamma}(x).$
Therefore, it is clear that the stochastic proximal algorithm can be seen as an instance of SGD, whose step-size is exactly the optimal step-size:
$x_{t+1}=\operatorname{prox}_{\gamma f_{i}}(x_{t})=x_{t}-\gamma\nabla f_{i}^{\gamma}(x_{t}).$
This connection between (SProx) and (SGDE) is not new, and was already discussed and exploited in patrascu_nonasymptotic_2018, necoara_randomized_2019. The novelty in our paper is that we are able to exploit easily this connection, since we were able to derive complexity rates for SGD when $\gamma L=1$.
We can see that the stochastic proximal algorithm aims at minimizing another function, which is given by
$F^{\gamma}(x)\coloneqq\mathbb{E}\mathopen{}\left[{f_{i}^{\gamma}(x)}\right]\mathclose{}.$
The following result can immediately be derived from our bounds for SGD with optimal step-size in Theorem C.5.
Theorem H.2 (Bound for SProx - General case).
Consider Problem H.1. Let $x_{t}$ be generated by (SProx), with any step-size $\gamma>0$. Then, for all $T\geq 1$,
$\mathbb{E}\left[F^{\gamma}(\bar{x}_{T})-\min F^{\gamma}\right]\leq\frac{\|x_{0}-x_{*}^{\gamma}\|^{2}}{(2-\varepsilon)\gamma T}+\frac{(2+\varepsilon)\gamma\sigma^{2}_{*}(\gamma)}{\varepsilon(2-\varepsilon)},$
where $\sigma^{2}_{*}(\gamma)=\mathbb{E}\left[\|\nabla f_{i}^{\gamma}(x_{*}^{\gamma})\|^{2}\right]$, and $x_{*}^{\gamma}\in{\rm{argmin}}~{}F^{\gamma}$. If we further assume that each $f_{i}$ is $\mu$-strongly convex ($\mu>0$), then with $\mu_{\gamma}\coloneqq\tfrac{\mu}{1+\gamma\mu}$, we have
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}^{\gamma}\|^{2}}\right]\mathclose{}\leq\left(1-\tfrac{\mu_{\gamma}}{L}\right)^{2T}\|x_{0}-x_{*}^{\gamma}\|^{2}+\frac{2}{\mu_{\gamma}L(2-\tfrac{\mu_{\gamma}}{L})}\sigma_{*}^{2}(\gamma).$
Proof.
Apply Theorem C.5 and Corollary D.5 with $\gamma L=1$. For the strongly convex case, we use the fact that if the $f_{i}$ are $\mu$-strongly convex, then $f_{i}^{*}$ is $\tfrac{1}{\mu}$-smooth. This in turn implies that (bauschke_convex_2017, Proposition 13.24.iii)
$(f^{\gamma}_{i})^{*}=f^{*}+\tfrac{\gamma}{2}\|\cdot\|^{2}$
is $\tfrac{1}{\mu}+\gamma$ smooth, allowing to conclude that $f^{\gamma}_{i}$ is strongly convex with constant $\tfrac{1}{{1}/{\mu}+\gamma}=\mu_{\gamma}$. ∎
The previous theorem might not appear to be satisfying in the convex case, as it yields a bound on the regularized function value gap which is not clearly related to our original problem. We will see that in some cases, however, it is. The rest of this section is dedicated to investigating those cases. We will see that in the case of interpolation, and for feasibility problems, we recover the standard bounds from literature. We will moreover investigate the Lipschitz and smooth cases, where we can recover meaningful bounds, even though they are less good than the existing literature.
It remains to discuss the strongly convex case (with no further assumptions). Our bound reduces to
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\left(1-\tfrac{\mu}{L}\tfrac{1}{1+\gamma\mu}\right)^{2T}\|x_{0}-x_{*}\|^{2}+O\left(\frac{1}{\mu}\right),$
which is comparable with the bound $\exp\left(-\mu T\right)+O\left({\mu}^{-1}\right)$ obtained in (asi_stochastic_2019, Proposition 5). Nevertheless, our result is clearly outperformed by (richtarik_unified_2024, Theorem 6.4) which obtains the cleaner bound (14), even if we take into account the fact that it requires differentiability of the functions $f_{i}$.
Before continuing, we stress the fact that, while we focus on the vanilla Stochastic Proximal algorithm, all our results remain valid for standard extensions such as non-uniform sampling or mini-batching. It suffices to combine our results with the techniques from Section G.
H.3 The case of interpolation
Recall from Definition A.11 that interpolation corresponds to the case where almost every function $f_{i}$ share a common minimizer. Note that in Section A.3 we considered differentiable functions, but the results of Lemma A.12 trivially extend to the setting of Problem H.1.
Proposition H.3.
Consider Problem H.1, and assume that interpolation holds for the family $(f_{i})_{i\in\mathcal{I}}$. Then
  1. 1.
    Interpolation holds for the family $(f_{i}^{\gamma})_{i\in\mathcal{I}}$.
  2. 2.
    $\inf F^{\gamma}=\inf f$.
  3. 3.
    ${\rm{argmin}}~{}F^{\gamma}={\rm{argmin}}~{}f$.
Proof.
It is well-known that $\inf f_{i}^{\gamma}=\inf f_{i}$ and ${\rm{argmin}}~{}f_{i}^{\gamma}={\rm{argmin}}~{}f_{i}$, see e.g. (bauschke_convex_2017, Proposition 12.9). Our assumption that interpolation holds for the family $(f_{i})_{i\in\mathcal{I}}$ means that $\cap_{i}{\rm{argmin}}~{}f_{i}\neq\emptyset$, from which we deduce that $\cap_{i}{\rm{argmin}}~{}f_{i}^{\gamma}\neq\emptyset$, proving the first point. Then we can use Lemma A.12 to write that
${\rm{argmin}}~{}F^{\gamma}=\cap_{i\in\mathcal{I}}{\rm{argmin}}~{}f_{i}^{\gamma}=\cap_{i\in\mathcal{I}}{\rm{argmin}}~{}f_{i}={\rm{argmin}}~{}f,$
and
$\inf F^{\gamma}=\mathbb{E}\mathopen{}\left[{\inf f_{i}^{\gamma}}\right]\mathclose{}=\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}=\inf f,$
concluding the proof. ∎
This result shows that the function value gap $F^{\gamma}(x)-\inf F^{\gamma}$ is a meaningful metric, in the sense that it provides a faithful and continuous measure of how far we are from optimality. This is particularly clear in the case of stochastic projections, which we will consider next. But first, we specialize our bound from Theorem H.2 to this interpolation setting.
Theorem H.4 (Bounds for SProx - Interpolation case).
Consider Problem H.1 and assume that interpolation holds. Let $x_{t}$ be generated by (SProx) with step-size $\gamma>0$. Then for every $x_{*}\in{\rm{argmin}}~{}f$,
$\mathbb{E}\left[F^{\gamma}(\bar{x}_{T})-\inf f\right]\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}.$
If we further assume that each $f_{i}$ is $\mu$-strongly convex ($\mu>0$), then with $\mu_{\gamma}\coloneqq\tfrac{\mu}{1+\gamma\mu}$ we have
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\left(1-\tfrac{\mu_{\gamma}}{L}\right)^{2T}\|x_{0}-x_{*}\|^{2}.$
Proof.
We apply Theorem H.2 taking into account that ${\rm{argmin}}~{}F^{\gamma}={\rm{argmin}}~{}f$ and $\sigma_{*}^{2}(\gamma)=0$ because interpolation holds for the regularized problem as well, as seen in Proposition H.3. In this case, we take $\varepsilon\to 0$ to obtain the result in the convex case. ∎
The above result is new in the convex case. We are not aware of any result providing bounds for (SProx) under the sole assumption of interpolation.
The result in the strongly convex case can be compared to (tovmasyan_revisiting_2025, Theorem 6.1), where the authors obtain a worse rate, under the additional restrictions that $\gamma\leq\tfrac{\mu}{2}$, that the functions are differentiable, and that a bound on the gradient variance called star-similarity is verified. Note nevertheless that they only need the strong convexity of the overall function $f$, and are able to obtain a result without interpolation in (tovmasyan_revisiting_2025, Theorem 6.4), which we are not.
H.4 The case of indicator functions
Here, we assume that $f_{i}=\delta_{C_{i}}$, the indicator function of some nonempty closed convex set $C_{i}\subset\mathcal{H}$. We recall that $\delta_{C_{i}}(x)$ is 00 if $x\in C_{i}$ and is +$\infty$+ ∞ if $x\notin C_{i}$. In this case, (SProx) specializes into the Stochastic Projection algorithm:
$x_{t+1}=\operatorname{proj}_{C_{i_{t}}}(x_{t}).$ (SProj)
A first observation to make in this version of Problem H.1 is that, for the problem to be well-posed, that is for ${\rm{argmin}}~{}f$ to be nonempty, we need $C\coloneqq\cap_{i}C_{i}$ to be nonempty. But $C_{i}={\rm{argmin}}~{}f_{i}$, so this means that interpolation hold for well-posed feasibility problems. Note that one could also investigate (SProj) when the problem is not feasible, as done by nedic_random_2010, but this goes beyond the scope of this paper.
We may compute $f^{\gamma}(x)=\tfrac{1}{2\gamma}\operatorname{dist}(x;C_{i})^{2}$, and therefore the regularized function value gap becomes
$F^{\gamma}(x)-\inf F^{\gamma}=\frac{1}{2\gamma}\mathbb{E}\mathopen{}\left[{\operatorname{dist}(x;C_{i})^{2}}\right]\mathclose{},$ (15)
which clearly is a good measure of how far $x$ is from the set of solutions $C=\cap_{i}C_{i}$. We can now directly obtain a bound for (SProj), which is exactly the one obtained in (nedic_random_2010, Proposition 6).
Corollary H.5 (Bounds for SProj).
Consider Problem H.1 where $f_{i}=\delta_{C_{i}}$, and let $C=\cap_{i}C_{i}$. Let $x_{t}$ be generated by (SProj). Then,
$\mathbb{E}\mathopen{}\left[{\operatorname{dist}(\bar{x}_{T};C_{i})^{2}}\right]\mathclose{}\leq\frac{\operatorname{dist}(x_{0};C)^{2}}{T}.$
Proof.
Apply Theorem H.4 with (15). Moreover, we take $x_{*}$ to be the projection of $x_{0}$ onto $C$. ∎
H.5 The case of functions with finite values
If we assume that the functions $f_{i}$ take finite value, the previously mentioned issue of the function value gap $f(x)-\inf f$ not being a good metric is no longer applicable. The next proposition shows that the former can be controlled by the regularized function gap $F^{\gamma}(x)-\inf F^{\gamma}$.
In what follows, we will denote by $\partial f(x)$ the subdifferential of a function $f$ at a point $x$:
$\partial f(x)=\{g\in\mathcal{H}\ |\ (\forall y\in\mathcal{H})\ f(y)-f(x)-\langle g,y-x\rangle\geq 0\}.$
We will further note $\partial^{0}f(x)$ to denote the element of minimal norm within $\partial f(x)$. The latter is well-defined because $\partial f(x)$ is nonempty (bauschke_convex_2017, Proposition 16.27), convex and closed (bauschke_convex_2017, Proposition 16.4).
Proposition H.6.
Consider problem H.1, and assume that each function $f_{i}$ takes finite values. Then, for all $x\in\mathcal{H}$,
$f(x)-\inf f\leq F^{\gamma}(x)-\inf F^{\gamma}+\frac{\gamma}{2}\mathbb{E}\mathopen{}\left[{\|\partial^{0}f_{i}(x)\|^{2}}\right]\mathclose{}.$
Proof.
First, it is a standard result that $f_{i}^{\gamma}\leq f_{i}$. So we deduce easily that $F^{\gamma}\leq f$ and that $\inf F^{\gamma}\leq\inf f$. The rest of the proof is devoted to prove that $f(x)\leq F^{\gamma}(x)+\tfrac{\gamma}{2}\mathbb{E}\mathopen{}\left[{\|\partial^{0}f_{i}(x)\|^{2}}\right]\mathclose{}.$ We start by writing
$f_{i}(x)-f_{i}^{\gamma}(x)$ $=$ $\sup_{y\in\mathcal{H}}f_{i}(x)-f_{i}(y)-\frac{1}{2\gamma}\|y-x\|^{2}.$
As we assumed the functions $f_{i}$ to be convex, we know that
$f_{i}(x)-f_{i}(y)\leq-\langle\partial^{0}f_{i}(x),y-x\rangle.$
Combining all this, we obtain
$f_{i}(x)-f_{i}^{\gamma}(x)$$\leq$ $\sup_{y\in\mathcal{H}}-\langle\partial^{0}f_{i}(x),y-x\rangle-\frac{1}{2\gamma}\|y-x\|^{2}$
$=$ $\sup_{z\in\mathcal{H}}-\langle\partial^{0}f_{i}(x),z\rangle-\frac{1}{2\gamma}\|z\|^{2}.$
Writing the optimality conditions, we deduce that the optimal $z$ is $z=-\gamma\partial^{0}f_{i}(x)$. Injecting this optimal solution leads to
$f_{i}(x)-f_{i}^{\gamma}(x)\leq\frac{\gamma}{2}{\|\partial^{0}f_{i}(x)\|^{2}},$
and the desired inequality follows after taking expectation. ∎
Corollary H.7 (Bounds for SProx - Finite case).
Consider Problem H.1 and assume that each function $f_{i}$ takes finite values. Let $x_{t}$ be generated by (SProx), with step-size $\gamma>0$. Then
$\mathbb{E}\left[f(\bar{x}_{T})-\inf f\right]\leq\frac{\|x_{0}-x_{*}^{\gamma}\|^{2}}{\gamma T}+3\gamma\sigma_{*}^{2}(\gamma)+\frac{\gamma}{2}\mathbb{E}\mathopen{}\left[{\|\partial^{0}f_{i}(\bar{x}_{T})\|^{2}}\right]\mathclose{},$
where $x_{*}^{\gamma}\in{\rm{argmin}}~{}F^{\gamma}$ and $\sigma^{2}_{*}(\gamma)=\mathbb{E}\left[\|\nabla f_{i}^{\gamma}(x_{*}^{\gamma})\|^{2}\right]$.
Proof.
Combine Theorem H.2 with Proposition H.6, and take $\varepsilon=1$ to simplify terms. ∎
This general bound, albeit totally new under such general assumptions, is not very useful. The main flaw is that it depends on $\mathbb{E}\mathopen{}\left[{\|\partial^{0}f_{i}(\bar{x}_{T})\|^{2}}\right]\mathclose{}$ which, without further assumptions, we cannot control. But under an additional assumption of Lipschitzness or smoothness, this is possible.
H.6 The case of Lipschitz problems
We first translate the bound in Proposition H.6 to the setting of Lipschitz functions.
Proposition H.8.
Consider problem H.1, and assume that each function $f_{i}$ is $G$-Lipschitz continuous. Then, for all $x\in\mathcal{H}$
$f(x)-\inf f\leq F^{\gamma}(x)-\inf F^{\gamma}+\frac{\gamma G^{2}}{2}.$
Proof.
This is a direct consequence of Proposition H.6, where we further use the fact that the $f_{i}$ are $G$-Lipschitz. This means that their subgradients are all bounded by $G$, and in particular that $\|\partial^{0}f_{i}(x)\|\leq G$, which is what we need to prove the claim. ∎
This thus allows us to obtain bound for the stochastic proximal algorithm in the same setting.
Corollary H.9 (Bounds for SProx - Lipschitz case).
Consider Problem H.1 and assume that each function $f_{i}$ is $G$-Lipschitz continuous. Let $x_{t}$ be generated by (SProx), with step-size $\gamma>0$. Then
$\mathbb{E}\left[f(\bar{x}_{T})-\inf f\right]\leq\frac{\|x_{0}-x_{*}^{\gamma}\|^{2}}{\gamma T}+4\gamma G^{2},$
where $x_{*}^{\gamma}\in{\rm{argmin}}~{}F^{\gamma}$.
Proof.
Combine Theorem H.2 with Proposition H.6, and take $\varepsilon=1$ to simplify terms. Moreover, we use the fact that
$\sigma^{2}_{*}(\gamma)=\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}^{\gamma}(x_{*}^{\gamma})\|^{2}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*}^{\gamma})\|^{2}}\right]\mathclose{}\leq G^{2},$
where we used the property that $\|\nabla f_{i}^{\gamma}(x)\|\leq\|\nabla f_{i}(x)\|$, a standard fact which can be, for instance, found in (patrascu_nonasymptotic_2018, Lemma 3). ∎
We do not recover exactly the bound (13) from (bertsekas_incremental_2011, Theorem 5). In particular, our constants are worse by a factor $8$. This should not come as a surprise, as we are relying on a result for optimal step-sizes, namely Theorem C.5, which has a singularity for $\gamma L=1$ without uniform variance bound, and in a second time we assume that the functions have bounded gradients. It would be more efficient to exploit the bound on the gradients from the start, as done in (bertsekas_incremental_2011).
H.7 The case of smooth problems
We again translate the bound in Proposition H.6 to the setting of smooth functions.
Proposition H.10.
Consider problem H.1, and assume that each function $f_{i}$ is $L$-smooth and bounded from below. If $\gamma<\tfrac{1}{L}$, then
$f(x)-\inf f\leq\frac{1}{1-\gamma L}\left(F^{\gamma}(x)-\inf F^{\gamma}\right)+\frac{\gamma\Delta_{*}}{1-\gamma L},$
where $\Delta_{*}\coloneqq\inf f-\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}$.
Proof.
We start from Proposition H.6. If the functions are $L$-smooth, a now standard variance transfer argument shows that (see (garrigos_handbook_2024, Lemma 4.19))
$\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x)\|^{2}}\right]\mathclose{}\leq 2L(f(x)-\inf f)+2\Delta_{*},\quad\Delta_{*}\coloneqq\inf f-\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}.$
Using this bound gives us
$f(x)-F^{\gamma}(x)\leq\gamma L(f(x)-\inf f)+\gamma\Delta_{*}.$
In the end, we obtained
$(1-\gamma L)(f(x)-\inf f)\leq F^{\gamma}(x)-\inf F^{\gamma}+\gamma\Delta_{*}.$
The conclusion follows after dividing by $1-\gamma L$. ∎
Note that the constant $\Delta_{*}$ appearing in our bound has the same status as $\sigma^{2}_{*}$, in the sense that it is an interpolation constant, see garrigos_handbook_2024 for more details. In particular, $\Delta_{*}=0$ if, and only if, interpolation holds, and, under a $L$-smoothness assumption, it holds that $\sigma_{*}^{2}\leq 2L\Delta_{*}$.
Corollary H.11 (Bounds for SProx - Smooth case).
Consider Problem H.1 and assume that each function $f_{i}$ is $L$-smooth. Let $x_{t}$ be generated by (SProx), with step-size $\gamma\in(0,\tfrac{1}{L})$. Then, for all $T\geq 1$,
$\mathbb{E}\left[f(\bar{x}_{T})-\inf f\right]\leq\frac{\|x_{0}-x_{*}^{\gamma}\|^{2}}{(1-\gamma L)\gamma T}+\gamma\frac{6L+1}{1-\gamma L}\Delta_{*},$
where $x_{*}^{\gamma}\in{\rm{argmin}}~{}F^{\gamma}$ and $\Delta_{*}=\inf f-\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}$.
Proof.
Combine Proposition H.10 and Theorem H.2, and take $\varepsilon=1$ to simplify terms. To conclude, it remains to bound $\sigma_{*}^{2}(\gamma)$ with $\Delta_{*}$. To do this, it is enough to take $x_{i}\in{\rm{argmin}}~{}f_{i}^{\gamma}={\rm{argmin}}~{}f_{i}$ and to use the convexity and smoothness of $f_{i}$ through Lemma A.3 to get
$\|\nabla f_{i}^{\gamma}(x_{*}^{\gamma})\|^{2}=\|\nabla f_{i}^{\gamma}(x_{*}^{\gamma})-\nabla f_{i}^{\gamma}(x_{i})\|^{2}=2L\left(f_{i}^{\gamma}(x_{*}^{\gamma})-f_{i}^{\gamma}(x_{i})\right)=2L\left(f_{i}^{\gamma}(x_{*}^{\gamma})-\inf f_{i}\right).$
After taking expectation
$\sigma_{*}^{2}(\gamma)=2L\left(\inf F^{\gamma}-\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}\right)\leq 2L\left(\inf f-\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}\right)=2L\Delta_{*},$
which yields the wanted bound. ∎
Here we can see that our bounds are far from being as good as the ones in traore_variance_2024. Not only do we have a limitation on the step-size that our competitors do not have, but our constants are worse.
This clearly shows that, while our approach proved to be successful in the interpolation regime, there remains work to be done to properly tackle regularity assumptions on the functions.