### v2

# 3Université Paris Cité and Sorbonne Université, CNRS, Laboratoire de Probabilités, Statistique et Modélisation, F-75013 Paris, France

Abstract
The non-asymptotic analysis of Stochastic Gradient Descent (SGD) typically yields bounds that decompose into a bias term and a variance term. In this work, we focus on the bias component and study the extent to which SGD can match the optimal convergence behavior of deterministic gradient descent. Assuming only (strong) convexity and smoothness of the objective, we derive new bounds that are bias-optimal, in the sense that the bias term coincides with the worst-case rate of gradient descent. Our results hold for the full range of constant step-sizes $\gamma L\in(0,2)$, including critical and large step-size regimes that were previously unexplored without additional variance assumptions. The bounds are obtained through the construction of a simple Lyapunov energy whose monotonicity yields sharp convergence guarantees. To design the parameters of this energy, we employ the Performance Estimation Problem framework, which we also use to provide numerical evidence for the optimality of the associated variance terms.
1 Introduction
Given a family $\{f_{i}\}_{i\in\mathcal{I}}$ of convex and smooth real-valued functions $f_{i}\colon\mathcal{H}\to\mathbb{R}$, we consider the problem given by
$\min_{x\in\mathcal{H}}f(x),\quad\text{where }f\coloneqq\mathbb{E}_{{i\sim\mathcal{D}}}\mathopen{}\left[{f_{i}}\right]\mathclose{},$ ($\mathbb{E}$-minimization)
where $\mathcal{D}$ is a probability distribution over $\mathcal{I}$. We assume $\operatorname{argmin}(f)$ to be nonempty. When the indexing set $\mathcal{I}$ is finite, note that we recover the typical finite sum minimization problem.
The Stochastic Gradient Descent (SGD) algorithm (robbins_stochastic_1951, ) is one of the most popular approaches for solving Problem ($\mathbb{E}$-minimization), and is widely used for large-scale machine learning and stochastic optimization problems. In its standard version, at each iteration, SGD performs a gradient step using one of the functions, which is chosen at random. More precisely, given the iterate $x_{t}$, it picks $i_{t}$ independently and identically distributed according to $\mathcal{D}$, and computes $x_{t+1}$ via
$x_{t+1}=x_{t}-\gamma\nabla f_{i_{t}}(x_{t}),$ (SGD)
where $\gamma>0$ is the step-size. Intuitively, SGD is able to progress towards a solution of the problem because it follows a direction that, on average, is an unbiased estimator for $\nabla f$. The simplicity and efficiency of SGD make it the go-to algorithm for training deep neural networks and other models on massive datasets, where computing the entire gradient $\nabla f$ is intractable or too expensive. While we keep our focus on vanilla SGD, all the results in this paper also apply to mini-batching and non-uniform sampling. This is a simple consequence of the the arguments laid out in gower_sgd_2019 , which we also provide in Appendix C for completeness and for the readers convenience.
Complexity of SGD. Quantitative results on the behavior of SGD consist in upper bounds for a metric $\Delta(T)$, which is to be interpreted as an optimality gap at iteration $T$. Typical choices for such metric include the expected solution gap $\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}$, the expected function value gap $\mathbb{E}\mathopen{}\left[{f(x_{T})-\inf f}\right]\mathclose{}$, the expected average function value gap $\tfrac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\inf f}\right]\mathclose{}$, or the expected ergodic function value gap $\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\inf f}\right]\mathclose{}$, where $\bar{x}_{T}$ is the Cesàro average of the iterates. In most known results garrigos_handbook_2024 , these upper bounds are expressed as the sum of two terms, taking the form
$\Delta(T)\leq{\rm Bias}(T)+{\rm Variance}(T),$
where, typically, the bias term vanishes when $T\to+\infty$ and the variance term can be made arbitrarily small by taking a small enough step-size. For strongly convex and smooth problems, the metric $\Delta(T)$ will typically be the expected solution gap. For convex smooth problems, most proofs construct a bound on the expected average function value gap, and then derive a bound on the expected ergodic function value gap through Jensen’s inequality.
This paper discusses and addresses the following question: what are the best non-asymptotic bounds for SGD? In recent years, remarkable efforts have been made to answer this question for various deterministic optimization methods. These results include the optimal convergence rates of gradient descent (drori_performance_2014, ; taylor_smooth_2017, ; rotaru_exact_2024, ), proximal gradient descent (taylor_exact_2018, ), gradient descent with line-search (de_klerk_worst-case_2017, ), to name a few. In contrast, obtaining an optimal non-asymptotic complexity of stochastic algorithms is significantly more challenging and the optimal complexity of SGD remains nowadays unknown. In this work, we establish new and improved bounds for SGD, that are provably optimal in some sense. These bounds take the form
$\Delta(T)\leq{\rm Bias}(T)\cdot\|x_{0}-x_{*}\|^{2}+{\rm Variance}(T)\cdot\sigma_{*}^{2},$
where $x_{*}$ is a solution of the problem and $\sigma^{2}_{*}=\mathbb{V}[\nabla f_{i}(x_{*})]$ is the solution gradient variance constant (see more details in Section 2). We note that in the deterministic setting, where SGD reduces to GD and where $\sigma^{2}_{*}=0$, such bound reduces to $\Delta(T)\leq\text{Bias}(T)\cdot\|x_{0}-x_{*}\|^{2}$. We will therefore say that a bound is bias-optimal if the bias term matches the optimal convergence rate of GD. In particular, a bias-optimal bound would guarantee that SGD attains an unimprovable convergence rate of the metric $\Delta(T)$ in the interpolation regime (that is when $\sigma^{2}_{*}=0$). We note, however, that it is not clear whether general bounds for SGD always take the form of a gradient descent bound plus an additional variance term. In this work, we solely focus on bounds of the given form.
Contributions. In the convex case, we establish new and improved bounds in expectation for the average function value gap, which automatically convert into bounds for the more standard ergodic function value gap. Our bounds, obtained in Lemmas 4.1, 4.2 and 4.3, with simplified notation for the ease of presentation, are, for arbitrary $\varepsilon\in(0,2)$,
[…]
For short step-sizes $\gamma L\in(0,1)$, our bounds provide better constants than what was previously known. More interesting, we prove that our bound is bias-optimal for both the expected average function value gap and the expected ergodic function value gap. The setting $\gamma L\in[1,2)$ is more delicate to summarize. Our bounds are totally new, in the sense that this setting was never studied before. More importantly, they reveal and suggest non-trivial interactions between the bias and variance terms which, up to our knowledge, were never observed before. We now split our discussion between the critical and large step-sizes.
For large step-sizes $\gamma L\in(1,2)$, we prove that our bound is bias-optimal for the expected average function value gap. A striking feature of this bound is that the variance term has an impractical exponential grow with time, which is a phenomenon previously unobserved. We were not able to prove formally that this variance term cannot be improved, but our numerical experiments (see more details in the dedicated section) suggest that one cannot avoid a variance term diverging with time whilst keeping an optimal bias term. However, if we allow for a slightly worse bias term, hence deviating from the bias-optimal setting, this variance may be uniformly bounded with respect to $T$ (see Lemma 4.4 for more details). Of course, those considerations about the variance are irrelevant in the interpolation regime, where $\sigma_{*}^{2}=0$.
Regarding the critical step-size $\gamma L=1$, it is natural to conjecture that the optimal bias should be $\tfrac{L}{2T}$, which is the limiting optimal bias as $\gamma L\to 1$. This constant $\tfrac{L}{2T}$ plays a particular role, because it is also the infimum of the optimal biases for $\gamma L\in(0,2)$, and is also the best constant for GD. We prove that it is possible to obtain a bound for which the bias term is arbitrarily close to $\tfrac{L}{2T}$, but that it comes at the cost of a constant variance term which diverges. We moreover conjecture that it is not possible to derive a bound with a bias term equal to $\tfrac{L}{2T}$ when $\gamma L=1$, which we support by numerical evidence.
We highlight that the study of the critical step-size $\gamma L=1$ in the convex smooth setting is novel. This in turn unlocks the study of the stochastic proximal algorithm, by viewing it as an instance of SGD applied to regularized functions. To the best of our knowledge, we establish the first complexity guarantees for this method in a general convex nonsmooth setting, considering possibly functions which do not have a full domain. This is presented in Section 6.
Let us now turn to the smooth strongly convex case, where each function $f_{i}$ is $L$-smooth and $\mu$-strongly convex. Our performance metric is the expected distance to the solution, namely $\Delta(T)=\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}$. Conceptually, our results are of the same nature as in the convex setting. Firstly, we obtain bounds for the full range of step-sizes $\gamma L\in(0,2)$, while the existing literature (bach_non-asymptotic_2011, ; needell_stochastic_2016, ; gower_sgd_2019, ) considered at best $\gamma L<1$. Extending the range of valid step-sizes allows us to consider the critical step-size $\gamma=\tfrac{2}{\mu+L}>1/L$, which is known to be optimal in the deterministic setting. Second, we show that our bounds are bias-optimal, except for the critical step-size for which we observe again a singularity. For every non-critical step-size we obtain ${\rm Bias}(T)=\phi^{2T}$, where $\phi$ defines the optimal constant for GD, namely
$\phi=\max\{1-\gamma\mu;\gamma L-1\}.$
For the critical step-size, we show that $\phi^{2T}$ can be approached with an arbitrary tolerance, with the cost of a variance growing to infinity. Supported by numerical results, we again conjecture that this optimal bias cannot be obtained.
Proof Strategy. For each fixed time horizon $T$, we define a time-dependent random energy sequence
$E_{t}=a_{t}\cdot\|x_{t}-x_{*}\|^{2}+\rho\cdot\sum_{s=0}^{t-1}\big(f(x_{s})-\min f\big)-\sum_{s=0}^{t-1}e_{s}\sigma_{*}^{2}\quad\forall t=0,\ldots,T,$ (Lyapunov)
where $\rho$, $(a_{t})$ and $(e_{t})$ are nonnegative parameters. The first term in the Lyapunov energy is the distance to the solution $\|x_{t}-x_{*}\|^{2}$, a classical term which decreases for deterministic gradient dynamics. The second term involves the function gap $f(x_{t})-\min f$, which also typically decreases for gradient descent. We chose to replace the standard term $t(f(x_{t})-\min f)$ with the sum of the past function gaps, which is of the same order in time. Up to our knowledge, this choice is new in the context of stochastic optimization, and we believe this choice to be critical to be able to derive bias-optimal bounds. The third term is a negative cumulated sum, introduced to compensate the fluctuations caused by the variance in the SGD algorithm, and hence to allow the Lyapunov energy to decrease.
A decrease of $\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ along the iterations implies that $\mathbb{E}\mathopen{}\left[{E_{T}}\right]\mathclose{}\leq E_{0}$, which translates into a bound for SGD (see Section 3.1 for more details). Our strategy is therefore to identify admissible Lyapunov parameters, which are those that make the energy decrease in expectation. Among those admissible parameters, our goal is to find the ones minimizing, whenever possible, the bias term in our bound.
To this end, we rely on the approach followed in taylor_stochastic_2019 (see also lessard_analysis_2016 ; fercoq_defining_2024 ; upadhyaya_automated_2025 ) to study Lyapunov potentials in a systematic fashion. It is based on the Performance Estimation Problem (PEP) methodology, which originally appeared in drori_performance_2014 , and was further developed in taylor_smooth_2017 . The PEP framework allows to find numerically admissible Lyapunov parameters, but also the ones minimizing the corresponding bias, by reformulating such problem into a semidefinite program. It helped us in two ways. First it helped us to navigate the space of admissible Lyapunov parameters and to obtain initial guesses for some parameter sequences, which eventually led us to our theoretical proofs. The reader curious about this process may check the proofs in the appendix for a more in-depth description. Second, it allowed us to numerically assess certain conjectures, such as the bias-optimality of our bounds, or the singularity of critical step-sizes.
Related Work. In early studies of SGD in the convex (smooth) framework, it was standard to make strong assumptions on the stochastic gradients $\nabla f_{i_{t}}(x_{t})$. The oldest, and possibly the best-known, are the uniform boundedness of the variance, and its variant, the uniform boundedness of the gradients, namely (robbins_stochastic_1951, ; nemirovski_robust_2009, ; rakhlin_making_2012, ; schmidt_fast_2013, )
$\sup_{x\in\mathcal{H}}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x)-\nabla f(x)\|^{2}}\right]\mathclose{}<+\infty\quad\text{or}\quad\sup_{x\in\mathcal{H}}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x)\|^{2}}\right]\mathclose{}<+\infty.$
Such assumptions date back to the early works of the 1950s (chung_stochastic_1954, ; sacks_asymptotic_1958, ) and continued to be employed until recently (polyak_acceleration_1992a, ; nemirovski_robust_2009, ). A culminating contribution to the nonasymptotic analysis of SGD under bounded variance assumptions is taylor_stochastic_2019 , where new and sharp convergence rates were established. Over time, these variance assumptions have been progressively relaxed and replaced by conditions that control the gradient or its variance through quantities such as $\|x\|$ or $\|\nabla f(x)\|$, we refer to the overview in alacaoglu_towards_2025 and the references therein. Nevertheless, variance-based assumptions can be unrealistic in practice (bottou_optimization_2018, ; nguyen_sgd_2018, ). Moreover, uniform boundedness of the gradients can never be satisfied for strongly convex functions.
A new line of research was initiated in bach_non-asymptotic_2011 by removing any such variance assumptions, followed by the more recent works (needell_stochastic_2016, ; nguyen_sgd_2018, ; gower_sgd_2019, ; gower_sgd_2021, ; khaled_better_2023, ; demidovich_guide_2023, ). The core idea is to exploit the convexity and smoothness to obtain a variance transfer inequality, bounding the variance at any point with a term involving $\sigma^{2}_{*}$. The new gold standard for the analysis of SGD quickly became to not make any variance assumption when assuming convexity and smoothness, a framework into which our work fits.
We point out that convexity and smoothness is a sufficient but not a necessary condition to obtain complexity results: all one needs is for some variance transfer inequality to hold. We mention for instance the expected smoothness assumption (gower_stochastic_2021, ; gower_sgd_2019, ; khaled_better_2023, ) and the ABC condition (demidovich_guide_2023, ), which were shown to be sufficient to establish performance guarantees. These assumptions are strictly weaker than convexity and smoothness, enabling the analysis of SGD in a nonconvex framework. While our analysis could also be carried out under weaker assumptions than convexity and smoothness, it is unclear how practical such assumptions would be, and we do not delve further into this in this work.
Note on Concurrent Work. In the convex setting, our work focuses on rates on the ergodic function value gap $f(\bar{x}_{T})-\min f$, which is currently the standard for SGD without uniformly bounded variance. In contrast, when making a stronger assumption on the stochastic gradients, it is standard to obtain rates on the last-iterate function value gap bach_non-asymptotic_2011 ; taylor_stochastic_2019 ; liu_revisiting_2023 . At the time of writing this document, it was unclear whether SGD without stronger variance assumptions could enjoy last-iterate rates, hence justifying our choice of performance metric. During the review phase of the manuscript, two independent works have answered this question positively attia_fast_2025 ; garrigos_lastiterate_2025 , opening the door for the possibility to extend our results to the setting of last-iterate rates.
Paper Organization. This paper is organized as follows: We present the formal problem and the assumptions in Section 2. In Section 3, we outline the proof format of our main results. Sections 4 and 5 are devoted to our results (new bounds, most of which being bias-optimal) in the convex and strongly convex settings, respectively. In Section 6 we present the application of our results to the Stochastic Proximal algorithm. Section 7 discusses our PEP framework and the numerical optimality of our variance terms. The technical proofs are gathered in the appendices.
2 Preliminaries
We first formalize the problem statement presented in Problem ($\mathbb{E}$-minimization).
Problem 1(Main problem)
Let $\mathcal{H}$ be a real Hilbert space with associated inner product $\langle\cdot,\cdot\rangle$ and induced norm $\|\cdot\|$. Let $\{f_{i}\}_{i\in\mathcal{I}}$ be a family of real-valued functions $f_{i}\colon\mathcal{H}\to\mathbb{R}$, where $\mathcal{I}$ is a (possibly infinite) set of indices. We consider the problem of minimizing $f\coloneqq\mathbb{E}\mathopen{}\left[{f_{i}}\right]\mathclose{}$, where the expectation is taken over the indices $i\in\mathcal{I}$, with respect to some probability distribution $\mathcal{D}$ over $\mathcal{I}$. We assume that
  1. 1.
    the problem is well-defined, in the sense that $i\mapsto f_{i}(x)$ is $\mathcal{D}$-measurable, and that $\mathbb{E}\mathopen{}\left[{f_{i}(x)}\right]\mathclose{}$ is finite for every $x\in\mathcal{H}$;
  2. 2.
    the problem is well-posed, in the sense that ${\rm{argmin}}~f\neq\emptyset$;
  3. 3.
    the problem is differentiable, in the sense that each $f_{i}$ is differentiable, and so is $f$, with $\nabla f(x)=\mathbb{E}\mathopen{}\left[{\nabla f_{i}(x)}\right]\mathclose{}$.
Within the context of Problem 1 the (SGD) algorithm is well defined. To analyze its performance, we will further assume the problem to be convex and smooth. We define a setting combining both convexity and strong convexity, which allows us to unify parts of the analyses.
{assumption}
[Smoothness and convexity] Considering Problem 1, we assume that
  • $\bullet$
    there exists $L\in(0,+\infty)$ such that each $f_{i}$ is $L$-smooth, i.e. $\nabla f_{i}\colon\mathcal{H}\to\mathcal{H}$ is $L$-Lipschitz continuous;
  • $\bullet$
    there exists $\mu\in[0,+\infty)$ such that each function $f_{i}$ is $\mu$-strongly convex. When $\mu=0$, this means that $f_{i}$ is convex.
Smoothness and strong convexity can be characterized by the means of a variational inequality, which will be central to the derivation of our results.
Lemma 1(Expected cocoercivity)
Let Assumption 1 hold. Then, for every $(x,y)\in\mathcal{H}^{2}$,
$\frac{1}{2}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(y)-\nabla f_{i}(x)\|^{2}}\right]\mathclose{}+\frac{\mu L}{2}\|y-x\|^{2}\leq(L-\mu)\left(f(y)-f(x)\right)+\langle\mu\nabla f(y)-L\nabla f(x),y-x\rangle.$ (EC)
Proof
Because each $f_{i}$ is $L$-smooth and $\mu$-strongly convex, we can use (taylor_smooth_2017, , Theorem 4) and take an expectation over $i$.
We now turn to the constant $\sigma_{*}^{2}$ which controls the complexity of the SGD algorithm.
{assumption}
[Finite solution variance] Considering Problem 1, we assume that the variance at the solution exists, namely that $\sigma^{2}_{*}\coloneqq\inf\limits_{x_{*}\in{\rm{argmin}}~f}\mathbb{E}\left[\|\nabla f_{i}(x_{*})\|^{2}\right]<+\infty$.
We note that, under Assumption 1, the variance term $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$ does not depend on the choice of $x_{*}\in\operatorname{argmin}f$, as was already observed in (garrigos_handbook_2024, , Lemma 4.17).
Lemma 2
Let Assumption 1 hold. Then $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$ is constant over ${\rm{argmin}}~f$.
Proof
Let $x_{*},x_{*}^{\prime}$ be two minimizers of $f$. From Lemma 1 we know that (EC) holds true for the couple $(x_{*},x_{*}^{\prime})$, which means that
$\frac{1}{2}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*}^{\prime})-\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}\leq(L-\mu)\left(f(x_{*}^{\prime})-f(x_{*})\right)+\langle\mu\nabla f(x_{*}^{\prime})-L\nabla f(x_{*}^{\prime}),y-x\rangle.$
Because $f(x_{*}^{\prime})=f(x_{*})=\inf f$ and $\nabla f(x_{*}^{\prime})=\nabla f(x_{*})=0$, we deduce that
$\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*}^{\prime})-\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}=0.$
Then, almost surely with respect to the distribution $\mathcal{D}$ over the indices $i\in\mathcal{I}$, we have $\|\nabla f_{i}(x_{*}^{\prime})-\nabla f_{i}(x_{*})\|^{2}=0$. This means that $\nabla f_{i}(x_{*}^{\prime})=\nabla f_{i}(x_{*})$ almost surely, from which we conclude that $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*}^{\prime})\|^{2}}\right]\mathclose{}=\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$.
Assuming that the variance at the solution $\sigma_{*}^{2}$ exists is trivially satisfied in the case of the minimization of a finite sum.
Lemma 3
If Problem 1 treats a finite sum of functions, i.e. $\mathcal{I}$ is finite, then Assumption 2 holds true.
Even for true expectation-minimization problems, where $\mathcal{I}$ is infinite, Assumption 2 is very mild. In particular, this assumption is automatically verified in problems where the loss functions $f_{i}$ are nonnegative, which is standard for problems arising in inverse problems and machine learning.
Lemma 4
Consider Problem 1 and assume that the functions $f_{i}$ are $L$-smooth and are uniformly bounded from below functions: there exists $r\in\mathbb{R}$ such that $f_{i}(x)\geq r$ for every $i\in\mathcal{I}$ and $x\in\mathcal{H}$. Then Assumption 2 is true, with moreover the information that
$\sigma_{*}^{2}\leq 2L(\min f-r).$
Proof
Since each $f_{i}$ is $L$-smooth, we can write (garrigos_handbook_2024, , Lemma 2.28), for all $x\in\mathcal{H}$,
$\|\nabla f_{i}(x)\|^{2}\leq 2L(f_{i}(x)-\min f_{i})\leq 2L(f_{i}(x)-r).$
Setting $x=x_{*}\in{\rm{argmin}}~f$ and taking the expectation of this inequality, we obtain
$\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}\leq 2L(f(x_{*})-r)=2L(\min f-r)<+\infty,$
which concludes since $\sigma_{*}^{2}=\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$.
We now highlight the connection between the variance at the solution $\sigma^{2}_{*}$ with the notion of interpolation. The following facts can be found in garrigos_handbook_2024 , which we have extended here from a finite-sum problem to our expectation-minimization Problem 1.
Definition 1(Interpolation)
Consider Problem 1. We say that interpolation holds for the family of functions $\{f_{i}\}_{i\in\mathcal{I}}$ if there exists $x_{*}$ which belongs to ${\rm{argmin}}~f_{i}$ for $\mathcal{D}$-almost every $i$.
Lemma 5
In the context of Problem 1, if interpolation holds, then $\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}=\inf f$ and $\cap_{i\in\mathcal{I}}\ {\rm{argmin}}~f_{i}={\rm{argmin}}~f$, where the intersection is to be understood for $\mathcal{D}$-almost every $i\in\mathcal{I}$.
Proof
Because interpolation holds, we may select $x_{*}\in\cap_{i\in\mathcal{I}}\ {\rm{argmin}}~f_{i}$. Let us start by showing that $x_{*}\in{\rm{argmin}}~f$. For this, consider any $x\in\mathcal{H}$ and write
$f(x_{*})=\mathbb{E}\mathopen{}\left[{f_{i}(x_{*})}\right]\mathclose{}=\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{f_{i}(x)}\right]\mathclose{}=f(x).$
This proves that $x_{*}\in{\rm{argmin}}~f$. Now let us prove the first part, by observing that
$\inf f=f(x_{*})=\mathbb{E}\mathopen{}\left[{f_{i}(x_{*})}\right]\mathclose{}=\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}.$
Secondly, we prove the second point, for which we are only left with the reverse inclusion. Consider some $x\in{\rm{argmin}}~f$. Then
$f(x)=\inf f=\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}\ \implies\ \mathbb{E}\mathopen{}\left[{f_{i}(x)-\inf f_{i}}\right]\mathclose{}=0.$
Since $f_{i}(x)-\inf f_{i}\geq 0$ for all $i$, it must hold that $f_{i}(x)-\inf f_{i}=0$ for almost every $i\in\mathcal{I}$, which proves the claim.
Lemma 6
In the context of Problem 1, if interpolation holds, then $\sigma_{*}^{2}=0$. This becomes an equivalence if all the functions $f_{i}$ are convex.
Proof
If interpolation holds, then for every $x_{*}\in{\rm{argmin}}~f$ we have that $x_{*}\in{\rm{argmin}}~f_{i}$ for almost every $i\in\mathcal{I}$. For such $i\in\mathcal{I}$, the optimality condition gives $\nabla f_{i}(x_{*})=0$, which after taking the norm and expectation leads to $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}=0$.
Reciprocally, if we assume that $\sigma_{*}^{2}=0$, it means that there exists $x_{*}\in{\rm{argmin}}~f$ such that $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}=0$. This is an expectation of nonnegative terms, so we deduce that $\nabla f_{i}(x_{*})=0$ for almost every $i\in\mathcal{I}$. Exploiting our convexity assumption, we deduce that $x_{*}$ is a minimizer of $f_{i}$ for almost every $i\in\mathcal{I}$, which shows that interpolation holds.
3 Reducing a Lyapunov Analysis to the Resolution of a System of Inequalities
Our main approach to derive bounds for SGD is to rely on an analysis of a new parametrized family of Lyapunov energies. After introducing these Lyapunov energies, we will show how to deduce bounds for SGD from a Lyapunov decrease in Section 3.1. In a second time, we present our main technical result which states that obtaining such Lyapunov decrease can be reduced to solving a system of inequalities (see Section 3.2).
3.1 Bounds for SGD from a Lyapunov Decrease
Let $T\geq 1$ be fixed, let $x_{*}\in{\rm{argmin}}~f$, and let $(x_{t})_{t=0}^{T}$ be generated by the SGD algorithm for some fixed step-size $\gamma>0$. Given a set of parameters $\rho,a_{0},\dots,a_{T},e_{0},\dots,e_{T-1}\geq 0$, we define the following Lyapunov energy, for $t=0,\dots,T-1$:
$E_{t}\coloneqq a_{t}\|x_{t}-x_{*}\|^{2}+\rho\sum_{s=0}^{t-1}\left(f(x_{s})-\min f\right)-\sum_{s=0}^{t-1}e_{s}\sigma_{*}^{2},$
where, by convention, the empty sum $\sum_{s=0}^{-1}$ is equal to zero.
The first term of this Lyapunov energy is the distance to the solution $\|x_{t}-x_{*}\|^{2}$, a classical term which typically decreases for deterministic monotone gradient dynamics. The second term involves the function gap $f(x_{t})-\min f$, which also typically decreases for gradient descent. The standard Lyapunov for gradient descent usually contains the term $t(f(x_{t})-\min f)$, and such term was also considered for analyzing SGD in taylor_stochastic_2019 . Here we make a slightly different choice, replacing this term with the sum of the past function gaps (observe that both are of the same order in time). The last term for this Lyapunov energy is a negative cumulated sum, where the $e_{t}$’s measure the contribution of the variance. This term is meant to compensate the fluctuations caused by the uncertainty in the SGD algorithm, and will allow the Lyapunov energy to decrease.
A first result consists in deriving bounds on the average function value gap, provided the Lyapunov parameter $\rho$ is nonzero. This bound will typically be obtained for convex smooth problems. The proof is based on a telescopic sum and an application of Jensen’s inequality, as is often the case for convex smooth objectives, as done for instance in taylor_stochastic_2019 ; gower_sgd_2021 .
Lemma 7(Bound from Lyapunov decrease. Convex case)
Consider Problem 1, assume that $f$ is convex, and let $T\geq 1$. Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for every $t=0,\dots,T-1$, and that $\rho>0$. Then, if we note $\bar{e}=\tfrac{1}{T}\sum_{t=0}^{T-1}e_{t}$ and $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\min f}\right]\mathclose{}\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\bar{e}\sigma_{*}^{2}}{\rho}.$
Proof
Iterating $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for $t=0,\dots,T-1$ yields that $\mathbb{E}\mathopen{}\left[{E_{T}}\right]\mathclose{}\leq E_{0}$, or, equivalently,
$a_{T}\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}+\rho\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\min f}\right]\mathclose{}-\sum_{t=0}^{T-1}e_{t}\sigma_{*}^{2}\leq a_{0}\|x_{0}-x_{*}\|^{2}.$
Bounding $a_{T}\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\geq 0$ and dividing by $\rho T>0$ yields
$\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]\leq\frac{1}{T}\sum_{t=0}^{T-1}[f(x_{t})-\min f]\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{\rho T}+\frac{\bar{e}\sigma_{*}^{2}}{\rho},$
where in the first inequality we used Jensen’s inequality.
As a second result, a Lyapunov decrease can yield a bound of the square distance of the iterate $x_{t}$ to a minimizer, provided the Lyapunov parameter $a_{T}$ is nonzero. This is standard for strongly convex problems, as we will see in Section 5. The proof, also based on a telescopic argument, is again typical for this class of problems, see for instance gower_sgd_2019 .
Lemma 8(Bound from Lyapunov decrease. Strongly convex case)
Consider Problem 1, and let $T\geq 1$. Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for every $t=0,\dots,T-1$, and that $a_{T}>0$. Then, if we note $e_{T}^{sum}=\sum_{t=0}^{T-1}e_{t}$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\frac{a_{0}\|x_{0}-x_{*}\|^{2}}{a_{T}}+\frac{e_{T}^{sum}\sigma_{*}^{2}}{a_{T}}.$
Proof
Iterating $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for $t=0,\ldots,T-1$ yields $\mathbb{E}\mathopen{}\left[{E_{T}}\right]\mathclose{}\leq E_{0}$, or in other words that
$a_{T}\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}+\rho\sum_{t=0}^{T-1}\mathbb{E}\mathopen{}\left[{f(x_{t})-\min f}\right]\mathclose{}\leq a_{0}\|x_{0}-x_{*}\|^{2}+e_{T}^{sum}\sigma_{*}^{2}.$
Since $f(x_{t})-\min f\geq 0$ for all $t=0,\ldots,T-1$, and $\rho\geq 0$, this yields the wanted inequality after division by $a_{T}>0$.
3.2 Lyapunov Decrease from a System of Inequalities
Lemmas 7 and 8 reduce the problem of finding upper bounds for SGD to finding parameters $\rho,a_{0},\dots,a_{T},e_{0},\dots,e_{T-1}\geq 0$ that allow the Lyapunov energy to decrease. The next theorem provides sufficient conditions for this to be the case, and will be at the heart of our subsequent results. Indeed, this will allow to shift the focus from sequences generated by SGD to a system of inequalities on real numbers.
Theorem 3.1(Sufficient conditions for Lyapunov decrease)
Let Assumptions 1 and 2 hold true, and let $T\geq 1$. Assume that there exist parameters $(\alpha_{t},\beta_{t})_{t=0}^{T-1}$ such that, for every $t=0,\dots,T-1$, the following conditions are verified:
  1. 1.
    $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}\geq 0$,
  2. 2.
    $\rho\leq 2(L-\mu)(\alpha_{t}-\beta_{t})$,
  3. 3.
    $a_{t+1}\leq\mu L(\alpha_{t}+\beta_{t})+a_{t}$,
  4. 4.
    $a_{t+1}\gamma^{2}\leq\alpha_{t}+\beta_{t}$,
  5. 5.
    $(a_{t+1}\gamma-\alpha_{t}L-\beta_{t}\mu)^{2}\leq(\mu L(\alpha_{t}+\beta_{t})+a_{t}-a_{t+1})(\alpha_{t}+\beta_{t}-a_{t+1}\gamma^{2})$,
  6. 6.
    $a_{t+1}\gamma^{2}(\alpha_{t}+\beta_{t})\leq(\alpha_{t}+\beta_{t}-a_{t+1}\gamma^{2})e_{t}$.
Then $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for every $t=0,\ldots,T-1$.
Proof
Using the definition of $x_{t+1}$ allows us to write
$E_{t+1}-E_{t}=\rho\left(f(x_{t})-\min f\right)-e_{t}\sigma_{*}^{2}+(a_{t+1}-a_{t})\|x_{t}-x_{*}\|^{2}+a_{t+1}\gamma^{2}\|\nabla f_{i}(x_{t})\|^{2}-2a_{t+1}\gamma\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle.$ (1)
In what follows, we denote by $\mathbb{E}_{t}\coloneqq\mathbb{E}\mathopen{}\left[{\cdot\ |\ x_{0},\dots,x_{t}}\right]\mathclose{}$ the conditional expectation with respect to the iterates up to and including $x_{t}$. Multiplying the expected cocoercivity inequalities (Lemma 1) for the pair $(x_{t},x_{*})$ by $2\alpha_{t}\geq 0$ and for the pair $(x_{*},x_{t})$ by $2\beta_{t}\geq 0$, and summing them gives
$({\alpha_{t}+\beta_{t}})\mathbb{E}_{t}\mathopen{}\left[{\|\nabla f_{i}(x_{t})-\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}+{\mu L(\alpha_{t}+\beta_{t})}\|x_{t}-x_{*}\|^{2}$
$\leq$ $2(L-\mu)(\beta_{t}-\alpha_{t})\left(f(x_{t})-\min f\right)+2(\alpha_{t}L+\beta_{t}\mu)\langle\nabla f(x_{t}),x_{t}-x_{*}\rangle.$
Adding $\mathbb{E}_{t}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}-E_{t}$ on both sides and using the expression obtained in (1), taking the expectation and developing the squares yields
$\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}$ $\leq$ $(\rho+2(L-\mu)(\beta_{t}-\alpha_{t}))\mathbb{E}\left[f(x_{t})-\min f\right]$
$+\mathbb{E}\mathopen{}\left[{A{\|x_{t}-x_{*}\|^{2}}+B{\|\nabla f_{i}(x_{t})\|^{2}}+B^{\prime}{\|\nabla f_{i}(x_{*})\|^{2}}+2C{\langle\nabla f_{i}(x_{t}),x_{t}-x_{*}\rangle}+2D{\langle\nabla f_{i}(x_{t}),\nabla f_{i}(x_{*})\rangle,}}\right]\mathclose{}$
where we simplified the expression by introducing the following constants
$\begin{cases}A&=\mu L(\alpha_{t}+\beta_{t})+a_{t}-a_{t+1},\\ B&=\alpha_{t}+\beta_{t}-a_{t+1}\gamma^{2},\\ B^{\prime}&=e_{t}+\alpha_{t}+\beta_{t},\\ C&=a_{t+1}\gamma-\alpha_{t}L-\beta_{t}\mu,\\ D&=-(\alpha_{t}+\beta_{t}).\end{cases}$
For the Lyapunov energy to be decreasing in expectation, it is enough that the right-hand side is nonpositive. For the first term, because $f(x_{t})-\min f\geq 0$, it is enough to assume that $\rho\leq 2(L-\mu)(\alpha_{t}-\beta_{t})$, which corresponds to Condition 2. The remaining terms are simply the expectation of a quadratic polynomial in $X=x_{t}-x_{*}$, $Y=\nabla f_{i}(x_{t})$ and $Y^{\prime}=\nabla f_{i}(x_{*})$. Using elementary linear algebra, it can be shown (see Lemma 9 below) that for this polynomial to be nonnegative, it is enough to require
$A,B,B^{\prime}\geq 0,\quad C^{2}\leq AB,\quad\text{and}\quad D^{2}\leq BB^{\prime}.$
A simple calculation shows that these conditions correspond exactly to the remaining conditions of our theorem.
Remark 1
We note that Theorem 3.1, which is the key theorem for all of our subsequent results, holds provided that the inequality of Lemma 1 holds for any pair of points $(x,x_{*})$ and $(x_{*},x)$, where $x\in\mathcal{H}$. As such, we could replace Assumption 1 by such a weaker version, which includes a class of nonconvex functions. Moreover, in the case where $\beta_{t}\equiv 0$, it is only necessary for the conclusion of Lemma 1 to hold for any pair $(x,x_{*})$, where $x\in\mathcal{H}$. We note the similarities with the recently introduced expected smoothness assumption gower_stochastic_2021 . Although this nonconvex extension is of interest, we should not delve further into it in this work.
We end this section by proving the technical lemma used the previous proof.
Lemma 9(Nonnegative quadratic polynomial)
Let $A,B,B^{\prime},C,D\in\mathbb{R}$ be such that
$A,B,B^{\prime}\geq 0,\quad C^{2}\leq AB,\quad\text{and}\quad D^{2}\leq BB^{\prime}.$ (2)
Let $X,Y,$ and $Y^{\prime}$ be three random variables over $\mathcal{H}$ such that $\mathbb{E}\left[Y^{\prime}\right]=0$. Then
$\mathbb{E}\mathopen{}\left[{A\|X\|^{2}+B\|Y\|^{2}+B^{\prime}\|Y^{\prime}\|^{2}+2C\langle Y,X\rangle-2D\langle Y,Y^{\prime}\rangle}\right]\mathclose{}\geq 0.$
Proof
Let us note $P$ to be the quantity that we want to see nonnegative. We start by exploiting the fact that $\mathbb{E}\mathopen{}\left[{Y^{\prime}}\right]\mathclose{}=0$ to write that $P=P+\theta\mathbb{E}\mathopen{}\left[{\langle X,Y^{\prime}\rangle}\right]\mathclose{}$, for every $\theta\in\mathbb{R}$. Now, see that for $P\geq 0$ to be true, it is enough for
$\hat{P}_{\theta}(x,y,y^{\prime})=A\|x\|^{2}+B\|y\|^{2}+B^{\prime}\|y^{\prime}\|^{2}+2C\langle y,x\rangle-2D\langle y,y^{\prime}\rangle+2\theta\langle x,y^{\prime}\rangle$
to be nonnegative for every $x,y,y^{\prime}\in\mathcal{H}$ and for some $\theta\in\mathbb{R}$. Our polynomial $\hat{P}_{\theta}$ is equal to $\langle M_{\theta}z,z\rangle$ where $z=(x,y,y^{\prime})$ and
$M_{\theta}=\begin{pmatrix}A&C&\theta\\ C&B&-D\\ \theta&-D&B^{\prime}\end{pmatrix}$.
As such, $\hat{P}_{\theta}\geq 0$ if, and only if, $M_{\theta}\succeq 0$, which, by Sylvester’s criterion, is equivalent to
$A,B,B^{\prime}\geq 0,$ (3a)
$C^{2}\leq AB,\quad\theta^{2}\leq AB^{\prime},\quad D^{2}\leq BB^{\prime},$ (3b)
$A(BB^{\prime}-D^{2})\geq B^{\prime}C^{2}+B\theta^{2}+2CD\theta.$ (3c)
So $P\geq 0$ if there exists $\theta\in\mathbb{R}$ such that Inequalities (3) are verified. We can already see Equations (3a)-(3b) imply Equations (2). Let us prove that in fact the existence of a $\theta\in\mathbb{R}$ satisfying Equations (3) is equivalent to Equations (2).
To prove equivalence, we assume Equations (2) and show that there exists a $\theta\in\mathbb{R}$ such that
$AB^{\prime}\geq\theta^{2}\quad\text{ and }\quad A(BB^{\prime}-D^{2})\geq B^{\prime}C^{2}+B\theta^{2}+2CD\theta.$ (4)
The last inequality is quadratic in $\theta$, and has solutions if, and only if, its discriminant $\Delta$ is nonnegative. A straightforward computation shows that
$\Delta=4(BB^{\prime}-D^{2})(AB-C^{2}),$
which guarantees that $\Delta\geq 0$ given Equations (2). The quadratic may hence be rewritten as
$\frac{-2CD-\sqrt{\Delta}}{2B}\leq\theta\leq\frac{-2CD+\sqrt{\Delta}}{2B}.$
For (4) to have a solution, the above must hold simultaneously with $\sqrt{AB^{\prime}}\leq\theta\leq\sqrt{AB^{\prime}}$. This is equivalent to the intersection of two intervals to be nonempty, which is the case if
$\begin{cases}\frac{-2CD+\sqrt{\Delta}}{2B}\geq-\sqrt{AB^{\prime}}\quad&\text{if $CD\geq 0$},\\ \frac{-2CD-\sqrt{\Delta}}{2B}\leq\sqrt{AB^{\prime}}\quad&\text{if $CD\leq 0$}.\end{cases}$
or equivalently,
$2|CD|\leq 2B\sqrt{AB^{\prime}}+\sqrt{\Delta}.$
As all terms are nonnegative, this inequality is equivalent to
$4C^{2}D^{2}\leq 4B^{2}AB^{\prime}+4B\sqrt{AB^{\prime}\Delta}+\Delta,$
which holds true since $D^{2}\leq BB^{\prime}$, $C^{2}\leq AB$ and $\Delta\geq 0$.
4 Bias-Optimal Bounds in the Convex Smooth Setting
We now present bounds for SGD that follow directly from Lemma 7. These bounds are obtained by identifying specific Lyapunov parameters that satisfy the sufficient conditions of Theorem 3.1, with the objective of maximizing the value of $\rho$, which controls the bias term in Lemma 7. Our approach proceeds in two stages. First, using numerical tools (see Section 7 for further details), we determined the maximal admissible value of $\rho$. In a second time, we established a formal proof demonstrating that this value indeed yields an optimal bias term. We note that the Lyapunov parameters $e_{t}$ arising from these proofs are at times rather intricate. For clarity of exposition, we therefore present bounds based on simplified parameter choices in this section. The complete, more detailed bounds, as well as the full proofs and the underlying motivation for the parameter selection, are deferred to Appendix A.1.
The presentation of our results is split into three parts based on the normalized step-size $\gamma L$, namely Theorem 4.1 for short step-sizes $\gamma L\in(0,1)$, Theorem 4.2 for the critical step-size $\gamma L=1$, and Theorem 4.3 for large step-sizes $\gamma L\in(1,2)$.
4.1 Short Step-Sizes $\gamma L<1$
Theorem 4.1(Convex case, short step-sizes)
Let Assumptions 1 and 2 hold true with $\mu=0$, and let $(x_{t})$ be generated by (SGD) with $\gamma L\in(0,1)$. Then, for every $T\geq 1$, with $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[f(x_{t})-\min f\right]\leq\frac{L\|x_{0}-x_{*}\|^{2}}{2\gamma LT+2(1-\gamma L)}+\frac{\gamma\sigma_{*}^{2}}{2(1-\gamma L)}.$
Moreover, the above bound is bias-optimal, in the sense that there exists a problem for which $\sigma_{*}^{2}=0$ and both inequalities are equalities.
Proof
The first inequality follows from Jensen’s inequality. The second follows from Lemma 7 and Theorem 3.1, where in the latter we select the following parameters:
$\rho=2\gamma+\frac{2(1-\gamma L)}{LT},\quad a_{t}=\frac{T-t}{T}\frac{1+\gamma L(T-1)}{1+\gamma L(T-t-1)},\quad e_{t}=\frac{\gamma^{2}}{1-\gamma L},\quad\alpha_{t}\equiv\alpha=\frac{\rho}{2L},\quad\beta_{t}\equiv\beta=0.$
A detailed proof of the validity of those parameters, including the motivation for their choice, is available in Appendix A.1.1. Regarding the bias-optimality, we consider the convex and $L$-smooth Huber function, parameterized by $\eta>0$,
$\mathcal{H}_{\eta}(x)=\begin{cases}\eta L\|x\|-\frac{L}{2}\eta^{2}&\text{if}\quad\|x\|>\eta\\ \frac{L}{2}\|x\|^{2}&\text{if}\quad\|x\|\leq\eta.\end{cases}$
When $\eta$ is sufficiently small, namely $\eta\leq\frac{\|x_{0}\|}{1+(T-1)\gamma L}$, it can be shown that the first $T$ iterates remain within the region where $\mathcal{H}_{\eta}$ is linear. In this setting, a direct computation yields
$\mathcal{H}_{\eta}(\overline{x}_{T})-\inf\mathcal{H}_{\eta}=\frac{L\|x_{0}-x^{*}\|^{2}}{2(1+(T-1)\gamma L)},$
which matches the bias term of Theorem 4.1. The full argument is provided in the proof of Proposition 3.
Remark 2(Related works)
Theorem 4.1 provides a better bound, both in terms of bias and variance, than what can be found in the literature, while also extending the range of step-sizes for which the bounds apply. For instance, (gower_sgd_2021, , Theorem D.6) requires $\gamma L\in(0,\frac{1}{2})$ and derives the bound
$f(\bar{x}_{T})-\inf f\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma(1-2\gamma L)T}+\frac{\gamma}{1-2\gamma L}\sigma_{*}^{2}.$
Note that the bias term blows up when $\gamma L$ approaches the upper bound $\tfrac{1}{2}$, which is not the case for our bound when $\gamma L\to 1$. However, they rely on weaker assumptions than ours, namely the smoothness of $f$ (instead of each $f_{i}$) and the quasar-convexity of the $f_{i}$ (instead of the convexity), and cover the use of variable step-size, which we do not address in this work, see also (garrigos_handbook_2024, , Theorem 5.3). A less straightforward comparison can be made with (taylor_stochastic_2019, , Theorem 15), where the authors investigate a primal averaging variant of SGD, which at every iteration evaluates the gradient at $\bar{x}_{t}$ instead of $x_{t}$. For $\gamma L\in(0,1)$, they establish the upper bound
$\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\frac{\gamma}{2(1-\gamma L)}\sigma_{*}^{2}.$
As our bound is strictly better, although asymptotically equivalent, we see that we do not need to rely on primal averaging to obtain such bound.
Remark 3(About the bounded variance assumptions)
Our results indicate that the commonly imposed assumptions of uniformly bounded variance or uniformly bounded gradients are not necessary to obtain improved complexity guarantees for SGD. In particular, the worst-case function arising in the proof of Theorem 4.1 has bounded gradients, implying that the bias term cannot be improved by assuming gradient boundedness. Furthermore, (taylor_stochastic_2019, , Theorem 6) shows that, under the assumption of uniformly bounded variance and for step-sizes satisfying $\gamma L\in(0,\tfrac{1+\sqrt{5}}{2}]$, one obtains the upper bound
$\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\gamma\frac{1+L}{2}\sigma^{2},$
which is looser than the bound established in our work.
4.2 Critical Step-Size $\gamma L=1$
We now turn to the case $\gamma L=1$, which we call the critical step-size, because of the sudden qualitative change in behavior. While standard for GD, this case has surprisingly not been included in any complexity rates for SGD without bounded variance assumptions. The only result we are aware of is (DOLoiLarMit23a, , Theorem 3), which provides a bound for the averaged gradient norm $\tfrac{1}{T}\sum_{t=0}^{T-1}\|\nabla f(x_{t})\|^{2}$, and which is not a complexity bound since its variance term is independent of $\gamma$ and is essentially equal to $\sigma_{*}^{2}$ (see (garrigos_handbook_2024, , Lemma 4.18.1)). As far as we know, what follows is the first complexity bound for the ergodic function values provided in this setting.
Considering the result of Theorem 4.1 for short step-sizes and letting $\gamma L\to 1$, we note that the bias term tends to $\tfrac{1}{2\gamma T}$ while the variance term diverges since its denominator is a multiple of $1-\gamma L$. This suggests there might be a complication to obtain tight bounds in the case $\gamma L=1$. As discussed in Section 7, we have empirical results showcasing that it is not possible to obtain a bound for SGD where the bias term is $\tfrac{1}{2\gamma T}$ for $\gamma L=1$ whilst preserving a finite variance term. Therefore, in this section, we will prove bounds where the bias term is arbitrarily close to $\tfrac{1}{2\gamma T}$.
Theorem 4.2(Convex case, critical step-size)
Let Assumptions 1 and 2 hold true with $\mu=0$. Let $(x_{t})$ be generated by (SGD) with $\gamma L=1$. Then, for every $\varepsilon\in(0,1)$ and $T\geq 1$, with $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[f(x_{t})-\min f\right]\leq\frac{\|x_{0}-x_{*}\|^{2}}{2(1-\varepsilon)\gamma T}+\frac{1+\varepsilon}{2\varepsilon(1-\varepsilon)}\gamma\sigma_{*}^{2}.$
Proof
It suffices to select the following parameters in Theorem 3.1:
$\rho=2(1-\varepsilon)\gamma,\quad a_{t}\equiv a=1,\quad\alpha_{t}=\gamma^{2},\quad\beta_{t}=\gamma^{2}\varepsilon,\quad\text{and}\quad e_{t}=\gamma^{2}\cdot\frac{1+\varepsilon}{\varepsilon}.$
A detailed proof of their validity, including the motivation for their choice, is presented in the proof of Theorem A.1.
We moreover show that a better bias-term, corresponding to $\varepsilon=0$, is not achievable through our proof technique. The implications of this result are further discussed in Remark 4.
Proposition 1
The sufficient conditions of Theorem 3.1 cannot be verified with $\rho=2\gamma$, $a_{0}=1$ and $\gamma L=1$.
Proof
The full details are spelled out in Appendix A.1.2. The proofs proceeds considering two quadratrics of equal curvature centered at distinct points, which lead to a contradiction with the sufficient conditions of Theorem 3.1.
Remark 4(Conjecture regarding singularity)
Although it would seem like a flaw of our analysis that $\varepsilon=0$ is not feasible in Theorem 4.2, we conjecture that obtaining a bias term of $2\gamma$ is not possible for $\gamma L=1$. Our first evidence is numerical, and will be presented in Section 7. Our second evidence is that the sufficient conditions of Theorem 3.1 cannot be satisfied when both $\gamma L=1$ and $\rho=2\gamma$, see Proposition 1. Although Theorem 3.1 provides only sufficient conditions, it is important to note that these conditions are nonetheless sharp enough to yield bias-optimal bounds for all non-critical step-sizes. The resulting singularity at $\gamma L=1$ is therefore unexpected, and it suggests that achieving a bias term of order $\tfrac{L}{2T}$ is unreachable for SGD in the absence of variance assumptions. This stands in contrast to both deterministic gradient descent and stochastic gradient descent under a uniform bounded variance assumption, for which this bias term may be reached (taylor_stochastic_2019, , Theorem 6).
Remark 5(Stochastic proximal algorithm)
While the convergence rates for the deterministic proximal algorithm are well-studied, complexity results for its stochastic counterpart are unknown in a general setting. A classical trick allows to rewrite the stochastic proximal algorithm as a particular instance of SGD, for a problem with no bounded variance, with a step-size of $\gamma L=1$. Theorem 4.2 then enables us to derive the first complexity bound for this algorithm with no other assumption on the functions than convexity, avoiding, for instance, to impose finiteness and Lipschitzness as in davis_stochastic_2019 . For more details, see Section 6.
Naturally, when interpolation holds (i.e., when $\sigma_{*}=0$) the variance term vanishes, allowing us to take the limit $\varepsilon\to 0$ and thereby obtain an optimal convergence rate. We note that an identical rate was previously established in (taylor_stochastic_2019, , Theorem 8) for a modified version of SGD incorporating a primal averaging scheme.
Corollary 1(Convex case with interpolation, critical step-size)
Let Assumptions 1 and 2 hold true, with $\mu=0$, and assume that $\sigma_{*}^{2}=0$. Let $(x_{t})$ be generated by (SGD) with $\gamma L=1$. Then, for every $T\geq 1$, with $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[f(x_{t})-\min f\right]\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}.$
Moreover, the above bound is optimal, in the sense that there exists a problem for which $\sigma^{2}_{*}=0$ and both inequalities are equalities.
Proof
Apply Theorem 4.2 with $\varepsilon\to 0$ to get the bound. The bias-optimality is again achieved through the same Huber function as for $\gamma L\in(0,1)$, see the proof of Proposition 3.
4.3 Large Step-Sizes $\gamma L>1$
The case $\gamma L\in(1,2)$, referred to as large step-sizes, is yet unexplored for SGD without bounded variance assumptions. We present below the first result of this type.
Theorem 4.3(Convex case, large step-sizes, optimal bias)
Let Assumptions 1 and 2 hold true, with $\mu=0$, and let $(x_{t})$ be generated by (SGD) with $\gamma L\in(1,2)$. Then, for every $T\geq 1$, with $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[f(x_{t})-\min f\right]\leq\frac{\delta_{T}\|x_{0}-x_{*}\|^{2}}{2\gamma(2-\gamma L)T}+\frac{\gamma\bar{e}_{T}^{\prime}\sigma_{*}^{2}}{2(2-\gamma L)^{3}},$
where $\delta_{T}=1-(\gamma L-1)^{2T}\in(0,1)$, and $\bar{e}_{T}^{\prime}$ grows exponentially like $\frac{1}{T(\gamma L-1)^{2T-2}}$. Moreover, this bound is bias-optimal, in the sense that there exists a problem for which $\sigma_{*}^{2}=0$ and the second inequality is an equality.
Proof
It suffices to select the following parameters in Theorem 3.1, with $\theta=(\gamma L-1)^{2}$:
$\rho=\frac{2\gamma(2-\gamma L)}{1-\theta^{T}},\quad a_{t}=\frac{1-\theta^{T-t}}{1-\theta^{T}},\quad\beta_{t}=\frac{\gamma(\gamma L-1)(1-\theta^{T-t-1})}{L(1-\theta^{T})},\quad\alpha_{t}=\beta_{t}+\frac{\rho}{2L},$
$e_{t}=\frac{\gamma^{2}}{2-\gamma L}\frac{1-\theta^{T-t-1}}{1-\theta^{T}}\left(\frac{\gamma L}{\theta^{T-t-1}}-2(\gamma L-1)\right).$
A detailed proof of their validity, including the motivation for their choice, is presented in Appendix A.1.3. Proving that the bias term is optimal can be done by considering $f(x)=\tfrac{L}{2}\|x\|^{2}$, see Proposition 7 in the appendix for more details.
Remark 6(About bias-optimality)
The bound above is bias-optimal for the average function value gap, but not for the ergodic function value gap $f(\bar{x}_{T})-\inf f$. This limitation arises from the fact that our Lyapunov analysis is inherently designed to produce bounds on the former quantity, rather than the latter. In the short step-size regime, we benefited from the fortunate circumstance that the worst-case objective is locally linear, which makes the two performance metrics coincide along the iterates. However, this argument does not extend to the present setting, and we are not aware of any existing proof technique capable of establishing SGD bounds directly for the ergodic function value gap.
Remark 7(About the exponential variance)
The variance term in Theorem 4.3 increases exponentially with $T$, making our bound unpractical unless interpolation holds. A couple of comments must be made here. First, we believe that such exponential growth cannot be avoided when looking for a bias-optimal bound, based on a numerical experiment described in Section 7. Second, this degenerate variance term is akin to what happened in the critical step-size regime, where the bias-optimal bound is seemingly impossible to attain due to an infinite variance term. While the price to pay for bias-optimality is too high here, one can obtain a bounded variance term provided the bias term is allowed to be suboptimal. A couple months after our paper was made available online, a concurrent work independently derived a bound for SGD in the large step-size regime (attia_fast_2025, , Theorem 8), where a suboptimal bias was traded with a bounded variance term. We illustrate below that such result can easily be derived within the Lyapunov framework of Theorem 3.1, by using exactly the same Lyapunov coefficients than in the critical step-size regime.
Theorem 4.4(Convex case, large step-sizes, suboptimal bias)
Let Assumptions 1 and 2 hold true with $\mu=0$, and let $(x_{t})$ be generated by (SGD) with $\gamma L\in(1,2)$. Then, for every $\varepsilon\in(0,1)$, every $T\geq 1$, with $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\mathopen{}\left[{f(\bar{x}_{T})-\min f}\right]\mathclose{}\leq\frac{1}{T}\sum_{t=0}^{T-1}\mathbb{E}\left[f(x_{t})-\min f\right]\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma(1-\varepsilon)(2-\gamma L)T}+\frac{\gamma L+\varepsilon(2-\gamma L)}{2\varepsilon(1-\varepsilon)(2-\gamma L)^{2}}\gamma\sigma_{*}^{2}.$
Proof
It suffices to select the following parameters in Theorem 3.1:
$\rho=2\gamma(1-\varepsilon)(2-\gamma L),\quad a_{t}=1,\quad\alpha_{t}=\frac{\gamma}{L},\quad\beta_{t}=\frac{\gamma}{L}\left[(\gamma L-1)+\varepsilon(2-\gamma L)\right],\quad e_{t}=\gamma^{2}\frac{\gamma L+\varepsilon(2-\gamma L)}{\varepsilon(2-\gamma L)}.$
A detailed proof of their validity, including the motivation for their choice, is presented in Appendix A.1.3.
Remark 8(Related works)
As far as we are aware of, Theorem 4.3 is the first bound provided for SGD with large step-sizes in the convex smooth case. We believe the closest anterior result to be (taylor_stochastic_2019, , Theorem 8) where the authors consider a primal averaging modification of SGD, further assume interpolation to hold, and obtain a convergence rate slightly larger than the bias term of Theorem 4.3. This suggests again that such primal averaging is not needed to derive satisfying bounds for SGD. The result in (attia_fast_2025, , Theorem 8), published after the first version of our manuscript, is comparable to ours and corresponds to Theorem 4.4 with $\varepsilon=3/4$ and a variance term slightly larger than ours. We also observe that their variance can be directly recovered within the framework of our Theorem 3.1, by using the exact same Lyapunov parameters, except for $\beta_{t}$ which is taken as $\gamma^{2}/2$.
5 Bias-Optimal Bounds in the Strongly Convex Smooth Setting
As in the convex case, results for strongly convex problems are obtained after carefully choosing Lyapunov parameters via Theorem 3.1, in view to derive rates with Lemma 8. As in the convex case, the results in this section are presented with slight simplifications, for the ease of presentation. The more detailed bounds may be found in the appendix. Our first Theorem 5.1 provides a new bias-optimal bound which is valid for almost all step-sizes $\gamma L\in(0,2)$.
Theorem 5.1(Strongly convex case, non-critical step-size)
Let Assumptions 1 and 2 hold, with $\mu>0$, and let $(x_{t})$ be generated by (SGD) with $\gamma L\in(0,2)$. If $\gamma\neq\gamma_{\operatorname{crit}}:=\tfrac{2}{\mu+L}$ then, for every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\phi^{2T}\cdot\|x_{0}-x_{*}\|^{2}+\frac{1-\phi^{2T}}{1-\phi^{2}}\frac{\phi\gamma_{\operatorname{crit}}}{|\gamma-\gamma_{\operatorname{crit}}|}\gamma^{2}\sigma_{*}^{2},$
where $\phi=\max\{1-\gamma\mu;\gamma L-1\}\in[0,1)$. Moreover the above bound is bias-optimal, in the sense that there exists a problem for which $\sigma_{*}^{2}=0$ and the inequality is an equality.
Proof
The proof distinguishes two cases. If $L>\mu$, we select the following parameters in Theorem 3.1:
$\rho=0,\quad a_{t}=\phi^{2(T-t)},\quad\alpha_{t}=\beta_{t}=\frac{\gamma\phi}{L-\mu}\cdot a_{t+1},\quad e_{t}=\frac{\phi\gamma_{\operatorname{crit}}}{|\gamma-\gamma_{\operatorname{crit}}|}\cdot\gamma^{2}a_{t+1}.$
If $L=\mu$, we proceed with a limiting argument. We fix $\varepsilon\in(0,1-\phi^{2})$ and select the following parameters in Theorem 3.1:
$\rho=0,\quad a_{t}=(\phi^{2}+\varepsilon)^{T-t},\quad\alpha_{t}=\beta_{t}=\alpha a_{t+1},\quad e_{t}=\frac{2\gamma^{2}\alpha}{2\alpha-\gamma^{2}}\cdot a_{t+1}$
where
$\alpha=\max\left\{\frac{\gamma^{2}}{2}\left(1+\frac{\phi^{2}}{\varepsilon}\right),\frac{1-\phi^{2}-\varepsilon}{2L^{2}}\right\}.$
We thus obtain a bound from Lemma 8, and then take $\varepsilon\to 0$. A detailed proof of the validity of those parameters, including the motivation for their choice, is presented in Appendix A.2. The bias-optimality of the bound follows from the standard fact that the functions $f(x)=\tfrac{L}{2}\|x\|^{2}$ and $\tfrac{\mu}{2}\|x\|^{2}$ provide worst-case rates (see Appendix A.2 for the details).
Remark 9(Related works)
As for the convex case, Theorem 5.1 provides a better bound than what can be found in the literature, providing a bias term of the order $\phi^{2T}$, where the geometrical rate $\phi^{2}$ corresponds exactly to the optimal rate for GD. The result also extends the range of step-sizes to $\gamma L\in(0,2)$, which was not done in previous works. The first boundsderived under the same assumptions as Theorem 5.1 were obtained in (bach_non-asymptotic_2011, , Theorem 1). For $\gamma L\in(0,\tfrac{\mu}{L})\subset(0,1)$, they obtain a bias term $\varphi^{T}$, where $\varphi=1-2\mu\gamma+2L^{2}\gamma^{2}$ is larger than our $\phi^{2}$, which is $(1-\gamma\mu)^{2}$ in this context. Those results were improved in (needell_stochastic_2016, , Theorem 2.1), where the authors extend the analysis to $\gamma L\in(0,1)$ with a bound governed by $\varphi^{T}$, where $\varphi=1-2\gamma\mu(1-\gamma L)$ is again larger than our $\phi^{2}$. A similar bound was obtained more recently in (gower_sgd_2019, , Theorem 3.1), where the authors notably showed that their result remains valid in a nonconvex setting.
Remark 10(Conjecture regarding singularity)
Our result is valid for every non-critical step-size $\gamma L\in(0,2)$. When $\gamma\to\gamma_{\operatorname{crit}}$ the bias term converges to $\left(\tfrac{L-\mu}{L+\mu}\right)^{2T}$, which is the best worst-case rate for GD. Unfortunately this is unattainable in our analysis as the variance term tends to $\infty$. It is remarkable that covering large step-sizes allows us to observe a singularity phenomenon which, to the best of our knowledge, was never observed before. We conjecture again that, in general, one cannot derive a bound for SGD for which the bias term is equal to the best worst-case rate of GD among all possible constant step-sizes. It seems reasonable to think that achieving the optimal rate for a deterministic method requires such precision that introducing a small perturbation or variance might break it. Such conjecture seems to be validated by our numerical experiments presented in Section 7. Our next Theorem 5.2 provides a bound for the critical step-size, where a sub-optimal bias term is traded with a bounded variance term.
Theorem 5.2(Strongly convex case, critical step-size, suboptimal bias)
Let Assumptions 1 and 2 hold true, with $\mu>0$, and let $\gamma_{\operatorname{crit}}\coloneq\tfrac{2}{\mu+L}$, $\phi_{\operatorname{crit}}=\tfrac{L-\mu}{L+\mu}$. Let $(x_{t})$ be generated by (SGD) with $\gamma=\gamma_{\operatorname{crit}}$. Then, for any $\phi\in(\phi_{\operatorname{crit}},1)$ and every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\phi^{2T}\cdot\|x_{0}-x_{*}\|^{2}+\frac{1-\phi^{2T}}{1-\phi^{2}}\frac{\phi^{2}}{\phi^{2}-\phi_{\operatorname{crit}}^{2}}\gamma^{2}\sigma_{*}^{2}.$
Proof
If $L>\mu$, it suffices to select the following parameters in Theorem 3.1:
$\rho=0,\quad a_{t}=\phi^{2(T-t)},\quad\alpha_{t}=\beta_{t}=\frac{2\phi^{2}}{(L-\mu)^{2}}\cdot a_{t+1},\quad e_{t}=\frac{\phi^{2}}{\phi^{2}-\phi_{\operatorname{crit}}^{2}}\cdot\gamma^{2}a_{t+1}.$
On the other hand, if $L=\mu$, it suffices to select the following parameters in Theorem 3.1:
$\rho=0,\quad a_{t}=\phi^{2(T-t)},\quad\alpha_{t}=\beta_{t}=\alpha a_{t+1},\quad e_{t}=\frac{2\gamma^{2}\alpha}{2\alpha-\gamma^{2}}a_{t+1},$
where
$\alpha=\max\left\{\frac{\gamma^{2}}{2}\left(1+\frac{\phi_{opt}^{2}}{\varepsilon}\right),\frac{1-\phi^{2}_{opt}-\varepsilon}{2L^{2}}\right\},$
and take $\varepsilon\to 0$. A detailed proof of the validity, including the motivation for their choice, is presented in Appendix A.2.
As in the convex case, when interpolation holds the variance blowing up is no longer a problem, allowing us to take an optimal bias.
Corollary 2(Strongly convex case with interpolation, critical step-size)
Let Assumptions 1 and 2 hold true, with $\mu>0$, and assume that $\sigma_{*}^{2}=0$. Let $(x_{t})$ be generated by (SGD) with $\gamma=\tfrac{2}{\mu+L}$. Then, for every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\left(\frac{L-\mu}{L+\mu}\right)^{2T}\|x_{0}-x_{*}\|^{2}.$
Moreover the above rate is optimal, in the sense that there exists a problem for which $\sigma_{*}^{2}=0$ and the inequality is an equality.
6 Stochastic Proximal Algorithm
6.1 Main Result
In the deterministic setting, one can approximate the minimizers of a nonsmooth (and possibly not everywhere finite) convex function by means of the Proximal Point Algorithm (martinet_breve_1970, ; rockafellar_monotone_1976, ). In this section, we explain how the study of SGD for the critical step-size $\gamma L=1$ allows us to study the convergence of the Stochastic Proximal Algorithm as a byproduct, and discuss how the results obtained compare to existing ones.
Problem 2
Let $\mathcal{H}$ be a real Hilbert space with inner product $\langle\cdot,\cdot\rangle$ and induced norm $\|\cdot\|$. Let $\{f_{i}\}_{i\in\mathcal{I}}$ be a family of extended real-valued functions $f_{i}\colon\mathcal{H}\to\mathbb{R}\cup\{$$\infty\}$, where $\mathcal{I}$ is a (possibly infinite) set of indices. We consider the problem of minimizing $f\coloneqq\mathbb{E}\mathopen{}\left[{f_{i}}\right]\mathclose{}$, where the expectation is taken with respect to some probability distribution $\mathcal{D}$ over $\mathcal{I}$, such that $i\mapsto f_{i}(x)$ is $\mathcal{D}$-measurable. We assume that each $f_{i}$ is proper, convex, and lower semi-continuous, and that ${\rm{argmin}}~f\neq\emptyset$.
An algorithm of choice to solve Problem 2 is the Stochastic Proximal Algorithm, defined as
$x_{t+1}=\operatorname{prox}_{\gamma f_{i_{t}}}(x_{t}),$ (SProx)
where $i_{t}\in\mathcal{I}$ is sampled i.i.d. from the distribution $\mathcal{D}$, and where the proximal operator $\operatorname{prox}_{\gamma f_{i}}$ is defined by
$\operatorname{prox}_{\gamma f_{i}}(x)=\underset{y\in\mathcal{H}}{\rm{argmin}}~\left\{f_{i}(y)+\frac{1}{2\gamma}\|y-x\|^{2}\right\},$
with a step-size $\gamma>0$. Since $f_{i}$ is proper, convex and lower semi-continuous, $\operatorname{prox}_{\gamma f_{i}}$ is well-defined and single-valued.
Barriers to a Unified Analysis. The complexity results available for (SProx) can be divided in four categories: convex feasibility problems; problems where the functions are everywhere finite and uniformly Lipschitz-continuous; smooth problems, where the gradients are uniformly Lipschitz-continuous; and finally strongly convex problems. These range from functions intrinsically taking the value $\infty$, to finite functions with various levels of regularity. As opposed to the deterministic case, there is no obvious performance metric that makes sense in full generality. The first natural candidate is the expected function value gap, with or without averaging, namely $\mathbb{E}\left[f(\bar{x}_{T})-\min f\right]$ or $\mathbb{E}\left[f(x_{T})-\min f\right]$. However, we cannot hope to bound such gap for the Stochastic Proximal Algorithm, since nothing guarantees that the iterates $x_{T}$ (or their averages $\bar{x}_{T}$) remain in the domain of $f$. An extreme case is the convex feasibility problem (see Example 1), where $\bar{x}_{T}$ must already be a solution in order for $f(\bar{x}_{T})$ to be finite. Now, one could search for a bound on the distance between the current iterate and a solution of the problem (or the distance to the set of solutions), but such distance can decrease arbitrarily slowly even in the deterministic case (garrigos_convergence_2023, ), which is why we usually do not have bounds on $\|x_{t}-x_{*}\|$ for (non strongly) convex problems. Our analysis will provide bounds for a different metric, which is always well-defined, and can be connected to standard performance metrics.
Stochastic Proximal Algorithm as SGD. It is common knowledge that the proximal algorithm is a particular case of the gradient descent algorithm. This can be stated formally, by introducing the notion of the Moreau envelope of a convex function. Given a proper, convex, and lower semi-continuous function $f\colon\mathcal{H}\to\mathbb{R}\cup\{$$\infty\}$, and a parameter $\gamma>0$, we define $f^{\gamma}\colon\mathcal{H}\to\mathbb{R}$ by
$f^{\gamma}(x)=\inf\limits_{y\in\mathcal{H}}\left\{f(y)+\frac{1}{2\gamma}\|y-x\|^{2}\right\}.$
The Moreau envelope $f^{\gamma}$ is finite and differentiable. Its gradient, given by $\nabla f^{\gamma}(x)=\frac{x-\operatorname{prox}_{\gamma f}(x)}{\gamma}$, is Lipschitz continuous with constant $\tfrac{1}{\gamma}$ (bauschke_convex_2017, , Proposition 12.30). In particular, computing $\operatorname{prox}_{\gamma f}(x)=x-\gamma\nabla f^{\gamma}(x)$ is the same as computing one step of gradient descent for the function $f^{\gamma}$, with step-size $\gamma$. Since $\gamma\operatorname{Lip}(\nabla f^{\gamma})=1$, (SProx) can be seen as an instance of (SGD) applied to the function
$F^{\gamma}(x)\coloneqq\mathbb{E}\mathopen{}\left[{f_{i}^{\gamma}(x)}\right]\mathclose{},$
with the step-size being critical. This connection was already discussed and exploited in patrascu_nonasymptotic_2018 ; necoara_randomized_2019 . The novelty of our approach is that we are able to derive complexity rates for (SProx) from the ones we obtain from (SGD) when $\gamma L=1$.
Theorem 6.1(Bound for SProx. General case)
Considering Problem 2, let $\gamma>0$ and assume that ${\rm{argmin}}~F^{\gamma}\neq\emptyset$. Let $(x_{t})$ be generated by (SProx) with step-size $\gamma$. Then, for every $T\geq 1$, with $\bar{x}_{T}=\tfrac{1}{T}\sum_{t=0}^{T-1}x_{t}$,
$\mathbb{E}\left[F^{\gamma}(\bar{x}_{T})-\min F^{\gamma}\right]\leq\frac{\|x_{0}-x_{*}^{\gamma}\|^{2}}{\gamma T}+3\gamma\sigma^{2}_{*}(\gamma),$
where $\sigma^{2}_{*}(\gamma)=\mathbb{E}\left[\|\nabla f_{i}^{\gamma}(x_{*}^{\gamma})\|^{2}\right]$, and $x_{*}^{\gamma}\in{\rm{argmin}}~F^{\gamma}$. If we further assume that each $f_{i}$ is $\mu$-strongly convex ($\mu>0$), then
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}^{\gamma}\|^{2}}\right]\mathclose{}\leq\left(\frac{1}{1+\gamma\mu}\right)^{2T}\|x_{0}-x_{*}^{\gamma}\|^{2}+\frac{2(1+\gamma\mu)}{\mu}\gamma^{2}\sigma_{*}^{2}(\gamma).$
Proof
For the convex case, apply Theorem 4.2 with $\varepsilon=1/2$. For the strongly convex case, apply Theorem 5.1. Remember that each $f_{i}^{\gamma}$ is $L_{\gamma}$-smooth with $\gamma L_{\gamma}=1$, and it a standard duality exercise to show that each $f_{i}^{\gamma}$ is $\mu_{\gamma}$-strongly convex, with $\mu_{\gamma}=\tfrac{\mu}{1+\gamma\mu}$. It is then a matter of simplifying the obtained bound
$\phi_{\gamma}=1-\frac{\mu_{\gamma}}{L_{\gamma}}\quad\implies\quad\frac{1}{1-\phi_{\gamma}^{2}}\frac{\tfrac{2}{\mu_{\gamma}+L_{\gamma}}\phi_{\gamma}}{|\tfrac{2}{\mu_{\gamma}+L_{\gamma}}-\tfrac{1}{L_{\gamma}}|}=\frac{2(1+\gamma\mu)^{2}}{\mu(2+\gamma\mu)}\leq\frac{2(1+\gamma\mu)}{\mu},$
yielding the wanted bound.
This is the first time that a complexity bound on the Stochastic Proximal Algorithm is derived with no other assumption than convexity. We note some unusual terms in those bounds, such as the use of the regularized function $F^{\gamma}=\mathbb{E}\mathopen{}\left[{f_{i}^{\gamma}}\right]\mathclose{}$ for the metric instead of $f=\mathbb{E}\mathopen{}\left[{f_{i}}\right]\mathclose{}$, or the use of the regularized minimizer $x_{*}^{\gamma}$ instead of $x_{*}$. The presence of $x_{*}^{\gamma}$ is an artifact of our approach, that we were unable to avoid, except for interpolated problems, and leave this technical issue for a future work. In the following subsections we illustrate that the optimality metric $F^{\gamma}$ can be related to metrics used in the literature under standard additional assumptions, allowing us to recover several classical results, together with a novel result in the interpolation setting. One class of results remain out of our reach at the moment, which we discuss next.
Remark 11(Strongly Convex Problems)
As far as we know, Theorem 6.1 is the first bound available for (SProx) applied to strongly convex problems, without any further assumptions. Nevertheless, our bound controls the distance of the iterates to the minimizer $x_{*}^{\gamma}$ of the regularized objective, instead of the original objective $f$. Some works were able to provide similar bounds controlling $\mathbb{E}[\|x_{T}-x_{*}\|^{2}]$ directly, under additional regularity assumptions. For instance (asi_stochastic_2019, , Proposition 5) assumes that all the functions $f_{i}$ have a common and closed domain. Similarly, (patrascu_nonasymptotic_2018, , Theorem 10), which analyzes a slightly different variant of the stochastic proximal algorithm, relies on the same assumption. More recently (richtarik_unified_2024, , Theorem 2) considered functions $f_{i}$ having a full domain. We leave for a future work the question of connecting Theorem 6.1 to the above mentioned results.
6.2 Interpolated Problems
Let us remind, as stated in Definition 1, that interpolation holds when the functions $f_{i}$ share a common minimizer. Using standard convex analysis tools, one can show that if the family $(f_{i})_{i\in\mathcal{I}}$ interpolates, then so does $(f_{i}^{\gamma})_{i\in\mathcal{I}}$, which means that we can use Corollary 1. We further have ${\rm{argmin}}~F^{\gamma}={\rm{argmin}}~f$ and $\inf F^{\gamma}=\inf f$, which means that the function value gap $F^{\gamma}(x)-\inf F^{\gamma}$ is a meaningful metric for our original problem, in the sense that it provides a faithful measure of how far we are from optimality.
Theorem 6.2(Bounds for SProx. Interpolation case)
Consider Problem 2 and assume that interpolation holds. Let $(x_{t})$ be generated by (SProx) with step-size $\gamma>0$. Then for every $x_{*}\in{\rm{argmin}}~f$,
$\mathbb{E}\left[F^{\gamma}(\bar{x}_{T})-\inf f\right]\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}.$
If we further assume that each $f_{i}$ is $\mu$-strongly convex ($\mu>0$) then
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\left(\frac{1}{1+\gamma\mu}\right)^{2T}\|x_{0}-x_{*}\|^{2}.$
Proof
Let us first verify that interpolation for the original problem also applies to the regularized one. If interpolation holds for the family $(f_{i})_{i\in\mathcal{I}}$, then $\cap~{\rm{argmin}}~f_{i}\neq\emptyset$. As shown, for instance, in (bauschke_convex_2017, , Proposition 12.9), $\inf f_{i}^{\gamma}=\inf f_{i}$ and ${\rm{argmin}}~f_{i}^{\gamma}={\rm{argmin}}~f_{i}$. As such $\cap~{\rm{argmin}}~f_{i}^{\gamma}\neq\emptyset$ and interpolation holds for the regularized problem. We can then use Lemma 5 to see that
${\rm{argmin}}~F^{\gamma}=\bigcap{\rm{argmin}}~f_{i}^{\gamma}=\bigcap{\rm{argmin}}~f_{i}={\rm{argmin}}~f\quad\hbox{and}\quad\inf F^{\gamma}=\mathbb{E}\mathopen{}\left[{\inf f_{i}^{\gamma}}\right]\mathclose{}=\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}=\inf f.$
We can now apply Corollary 1 in the convex case, or Theorem 6.1 in the strongly convex case.
The above result is new in the convex case. We are not aware of any result providing bounds for (SProx) under the sole assumption of interpolation. The result in the strongly convex case can be compared to (tovmasyan_revisiting_2025, , Theorem 6.1), where the authors obtain a worse rate when the function $f$ is strongly convex, under the additional restrictions that $\gamma\leq\tfrac{\mu}{2}$, that the functions are differentiable, and that a bound on the gradient variance, called star-similarity, is verified.
Example 1(Feasibility problems)
Feasibility problems consists in finding $x_{*}\in C\coloneqq\cap_{i}C_{i}$ where $C_{i}$ are nonempty closed convex sets. A standard algorithm for solving this problem is the Stochastic Projection Algorithm, where the current iterate is projected onto a set $C_{i}$ is chosen at random: $x_{t+1}=\operatorname{proj}_{C_{i_{t}}}(x_{t})$. This is a particular case of the (SProx) algorithm applied to Problem 2 where $f_{i}$ is taken as the indicator function of $C_{i}$, which is 0 if $x\in C_{i}$ and is $\infty$ if $x\notin C_{i}$. A first important observation is that this problem is well-posed ($f$ has minimizers) if and only if interpolation holds. This is due to the fact that ${\rm{argmin}}~f=C\coloneqq\cap_{i}C_{i}=\cap_{i}{\rm{argmin}}~f_{i}$. A second point is that the regularized function value gap is simply $F^{\gamma}(x)-\inf F^{\gamma}=\tfrac{1}{2\gamma}\mathbb{E}\mathopen{}\left[{\operatorname{dist}(x;C_{i})^{2}}\right]\mathclose{}$ which is a standard measure of how far $x$ is from the set of solutions $C=\cap_{i}C_{i}$. We can therefore apply Theorem 6.2 and obtain that the Stochastic Projection Algorithm enjoys the rate (take for instance $\gamma=\tfrac{1}{2}$ and $x_{*}=\operatorname{proj}_{C}(x_{0})$)
$\mathbb{E}\mathopen{}\left[{\operatorname{dist}(\bar{x}_{T};C_{i})^{2}}\right]\mathclose{}\leq\frac{\operatorname{dist}(x_{0};C)^{2}}{T}.$
The above bound is well-known and corresponds exactly to what was obtained in (nedic_random_2010, , Proposition 6). It is remarkable that our analysis on SGD converted into (SProx) is able to provide exactly the best known rates for the Stochastic Projection Algorithm. Of course the story does not stop here, as it is possible to derive better bounds provided that the intersection $\cap_{i}C_{i}$ is regular Kru06 ; LewLukMal09 . Such regularity ensures that the problem enjoys a Łojasiewicz inequality AttBolSva13 ; bolte_error_2017 which can be exploited to derive linear rates of convergence (necoara_randomized_2019, , Theorem 3). We leave for future work the question whether our analysis can be adapted to exploit such regularity features.
6.3 Regular Problems: Lipschitz, Smooth
If the functions $f_{i}$ are everywhere finite, the function value gap $f(x)-\inf f$ becomes an appropriate performance metric. So it becomes reasonable to expect the results of Theorem 6.1 to hold for $f(x)-\inf f$ instead of $F^{\gamma}(x)-\inf F^{\gamma}$. We are going to show that both quantities are related when the problem is regular, which we decline in two flavors: Lipschitz and smooth problems.
Corollary 3(Bound for SProx. Regular cases)
Consider Problem 2, let $\gamma>0$ and assume that ${\rm{argmin}}~F^{\gamma}\neq\emptyset$. Let $(x_{t})$ be generated by (SProx), with step-size $\gamma>0$.
  1. 1.
    If each function $f_{i}$ is $G$-Lipschitz continuous, then
    $\mathbb{E}\left[f(\bar{x}_{T})-\inf f\right]\leq\frac{\|x_{0}-x_{*}^{\gamma}\|^{2}}{\gamma T}+4\gamma G^{2},\qquad\hbox{where}\quad x_{*}^{\gamma}\in{\rm{argmin}}~F^{\gamma}.$
  2. 2.
    If each function $f_{i}$ is $L$-smooth, if $\mathbb{E}[\inf f_{i}]>$-$\infty$, and if $\gamma L\in(0,1)$, then
    $\mathbb{E}\left[f(\bar{x}_{T})-\inf f\right]\leq\frac{\|x_{0}-x_{*}^{\gamma}\|^{2}}{(1-\gamma L)\gamma T}+\frac{7\gamma L}{1-\gamma L}\Delta_{*},\qquad\hbox{where}\quad x_{*}^{\gamma}\in{\rm{argmin}}~F^{\gamma},\quad\Delta_{*}=\inf f-\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}.$
Proof
In this proof we will use the subdifferential of $f_{i}$ at a point $x$, which is the set
$\partial f_{i}(x)\coloneqq\{g\in\mathcal{H}\ |\ (\forall y\in\mathcal{H})\ f_{i}(y)-f_{i}(x)-\langle g,y-x\rangle\geq 0\}.$
This set is nonempty, convex and closed (bauschke_convex_2017, , Propositions 16.4 and 16.27), in particular due to the fact that we assume $f_{i}$ to take finite values. We denote its least-norm element by $\partial^{0}f_{i}(x)$. Using the definition of subdifferential, we can write for $x\in\mathcal{H}$
$f_{i}(x)-f_{i}^{\gamma}(x)=\sup_{y\in\mathcal{H}}f_{i}(x)-f_{i}(y)-\frac{1}{2\gamma}\|y-x\|^{2}\leq\sup_{y\in\mathcal{H}}-\langle\partial^{0}f_{i}(x),y-x\rangle-\frac{1}{2\gamma}\|y-x\|^{2}$
The supremum on the right-hand side is attained at $y=x-\gamma\partial^{0}f_{i}(x)$, which yields $f_{i}(x)-f_{i}^{\gamma}(x)\leq\frac{\gamma}{2}{\|\partial^{0}f_{i}(x)\|^{2}}$. We further know that $f_{i}^{\gamma}\leq f_{i}$, which means that $F^{\gamma}\leq f$ and so $\inf F^{\gamma}\leq\inf f$. Combining all the above inequalities, we deduce that
$(\forall x\in\mathcal{H})\quad f(x)-\inf f\leq F^{\gamma}(x)-\inf F^{\gamma}+\frac{\gamma}{2}\mathbb{E}\mathopen{}\left[{\|\partial^{0}f_{i}(x)\|^{2}}\right]\mathclose{}.$ (5)
Combining this inequality with Theorem 6.1, we obtain
$\mathbb{E}\left[f(\bar{x}_{T})-\inf f\right]\leq\frac{\|x_{0}-x_{*}^{\gamma}\|^{2}}{\gamma T}+3\gamma\sigma_{*}^{2}(\gamma)+\frac{\gamma}{2}\mathbb{E}\mathopen{}\left[{\|\partial^{0}f_{i}(\bar{x}_{T})\|^{2}}\right]\mathclose{},$
where $x_{*}^{\gamma}\in{\rm{argmin}}~F^{\gamma}$ and $\sigma^{2}_{*}(\gamma)=\mathbb{E}\left[\|\nabla f_{i}^{\gamma}(x_{*}^{\gamma})\|^{2}\right]$. We are now ready to derive our bounds in the Lipschitz and smooth cases.
  1. 1.
    The $G$-Lipschitzness of $f_{i}$ guarantees that $\|\partial^{0}f_{i}(\bar{x}_{T})\|^{2}\leq G^{2}$. Moreover, $\|\nabla f_{i}^{\gamma}(x_{*}^{\gamma})\|\leq\|\partial^{0}f_{i}(x_{*}^{\gamma})\|$ (see (patrascu_nonasymptotic_2018, , Lemma 3)) so we also have $\|\nabla f_{i}^{\gamma}(x_{*}^{\gamma})\|^{2}\leq G^{2}$ and in turn $\sigma_{*}^{2}(\gamma)\leq G^{2}$. This leads to the desired variance term $3\gamma G^{2}+\tfrac{\gamma}{2}G^{2}\leq 4\gamma G^{2}$.
  2. 2.
    If the functions are $L$-smooth, it is easy to bound $\sigma_{*}^{2}(\gamma)$ with $\Delta_{*}$. To do so, take $x_{i}\in{\rm{argmin}}~f_{i}^{\gamma}={\rm{argmin}}~f_{i}$ and use the convexity and smoothness of $f_{i}$ through Lemma 1 to get
    $\|\nabla f_{i}^{\gamma}(x_{*}^{\gamma})\|^{2}=\|\nabla f_{i}^{\gamma}(x_{*}^{\gamma})-\nabla f_{i}^{\gamma}(x_{i})\|^{2}\leq 2L\left(f_{i}^{\gamma}(x_{*}^{\gamma})-f_{i}^{\gamma}(x_{i})\right)=2L\left(f_{i}^{\gamma}(x_{*}^{\gamma})-\inf f_{i}\right).$
    After taking expectation, we conclude that $\sigma_{*}^{2}(\gamma)\leq 2L\left(\inf F^{\gamma}-\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}\right)\leq 2L\left(\inf f-\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}\right)=2L\Delta_{*}$. Then we need to tackle the term $\mathbb{E}\mathopen{}\left[{\|\partial^{0}f_{i}(\bar{x}_{T})\|^{2}}\right]\mathclose{}=\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(\bar{x}_{T})\|^{2}}\right]\mathclose{}$. A now standard variance transfer argument (see (garrigos_handbook_2024, , Lemma 4.19)) shows that $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x)\|^{2}}\right]\mathclose{}\leq 2L(f(x)-\inf f)+2L\Delta_{*}$, where $\Delta_{*}\coloneqq\inf f-\mathbb{E}\mathopen{}\left[{\inf f_{i}}\right]\mathclose{}$. Reorganizing the term $f(\bar{x}_{T})-\inf f$ and dividing by $(1-\gamma L)$ leads to the desired complexity bound.
Remark 12(Convex Regular Problems)
The unifying approach that leads to Theorem 6.1 comes at a price. Directly exploiting the regularity of the functions allows to derive sharper bounds. For Lipschitz functions, both (asi_stochastic_2019, , Proposition 4) and (patrascu_nonasymptotic_2018, , Theorem 5) obtain bounds in which the bias and variance terms are both improved by constant multiplicative factors, namely
$\mathbb{E}\left[f(\overline{x}_{T})-f_{*}\right]\leq\frac{\|x_{0}-x_{*}\|^{2}}{2\gamma T}+\frac{\gamma G^{2}}{2}.$
Likewise, for smooth functions (traore_variance_2024, , Theorem 4.3) obtains the better bound
$\mathbb{E}\left[f(\overline{x}_{T})-f_{*}\right]\leq\frac{\|x_{0}-x_{*}\|^{2}}{\gamma T}+2\gamma\sigma^{2}_{*},$
for a step-size $\gamma<1/2L$. Additionally, note that the right hand side of the aforementioned bounds involve a true problem solution $x_{*}$, whilst ours unconventionally relies on the distance of the initialization $x_{0}$ to the regularized solution $x_{*}^{\gamma}$.
7 Numerical Results with PEP
In this section, we discuss both the proof strategy underlying our results and the tightness of the derived bounds. Our analysis relies on the Performance Estimation Problem (PEP) methodology, which played a central role throughout this work. The PEP framework allows one to reformulate the search for admissible Lyapunov parameters ensuring energy decrease as a semidefinite program that can be solved numerically. This approach not only yields numerical values for the optimal bias and variance terms, but also enables us to infer proofs of their optimality through the dual variables assocaited to the program. The insights provided by this framework directly inspired and guided our analytical proofs, by helping us conjecture the appropriate Lyapunov parameters, as made explicit in our proofs in Appendix A.
Section 7.1 provides a concise presentation of the PEP framework, while a more detailed discussion is deferred to Appendix B. We have provided lower bounds matching the bias term of our bounds, thus analytically proving that the bias term cannot be improved. Section 7.2 demonstrates that the variance terms obtained in our theoretical bounds match the numerically optimal ones identified through the PEP analysis.
7.1 Performance Estimation Problem Methodology
The PEP framework was initially introduced by drori_performance_2014 , further improved by taylor_smooth_2017 , and adapted by taylor_stochastic_2019 in the context of stochastic algorithms, that we will follow closely. We briefly introduce the PEP framework, specifically adapted to carry out our Lyapunov analysis, and refer the reader to Appendix B for a detailed explanation.
As demonstrated by Lemmas 7 and 8, a natural way to establish guarantees for SGD is to derive nonnegative parameters $\rho,(a_{t},e_{t})$ such that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for all $t=0,\ldots,T-1$ and for all functions $f_{1},\dots,f_{m}\in\mathcal{F}_{\mu,L}(\mathbb{R}^{d})$, where $E_{t}$ is given by (Lyapunov). We will refer to such parameters as admissible Lyapunov parameters. As our focus is on bias-optimal results, we aim to identify admissible Lyapunov parameters that minimize the bias. In particular Lemmas 7 and 8 show that the bias and variance are given by terms of the form
[…]
Thus, minimizing the bias within this framework amounts to solving
$\text{Bias}_{\text{opt}}\coloneq\inf\left\{\text{Bias}\colon\rho,\left(a_{t},e_{t}\right)\text{ are admissible Lyapunov parameters}\right\}.$
We note that this is an infinite-dimensional optimization problem, due to the functional constraints $f_{i}\in\mathcal{F}_{\mu,L}(\mathbb{R}^{d})$. Moreover, as both $a_{0}$ and $\rho$ ($a_{0}$ and $a_{T}$ in the strongly convex case) are variables of the problem, the optimization problem in inherently nonconvex. However, using standard tools from the PEP methodology, this problem can be re-cast into an equivalent finite-dimensional semidefinite program that can be solved numerically. In fact, we observe that
$\rho,\left(a_{t},e_{t}\right)\text{ are Lyapunov parameters}\iff B_{0}\leq 0,\ldots,B_{T}\leq 0,$
where
$B_{t}\coloneq\sup\left\{\mathbb{E}[E_{t+1}-E_{t}]\colon f_{1},\dots,f_{m}\in\mathcal{F}_{\mu,L}(\mathbb{R}^{d}),~~x_{0}\in\mathbb{R}^{d}\right\}.$
The essence of the PEP methodology is that the infinite-dimensional constraint $f_{i}\in\mathcal{F}_{\mu,L}(\mathbb{R}^{d})$, can be reduced to a finite collection of linear matrix inequalities. This reformulation allows the problems $(B_{t})$ to be re-cast as a conic convex optimization problems. Moreover, by considering its dual formulation, verifying whether $B_{t}\leq 0$ holds true amounts to solving a feasibility problem, denoted by $\tilde{B}_{t}$. These ideas may be summarized by the following high-level theorem.
Theorem 7.1
For all $t=0,\ldots,T$, there exists a convex conic problem $\tilde{B}_{t}$ such that $\tilde{B}_{t}$ is feasible if, and only if, $B_{t}\leq 0$. Consequently,
$\rho,\left(a_{t},e_{t}\right)\text{ are admissible Lyapunov parameters}\iff\tilde{B}_{0},\ldots,\tilde{B}_{T}\text{ are feasible.}$
Therefore, $\operatorname{Bias}_{{\operatorname{opt}}}$ admits a tight semidefinite reformulation.
It is known that the dual variables extracted when solving such semidefinite program help to derive the analytical proofs of the results. This observation motivated our proofs and allowed us to obtain analytical proofs showing bias-optimality. This process is further described in Appendix B.2.
Once $\text{Bias}_{\text{opt}}$ has been established, one can, as a by-product derive the variance parameters $(e_{t})$ that will yield the minimal variance by solving
$\text{Var}_{\text{opt}}\coloneq\inf\{\text{Var}\colon\rho,\left(a_{t},e_{t}\right)\text{ are Lyapunov parameters},\ \text{Bias}=\text{Bias}_{\text{opt}}\},$
which can, in a similar fashion, be equivalently reformulated as a semidefinite program.
7.2 Numerical Experiments
We have established the optimal bias parameters for all non-critical step-sizes in both the convex and strongly convex settings (Theorems 4.1, 4.3, and 5.1). However, our results suggest that these optimal biases are not attainable at the corresponding critical step-sizes (Theorems 4.2 and 5.2), as achieving it seem to result in an infinite variance term. In the following, we perform numerical experiments to examine the associated variance behavior when the bias is fixed to its theoretically optimal value. This allows us to verify numerically that our variance bounds are tight given the optimal bias, and, moreover, that the variance diverges as the step-size approaches the critical value. This suggests that the optimal bias cannot be achieved in these degenerate regimes. Specifically, we fix the bias term to be the optimal bias term obtained in Theorems 4.1 and 4.3 in the convex setting and in Theorem 5.1 in the strongly convex setting, and aim to minimize the associated variance term.
All of our numerical experiments [^1] Code available on https://github.com/DanielCortild/Bias-Optimal-SGD. were developed in Python 3.13 using the solvers MOSEK version 11.0.19 (mosek_mosek_2025, ) and Clarabel version 0.9.0 (goulart_clarabel_2024, ) to solve the resulting semidefinite programs, and run on Intel Xeon Platinum 8380 CPUs.
Convex Setting. The obtained minimal variance is plotted in Figure 1. For all $\gamma L\in(0,2)\backslash\{1\}$, the numerically optimal variance precisely matches the theoretically predicted one. This suggests, at least numerically, that the variance term in our bounds cannot be improved. It also demonstrates numerically that the variance diverges at the critical step-size $\gamma L=1$ hence leaving the bias to be at most $(2-\varepsilon)\gamma$ for an arbitrary $\varepsilon>0$, as presented in Theorem 4.2.
Figure 1: Theoretical and numerical average variance term, subject to optimal bias ($L=1$).
Strongly Convex Setting. As illustrated in Figure 2, the empirical and theoretical variances coincide perfectly for all $\gamma\neq\tfrac{2}{L+\mu}$, confirming the sharpness of our theoretical results. In contrast, the variance exhibits a clear divergence in the vicinity of the critical step-size, suggesting that, much like in the convex case, the theoretically optimal bias is not achievable at this critical step-size, thus justifying the need for a result with sub-optimal bias, obtained in Theorem 5.2.
Figure 2: Theoretical and numerical total variance term, subject to optimal bias ($L=1$, $\mu=0.25$ and $\tfrac{2}{\mu+L}=1.6$).
In both the convex and strongly convex settings, our analytical expressions for the variance align with the variance estimated through the PEP methodology. Although this was investigated, we were unable to identify an instance of SGD for which the variance accumulated throughout the execution of the algorithm attains the theoretical variance described in Theorems 4.1 and 4.3 for the convex case, and Theorem 5.1 for the strongly convex case. An alternative explanation could be that our Lyapunov framework, while being well suited to derive an optimal bias term in our bounds, struggles to capture the behavior of the optimal variance. Maybe a larger class of Lyapunov energies is required to obtain a lower variance term while keeping bias optimality. We leave that question open for a future work.
8 Conclusion and Perspectives
In this paper, we established new and improved upper bounds for (SGD) without imposing any variance assumptions beyond smoothness, both in the convex and strongly convex settings, and for a full range of step-sizes. Our analysis relies on the introduction of a novel Lyapunov energy (Lyapunov). By providing lower bounds that match our results in some sense, we demonstrated that the obtained bounds are bias-optimal. By doing so, our analysis raised intriguing questions regarding the presence of unexpected singularities at critical step-sizes, observed in both settings. The results, as well as their proofs, are inspired by the Performance Estimation Problem methodology.
Future Work. We leave the following questions and research directions open:
  • $\bullet$
    While this paper focuses on achieving optimal bias terms, this is not the only meaningful objective. In future work, we plan to investigate bounds that provide sharp complexity rates. We anticipate this to be more challenging, as quantities such as $\|x_{0}-x_{*}\|^{2}$ and $\sigma_{*}^{2}$ then become intrinsic problem parameters rather than fixed constants.
  • $\bullet$
    In this work, we focused on Lyapunov energies of the form (Lyapunov). It remains unclear whether sharper results could be obtained by considering more general classes of Lyapunov functions. While we showed that the bias term cannot be improved, no such guarantees are currently available for the variance term. In particular, it is conceivable that the singularities we observed are artifacts of the chosen Lyapunov structure. Investigating more complex energies therefore appears as a natural direction for future research.
  • $\bullet$
    In the convex setting, our results establish bias-optimality for the average function value gap in the large step-size regime. A natural next step would be to strengthen Theorem 4.3 in order to obtain bias-optimal bounds for the function value gap evaluated at the Cesàro average of the iterates.
  • $\bullet$
    Our analysis is restricted to constant step-sizes. Extending these results to varying or adaptive step-size schemes constitutes an interesting and potentially impactful line of work.
Acknowledgments. This work benefited from the support of the FMJH Program Gaspard Monge for optimization and operations research and their interactions with data science. Daniel Cortild acknowledges the support of the Clarendon Funds Scholarships. The authors also thank the Center for Information Technology of the University of Groningen for their support and for providing access to the Hábrók high performance computing cluster.
Appendix A Proofs of Main Results
A.1 Convex Setting
In this section, we derive parameters that satisfy the sufficient conditions of Theorem 3.1 in order to derive upper bounds for SGD in the convex setting. The propositions in this section can be verified symbolically, yet we include the proofs to provide intuition on how these parameters were derived.
A.1.1 Short Step-Sizes: $\gamma L<1$ (Proof of Theorem 4.1)
Proposition 2(Lyapunov parameters. Convex case, short step-sizes)
Let $\gamma L\in(0,1)$, and consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\bullet$
    $\rho=2\gamma+\tfrac{2(1-\gamma L)}{LT}$,
  • $\bullet$
    $a_{t}=\tfrac{T-t}{T}\tfrac{1+\gamma L(T-1)}{1+\gamma L(T-t-1)}$,
  • $\bullet$
    $\alpha_{t}\equiv\alpha=\tfrac{\rho}{2L}$,
  • $\bullet$
    $\beta_{t}\equiv\beta=0$,
  • $\bullet$
    $e_{t}=\frac{a_{t+1}\gamma^{2}\alpha_{t}}{\alpha_{t}-a_{t+1}\gamma^{2}}$.
These parameters satisfy the sufficient Lyapunov conditions of Theorem 3.1. Moreover, it holds that
$e_{t}=\frac{\gamma^{2}}{1-\gamma L}\frac{T-t-1}{T}\frac{(1-\gamma L)+\gamma LT}{(1-\gamma L)+\gamma L(T-t)}\leq\frac{\gamma^{2}}{1-\gamma L}.$
In particular, the bound in Theorem 4.1 holds true.
Proof
Notice that the sufficient conditions from Theorem 3.1 are homogeneous. Because our goal is to obtain bounds whose bias term is the smallest possible, we can without loss of generality impose that $a_{0}=1$, and try to maximize $\rho$. Through our numerical analysis, we empirically observe that we can take $\alpha_{t}\equiv\alpha>0$ and $\beta_{t}\equiv\beta=0$, thus justifying these choices. We dedicate the remainder of the proof to justify our remaining choices.
We start by focusing on Condition 5 from Theorem 3.1, which reads
$\left(a_{t+1}\gamma-L{\alpha}\right)^{2}\leq({\alpha}-a_{t+1}\gamma^{2})(a_{t}-a_{t+1})$
$\iff$ $\left(a_{t+1}\gamma-L{\alpha}\right)^{2}\leq(1-\gamma L)\alpha(a_{t}-a_{t+1})+(\gamma L\alpha-a_{t+1}\gamma^{2})(a_{t}-a_{t+1})$
$\iff$ $\left(a_{t+1}\gamma-L{\alpha}\right)^{2}-\gamma(L\alpha-a_{t+1}\gamma)(a_{t}-a_{t+1})\leq(1-\gamma L)\alpha(a_{t}-a_{t+1})$
$\iff$ $\left(L\alpha-a_{t+1}\gamma\right)(L\alpha-a_{t+1}\gamma-\gamma(a_{t}-a_{t+1}))\leq(1-\gamma L)\alpha(a_{t}-a_{t+1})$
$\iff$ $\left(L\alpha-a_{t+1}\gamma\right)(L\alpha-a_{t}\gamma)\leq(1-\gamma L)\alpha(a_{t}-a_{t+1}).$
Condition 3 boils down to $a_{t}\geq a_{t+1}$. By additionally imposing $a_{t}>a_{t+1}$, we may divide by $a_{t}-a_{t+1}$ to obtain
$\iff$ $\frac{\left(L\alpha-a_{t+1}\gamma\right)(L\alpha-a_{t}\gamma)}{a_{t}-a_{t+1}}\leq(1-\gamma L)\alpha$
$\iff$ $\frac{\left(L\alpha-a_{t+1}\gamma\right)(L\alpha-a_{t}\gamma)}{(L\alpha-a_{t+1}\gamma)-(L\alpha-a_{t}\gamma)}\leq\frac{(1-\gamma L)\alpha}{\gamma}.$
As $(a_{t})$ is decreasing, with $a_{0}=1$, it holds true that $a_{t}\leq 1$ for all $t=0,\ldots,T-1$. Specifically, if $L\alpha>\gamma$ (which we shall assume), then both $L\alpha-a_{t+1}\gamma>0$ and $L\alpha-a_{t}\gamma>0$, thus allowing us to write
$\iff$ $\frac{(L\alpha-a_{t+1}\gamma)-(L\alpha-a_{t}\gamma)}{\left(L\alpha-a_{t+1}\gamma\right)(L\alpha-a_{t}\gamma)}\geq\frac{\gamma}{(1-\gamma L)\alpha}$
$\iff$ $\frac{1}{L\alpha-a_{t}\gamma}-\frac{1}{L\alpha-a_{t+1}\gamma}\geq\frac{\gamma}{(1-\gamma L)\alpha}$
$\iff$ $\frac{1}{L\alpha-a_{t+1}\gamma}\leq\frac{1}{L\alpha-a_{t}\gamma}-\frac{\gamma}{(1-\gamma L)\alpha}.$
Introducing the temporary variables $u_{t}\coloneqq\tfrac{1}{L\alpha-\gamma a_{t}}$ and $c=\tfrac{\gamma}{(1-\gamma L)\alpha}$, we see that we obtained an arithmetic inequality $u_{t+1}\leq u_{t}-c$, leading us to $u_{t}\leq u_{0}-ct$. We know that $u_{0}=\tfrac{1}{L\alpha-\gamma}$, so we can write
$u_{t}\leq\frac{1}{L\alpha-\gamma}-\frac{t\gamma}{(1-\gamma L)\alpha}=\frac{(1-\gamma L)\alpha-t\gamma(L\alpha-\gamma)}{(1-\gamma L)\alpha(L\alpha-\gamma)}.$
Returning to the definition of $u_{t}$ we deduce the following bound for $a_{t}$:
$a_{t}$ $\leq$ $\frac{1}{\gamma}\left(L\alpha-\frac{(1-\gamma L)\alpha(L\alpha-\gamma)}{(1-\gamma L)\alpha-t\gamma(L\alpha-\gamma)}\right)$
$=$ $\frac{1}{\gamma}\frac{L\alpha^{2}(1-\gamma L)-L\alpha t\gamma(L\alpha-\gamma)-(1-\gamma L)\alpha(L\alpha-\gamma)}{(1-\gamma L)\alpha-t\gamma(L\alpha-\gamma)}$
$=$ $\frac{1}{\gamma}\frac{-L\alpha t\gamma(L\alpha-\gamma)+\alpha\gamma(1-\gamma L)}{(1-\gamma L)\alpha-t\gamma(L\alpha-\gamma)}$
$=$ $\alpha\frac{(1-\gamma L)-tL(L\alpha-\gamma)}{(1-\gamma L)\alpha-t\gamma(L\alpha-\gamma)}.$
We note that our Condition 1 requires that $a_{t}\geq 0$ for all $t=0,\ldots,T$, which is equivalent to having $a_{T}\geq 0$ as we assumed $(a_{t})$ to be decreasing. Specifically, we must have
$(1-\gamma L)-TL(L\alpha-\gamma)\geq 0\quad\iff\quad\alpha\leq\frac{1-\gamma L+\gamma TL}{TL^{2}}.$
As we want $\rho$ as large as possible, and due to Condition 2 which reads $\rho\leq 2L(\alpha-\beta)$ with $\beta=0$, we want $\alpha$ to be as large as possible. Therefore we fix
$\alpha=\frac{1-\gamma L+T\gamma L}{TL^{2}}\quad\text{ and }\quad\rho=2L\alpha=\frac{2(1-\gamma L)+2T\gamma L}{TL}.$
This choice respects the assumption $L\alpha>\gamma$ we made earlier, since $1-\gamma L>0$.
To get an expression for $a_{t}$, we simply set the inequalities to equalities, and replace $\alpha$ by its chosen value. We then obtain
$L\alpha-\gamma=\frac{(1-\gamma L)+\gamma LT-\gamma LT}{TL}=\frac{1-\gamma L}{TL}\Rightarrow u_{0}=\frac{TL}{1-\gamma L}.$
So
$u_{t}$ $=$ $u_{0}-tc=\frac{TL}{1-\gamma L}-\frac{t\gamma}{\alpha(1-\gamma L)}=\frac{\alpha TL-t\gamma}{\alpha(1-\gamma L)}=\frac{\tfrac{1}{L}(1-\gamma L)+\gamma T-t\gamma}{\alpha(1-\gamma L)}=\gamma\frac{T-t-1+\tfrac{1}{\gamma L}}{\alpha(1-\gamma L)}.$
We now use the fact that $\gamma a_{t}=L\alpha-u_{t}^{-1}$ to write
$a_{t}$ $=$ $\frac{L\alpha}{\gamma}-\frac{1}{\gamma u_{t}}=\frac{\alpha L}{\gamma}-\frac{1}{L\gamma}\frac{(1-\gamma L)\tfrac{\alpha L}{\gamma}}{T-t-1+\tfrac{1}{\gamma L}}$
$=$ $\frac{\alpha L}{\gamma}\left(1-\frac{1}{\gamma L}\frac{(1-\gamma L)}{T-t-1+\tfrac{1}{\gamma L}}\right)=\frac{\alpha L}{\gamma}\left(1-\frac{(1-\gamma L)}{\gamma L(T-t-1)+1}\right)$
$=$ $\frac{\alpha L}{\gamma}\frac{\gamma L(T-t)}{\gamma L(T-t-1)+1}=\frac{(1-\gamma L)+\gamma LT}{T\gamma L}\frac{\gamma L(T-t)}{\gamma L(T-t-1)+1}$
$=$ $\frac{(T-t)}{T}\frac{1+\gamma L(T-1)}{\gamma L(T-t-1)+1}.$
To conclude the proof, we need to check all the sufficient Lyapunov conditions of Theorem 3.1. Conditions 1, 2, 3 and 5 are already satisfied, by construction. It remains to check Conditions 4 and 6.
Condition 4 is equivalent to
$a_{t+1}\gamma^{2}\leq\alpha$ $\iff$ $\gamma^{2}\frac{(T-t-1)}{T}\frac{1+\gamma L(T-1)}{1+\gamma L(T-t-2)}\leq\frac{(1-\gamma L)+\gamma LT}{TL^{2}}$
$\iff$ $\frac{\gamma^{2}L^{2}(T-t-1)}{1+\gamma L(T-t-2)}(1+\gamma L(T-1))\leq 1+\gamma L(T-1)$
$\iff$ $\gamma^{2}L^{2}(T-t-1)\leq 1+\gamma L(T-t-2)=1+\gamma L(T-t-1)-\gamma L$
$\iff$ $0\leq 1-\gamma L+\gamma L(T-t-1)(1-\gamma L),$
which is true for every $t=0,\dots,T-1$.
Condition 6 is a lower bound on $e_{t}$, of which the equality case coincides with the chosen value of $e_{t}$. More specifically, this gives
$e_{t}=\frac{a_{t+1}\gamma^{2}\alpha}{\alpha-a_{t+1}\gamma^{2}}=\frac{\gamma^{2}}{\frac{1}{a_{t+1}}-\frac{\gamma^{2}}{\alpha}},$
where
$\frac{1}{a_{t+1}}-\frac{\gamma^{2}}{\alpha}$ $=$ $\frac{T}{T-t-1}\frac{(1-\gamma L)+\gamma L(T-t-1)}{(1-\gamma L)+\gamma LT}-\frac{T\gamma^{2}L^{2}}{(1-\gamma L)+\gamma LT}$
$=$ $\frac{T}{T-t-1}\frac{(1-\gamma L)+\gamma L(T-t-1)}{(1-\gamma L)+\gamma LT}-\frac{T\gamma^{2}L^{2}}{(1-\gamma L)+\gamma LT}$
$=$ $T\frac{(1-\gamma L)+\gamma L(T-t-1)-\gamma^{2}L^{2}(T-t-1)}{(T-t-1)((1-\gamma L)+\gamma LT)}$
$=$ $T\frac{(1-\gamma L)((1-\gamma L)+\gamma L(T-t))}{(T-t-1)((1-\gamma L)+\gamma LT)}.$
We thus get
$e_{t}$ $=$ $\frac{\gamma^{2}}{1-\gamma L}\frac{T-t-1}{T}\frac{(1-\gamma L)+\gamma LT}{(1-\gamma L)+\gamma L(T-t)}.$
We readily verify that $e_{t}$ is nonincreasing with respect to $t$. Indeed, its derivative with respect to $t$ has the same sign as
$(T-1)((1-\gamma L)+\gamma LT)\gamma L-((1-\gamma L)+\gamma LT)^{2},$
which is always nonpositive. Therefore $e_{t}\leq e_{0}$, where
$e_{0}=\frac{\gamma^{2}}{1-\gamma L}\frac{T-1}{T}\leq\frac{\gamma^{2}}{1-\gamma L},$
thus concluding the proof. The bound in Theorem 4.1 follows by combining Theorem 3.1 and Lemma 7.
Note that the bound in Theorem 4.1 may be sharpened by using the exact values of $e_{t}$ rather than their bound. Finally, we prove the bias-optimality claim in Theorem 4.1.
Proposition 3(Bias-optimality. Convex case, short step-sizes)
The bias obtained in the bound of Theorem 4.1 is optimal, in the sense that, for any $x_{0}\in\mathcal{H}$, there exists a convex and smooth problem for which $m=1$ and both inequalities are satisfied with equality. Moreover, this problem has bounded gradients.
Proof
We build our lower bound example in a similar to what was done in taylor_smooth_2017 . The idea is to consider Huber functions, which have minimizers while behaving linearly when far away from the minimizers. This linear behavior is not only slow, which fits our goal to find a worst-case function, but also allows for computing easily and explicitly the iterates. Given $\eta>0$, we define the function $\mathcal{H}_{\eta}$ as
$\mathcal{H}_{\eta}(x)=\begin{cases}\eta L\|x\|-\frac{L}{2}\eta^{2}&\text{if}\quad\|x\|>\eta\\ \frac{L}{2}\|x\|^{2}&\text{if}\quad\|x\|\leq\eta.\end{cases}$
This is exactly the classical Huber function multiplied by $L$, which is convex, $L$-smooth and has bounded gradients (bauschke_convex_2017, , Example 13.7). It now remains to prove the claim in item 3. As a preliminary, observe that if $\|x\|\geq\eta$ and if $x^{+}:=x-\gamma\nabla\mathcal{H}_{\eta}(x)$, then
[…]
Let $x_{0}\in\mathcal{H}$ and denote $\|x_{0}\|=D$, and let $(x_{t})$ be the iterates of gradient descent applied to $f=\mathcal{H}_{\eta}$, where the parameter $\eta$ is defined as $\eta=\tfrac{D}{1+(T-1)\gamma L}$. A simple induction argument shows that this choice of $\eta$ is small enough to ensure that $\|x_{t}\|\geq\eta$ for $t=0,\dots,T-1$, which in turn allows us to write $x_{t}=(1-t\tfrac{\gamma L\eta}{D})x_{0},$ and as such
$\overline{x}_{T}$ $=\frac{1}{T}\sum_{t=0}^{T-1}x_{t}=\frac{1}{T}\sum_{t=0}^{T-1}\left(1-t\gamma L\eta/D\right)x_{0}=\left(1-\frac{\gamma L(T-1)\eta}{2D}\right)x_{0}.$
A direct computation shows that $\|\bar{x}_{T}\|\leq\eta$, which means that
$f(\overline{x}_{T})-\inf f$ $=\eta L\left(D-\frac{\gamma L(T-1)\eta}{2}\right)-\frac{L}{2}\eta^{2}=\eta LD-L\eta^{2}\cdot\frac{1+\gamma L(T-1)}{2}=\eta LD-\frac{\eta LD}{2}=\frac{\eta LD}{2}=\frac{L\|x_{0}-x_{*}\|^{2}}{2(1+(T-1)\gamma L)},$
which indeed matches the bias term of Theorem 4.1. Moreover, the iterates $x_{0},\dots,x_{T}$ all belong to the segment $x_{0},x_{T}]$, on which $f$ is affine, which means that we also have $\frac{1}{T}\sum_{t=0}^{T-1}f(x_{t})-\inf f=f(\overline{x}_{T})-\inf f$, thus concluding.
A.1.2 Critical Step-Size: $\gamma L=1$
The proof of Theorem 4.2 is presented in Section A.1.4.
In Remark 4, we mentioned that we believe there to be a singularity at the critical step-size, making it such that the optimal bias cannot be achieved with a finite variance term. Moreover, this was confirmed numerically in Section 7. Theorem 3.1 provides sufficient conditions on the parameters $\left(\rho,(a_{t}),(e_{t})\right)$ for the Lyapunov decrease to happen. Those sufficient conditions are the ones we use throughout the paper, and allow us to theoretically match the bounds obtained numerically. We therefore have good reasons to believe that those sufficient conditions are very close to be necessary. In what follows, we show that the system of inequalities in Theorem 3.1 is not feasible for the optimal value $\rho=2\gamma$ when $\gamma L=1$.
Proposition 4
Assume that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for all $t=0,\dots,T-1$. If $\gamma L=1$, $a_{0}=1$ and $\rho=2\gamma$, then it must hold true that $a_{t}\geq 1$ for all $t=0,\ldots,T$.
Proof
We consider $d=1$ and $m=2$, and let $f\in\mathcal{F}_{0,L}(\mathbb{R})$ be defined through
$f(x)=\frac{1}{2}\left(f_{+}(x)+f_{-}(x)\right)=\frac{1}{2}\left(\frac{L}{2}(x+\delta)^{2}+\frac{L}{2}(x-\delta)^{2}\right).$
One can easily compute that
$f(x)=\frac{L\left(x^{2}+\delta^{2}\right)}{2},\quad\min f=\frac{L\delta^{2}}{2},\quad f(x)-\min f=\frac{Lx^{2}}{2},\quad x_{*}=0$
$\nabla f(x)=Lx,\quad\nabla f_{\pm}(x)=L(x\pm\delta),\quad\nabla f_{\pm}(x_{*})=\pm L\delta,$
We know that $\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}\leq 0$ for all $t=0,\ldots,T-1$ and for all $x_{0}\in\mathbb{R}$. Expanding the Lyapunov decrease and rearranging yields
$a_{t+1}\mathbb{E}_{t}\mathopen{}\left[{|x_{t+1}|^{2}}\right]\mathclose{}-a_{t}|x_{t}|^{2}+\rho\left(f(x_{t})-\min f\right)\leq e_{t}\sigma_{*}^{2}$
$\Longleftrightarrow$ $~a_{t+1}\left((1-\gamma L)^{2}|x_{t}|^{2}+(\gamma L)^{2}\delta^{2}\right)-a_{t}|x_{t}|^{2}+\rho L\left(\frac{|x_{t}|^{2}+\delta^{2}}{2}-\frac{\delta^{2}}{2}\right)\leq 2e_{t}\delta^{2}$
$\Longleftrightarrow$ $\left(a_{t+1}(1-\gamma L)^{2}-a_{t}+\frac{\rho L}{2}\right)|x_{t}|^{2}\leq\left(2e_{t}-a_{t+1}(\gamma L)^{2}\right)\delta^{2}.$
This inequality is to be satisfied for all $\delta$ and $x_{t}$. Specifically, the inequality must hold for $(x_{t},\delta)=(1,0)$, showing that $a_{t+1}(1-\gamma L)^{2}-a_{t}+\rho L/2\leq 0$. Setting $\rho=2\gamma$, and $\gamma L=1$ yields $a_{t}\geq 1$.
Proposition 5
The sufficient conditions of Theorem 3.1 cannot be verified with $\rho=2\gamma$, $a_{0}=1$ and $\gamma L=1$.
Proof
Let $\rho,(e_{t},a_{t})_{t=0}^{T}$ be parameters, normalized with $a_{0}=1$, and suppose $\gamma L=1$ and $\rho=2\gamma$. Suppose that there exists $(\alpha_{t},\beta_{t})$ such that the system of inequality in Theorem 3.1 holds. From Proposition 4, we know that $a_{t}\geq 1$ for all $t=0,\ldots,T$. Moreover, from Condition 3 and a simple induction argument, we deduce that $a_{t}\leq a_{0}=1$. Hence, $a_{t}\equiv 1$. This together with Condition 5 yields
$\left(\frac{1}{L}-\alpha_{t}L\right)^{2}\leq 0\iff\alpha_{t}=\frac{1}{L^{2}}.$
Plugging this value of $\alpha_{t}$ in Condition 2 gives
$\frac{2}{L}\leq 2L\left(\frac{1}{L^{2}}-\beta_{t}\right)\Longrightarrow\beta_{t}\leq 0\Longrightarrow\beta_{t}=0.$
Finally by injecting the values of $\alpha_{t}$ and $\beta_{t}$ into Condition 6, we obtain ${1}/{L^{4}}\leq 0,$ which contradicts $L>0$, thus concluding.
A.1.3 Large Step-Sizes: $\gamma L>1$ (Proof of Theorem 4.3)
Proposition 6(Lyapunov parameters. Convex case, large step-sizes)
Let $1<\gamma L<2$, and consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\bullet$
    $\rho=\tfrac{2\gamma(2-\gamma L)}{1-\theta^{T}}$ where $\theta=(1-\gamma L)^{2}$,
  • $\bullet$
    $a_{t}=\tfrac{1-\theta^{T-t}}{1-\theta^{T}}$,
  • $\bullet$
    $\beta_{t}=\tfrac{\gamma(\gamma L-1)(1-\theta^{T-t-1})}{L(1-\theta^{T})}$,
  • $\bullet$
    $\alpha_{t}=\gamma\tfrac{(2-\gamma L)+(\gamma L-1)(1-\theta^{T-t-1})}{L(1-\theta^{T})}$,
  • $\bullet$
    $e_{t}=\tfrac{\gamma^{2}}{2-\gamma L}\tfrac{1-\theta^{T-t-1}}{1-\theta^{T}}\left(\tfrac{\gamma L}{\theta^{T-t-1}}-2(\gamma L-1)\right)$.
These parameters satisfy the sufficient Lyapunov conditions of Theorem 3.1. Moreover, the averaged variance $\bar{e}_{T}\coloneqq\tfrac{1}{T}\sum_{t=0}^{T-1}e_{t}$ has the following asymptotic equivalence as $T\to+\infty$:
$\bar{e}_{T}\sim\frac{\gamma^{2}}{(2-\gamma L)^{2}}\frac{1}{T(1-\gamma L)^{2T-2}}\to+\infty.$
In particular, the bound in Theorem 4.3 holds true.
Proof
To determine the parameters, we will rely on insights from our numerical experiments. Specifically, we were able to guess that
$\beta_{t}=\frac{a_{t+1}\gamma(\gamma L-1)}{L},$
and that Condition 2 is satisfied with equality, meaning that
$\rho=2L(\alpha_{t}-\beta_{t})\iff\alpha_{t}=\frac{\rho}{2L}+\beta_{t}.$
Considering Condition 5, we have
$(a_{t+1}\gamma-L\alpha_{t})^{2}\leq(\alpha_{t}+\beta_{t}-\gamma^{2}a_{t+1})(a_{t}-a_{t+1}).$
The latter will give us an expression of $a_{t}$ in terms of $a_{t+1}$, from which we will deduce a general expression for $a_{t}$ by induction. By injecting the values of $\alpha_{t}$ and $\beta_{t}$, we obtain
$\frac{1}{L^{2}}\left(\frac{\rho L}{2}-a_{t+1}\gamma L(2-\gamma L)\right)^{2}\leq\frac{1}{L^{2}}\left(\frac{\rho L}{2}-a_{t+1}\gamma L(2-\gamma L)\right)(a_{t}-a_{t+1}).$
This is satisfied when $\rho>2\gamma L(2-\gamma L)$ and
$\frac{\rho L}{2}-a_{t+1}\gamma L(2-\gamma L)\leq a_{t}-a_{t+1}\iff a_{t}\geq\frac{\rho L}{2}+a_{t+1}(\gamma L-1)^{2}.$
We observe an arithmetic-geometric relation which can be summarized after setting $\theta=(\gamma L-1)^{2}$:
$a_{t}\geq\frac{\rho L}{2}+a_{t+1}\theta\iff a_{t+1}\leq\theta^{-1}a_{t}-\frac{\rho L}{2\theta}.$
A simple induction argument leads to
$a_{t}\leq\theta^{-t}a_{0}-\frac{\rho L}{2\theta}\frac{\theta^{-t}-1}{\theta^{-1}-1}.$
As done previously, without loss of generality, we impose $a_{0}=1$. We impose the sequence to be nonincreasing and nonnegative. The latter leads to
$a_{T}\geq 0\iff\frac{\rho L}{2\theta}\frac{\theta^{-T}-1}{\theta^{-1}-1}\leq\theta^{-T}\iff\frac{\rho L}{2}\leq\theta^{-T}\frac{1-\theta}{\theta^{-T}-1}=\frac{1-\theta}{1-\theta^{T}}.$
As we aim to maximize $\rho$, we impose
$\frac{\rho L}{2}=\frac{1-\theta}{1-\theta^{T}}\iff\rho=\frac{2(1-\theta)}{L(1-\theta^{T})}=\frac{2\gamma(2-\gamma L)}{1-\theta^{T}}.$
In particular, we obtain
$a_{t}\leq\theta^{-t}-\frac{\rho L}{2}\frac{\theta^{-t}-1}{1-\theta}=\theta^{-t}-\frac{1-\theta}{1-\theta^{T}}\frac{\theta^{-t}-1}{1-\theta}=\theta^{-t}-\frac{\theta^{-t}-1}{1-\theta^{T}}=\frac{1-\theta^{T-t}}{1-\theta^{T}}.$
To enforce that $(a_{t})$ is nonincreasing, we set the above to be an equality, and since $a_{T}\geq 0$, it is also nonnegative, as wanted.
We now verify the remaining conditions from Theorem 3.1. The nonnegativity of $a_{t}$ implies the nonnegativity of $\alpha_{t},\beta_{t}$. Conditions 1, 3, 5 are verified by construction. It remains to check that Conditions 4 and 6 hold true.
In the following, we introduce $k\coloneqq T-t-1$ to simplify notation. To check that Condition 4 is satisfied, we compute $\alpha_{t}+\beta_{t}$:
$\alpha_{t}+\beta_{t}=\frac{\rho}{2L}+2\beta_{t}=\frac{\gamma(2-\gamma L)+2\gamma(\gamma L-1)(1-\theta^{k})}{L(1-\theta^{T})}.$
Therefore, Condition 4, which states $\alpha_{t}+\beta_{t}-\gamma^{2}a_{t+1}\geq 0$, becomes
$0$ $\leq$ $\frac{\gamma(2-\gamma L)+2\gamma(\gamma L-1)(1-\theta^{k})}{L(1-\theta^{T})}-\gamma^{2}\frac{1-\theta^{k}}{1-\theta^{T}}$
$=$ $\frac{\gamma(2-\gamma L)+2\gamma(\gamma L-1)(1-\theta^{k})-\gamma^{2}L(1-\theta^{k})}{L(1-\theta^{T})}$
$=$ $\frac{\gamma(2-\gamma L)-(1-\theta^{k})\gamma(2-\gamma L)}{L(1-\theta^{T})}$
$=$ $\frac{\gamma(2-\gamma L)\theta^{k}}{L(1-\theta^{T})},$
which is true because $\gamma L\in(1,2)$. It remains to study Condition 6, which is satisfied by setting
$e_{t}$ $=$ $\gamma^{2}\frac{a_{t+1}(\alpha_{t}+\beta_{t})}{\alpha_{t}+\beta_{t}-\gamma^{2}a_{t+1}}$
$=$ $\gamma^{2}\frac{1-\theta^{k}}{1-\theta^{T}}\frac{\gamma(2-\gamma L)+2\gamma(\gamma L-1)(1-\theta^{k})}{L(1-\theta^{T})}\frac{L(1-\theta^{T})}{\gamma(2-\gamma L)\theta^{k}}$
$=$ $\gamma^{2}\frac{1-\theta^{k}}{1-\theta^{T}}\frac{(2-\gamma L)+2(\gamma L-1)(1-\theta^{k})}{(2-\gamma L)\theta^{k}}$
$=$ $\gamma^{2}\frac{1-\theta^{k}}{1-\theta^{T}}\frac{\gamma L-2(\gamma L-1)\theta^{k}}{(2-\gamma L)\theta^{k}}$
$=$ $\frac{\gamma^{2}}{2-\gamma L}\frac{1-\theta^{k}}{1-\theta^{T}}\left(\frac{\gamma L}{\theta^{k}}-2(\gamma L-1)\right).$
To conclude the proof, it remains to analyze the average of the $e_{t}$’s. Denote $e_{t}^{\prime\prime}=({1-\theta^{k}})\left(\frac{\gamma L}{\theta^{k}}-2(\gamma L-1)\right)$, such that
$\sum_{t=0}^{T-1}e_{t}^{\prime\prime}$ $=$ $\gamma L\sum_{t=0}^{T-1}\frac{1-\theta^{k}}{\theta^{k}}-2(\gamma L-1)\sum_{t=0}^{T-1}(1-\theta^{k})$
$=$ $\gamma L\sum_{t=0}^{T-1}(\theta^{-k}-1)-2(\gamma L-1)\sum_{t=0}^{T-1}(1-\theta^{k})$
$=$ $\gamma L\left(\sum_{t=0}^{T-1}\theta^{-k}\right)-T\gamma L-2(\gamma L-1)T+2(\gamma L-1)\left(\sum_{t=0}^{T-1}\theta^{k}\right)$
$=$ $\gamma L\left(\sum_{t=0}^{T-1}\theta^{-t}\right)-T(3\gamma L-2)+2(\gamma L-1)\left(\sum_{t=0}^{T-1}\theta^{t}\right)$
$=$ $\gamma L\frac{1-\theta^{-T}}{1-\theta^{-1}}-T(3\gamma L-2)+2(\gamma L-1)\frac{1-\theta^{T}}{1-\theta}$
$=$ $\frac{\theta}{2-\gamma L}(\theta^{-T}-1)-T(3\gamma L-2)+\frac{2(\gamma L-1)}{\gamma L(2-\gamma L)}(1-\theta^{T}),$
which is asymptotically equivalent to $\tfrac{\theta}{2-\gamma L}\theta^{-T}$. Therefore we can write
$\bar{e}_{T}=\frac{\gamma^{2}}{2-\gamma L}\frac{1}{1-\theta^{T}}\frac{1}{2-\gamma L}\bar{e}_{T}^{\prime},$
where $e_{t}^{\prime}=(2-\gamma L)e_{t}^{\prime\prime}$ and
$\bar{e}_{T}^{\prime}=\frac{1}{T}\sum_{t=0}^{T-1}e_{t}^{\prime}=\frac{\theta}{T}(\theta^{-T}-1)-(2-\gamma L)(3\gamma L-2)+\frac{2(\gamma L-1)}{\gamma LT}(1-\theta^{T}).$ (6)
In particular, one sees that $\bar{e}_{T}^{\prime}\sim\frac{1}{T\theta^{T-1}}$, where we recall that we defined $\theta=(\gamma L-1)^{2}\in(0,1)$. The bound in Theorem 4.3 now follows by combining Theorem 3.1 and Lemma 7.
We may moreover show the bias-optimality claim of Theorem 4.3 by providing a function that achieves the lower bound.
Proposition 7(Bias-optimality. Convex case, large step-sizes)
The bias obtained in the bound of Theorem 4.3 is optimal, in the sense that, for any $x_{0}\in\mathcal{H}$, there exists a convex and smooth problem for which $m=1$ and the inequality is satisfied with equality.
Proof
Consider the function $f(x)=\tfrac{L}{2}\|x\|^{2}$ with minimal value $\inf f=0$ and minimizer $x_{*}=\mathbf{0}$, and take any $x_{0}\in\mathcal{H}$. The iterates $(x_{t})$ of gradient descent are then given by $x_{t}=(1-\gamma L)^{t}x_{0}$, such that
$\frac{1}{T}\sum_{t=0}^{T-1}f(x_{t})-\inf f$ $=\frac{L}{2T}\sum_{t=0}^{T-1}(1-\gamma L)^{2t}\cdot\|x_{0}\|^{2}=\frac{1-(1-\gamma L)^{2T}}{2\gamma(2-\gamma L)T}\cdot\|x_{0}-x_{*}\|^{2}.$
This indeed matches the bias term in Theorem 4.3.
A.1.4 General Step-Size, Suboptimal Bias (Proof of Theorems 4.2 and 4.4)
We now provide a proof for our suboptimal bounds, which cover at the same time the critical and large step-size regimes.
Theorem A.1(Convex case, any step-size, suboptimal bounds)
Let $\gamma L\in(0,2)$ and $\varepsilon\in(0,1)$, and consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\bullet$
    $\rho=2\gamma(1-\varepsilon)\eta$ where $\eta=1-(\gamma L-1)_{+}$,
  • $\bullet$
    $a_{t}=1$,
  • $\bullet$
    $\alpha_{t}=\gamma/L$,
  • $\bullet$
    $\beta_{t}=\alpha_{t}-\tfrac{\rho}{2L}$,
  • $\bullet$
    $e_{t}=\gamma^{2}\frac{\alpha_{t}+\beta_{t}}{\alpha_{t}+\beta_{t}-\gamma^{2}}$.
These parameters satisfy the sufficient Lyapunov conditions of Theorem 3.1. Moreover, the following simplified bound
$\frac{\bar{e}_{T}}{\rho}\leq\gamma\cdot\frac{\gamma L+\varepsilon\eta}{2(1-\varepsilon)\varepsilon\eta^{2}}$
holds true. In particular, Theorems 4.2 and 4.4 hold true.
Proof
The goal is to find parameters satisfying the sufficient Lyapunov conditions of Theorem 3.1 with a suboptimal value for $\rho$. Therefore we start by defining $\rho=2\gamma(1-\varepsilon)\eta$ for some $\varepsilon\in(0,1)$ and where $\eta=1$ or $2-\gamma L$ depending whether $\gamma L\leq 1$ or $\gamma L\geq 1$. Taking inspiration from our numerical findings, we allow ourselves, in this suboptimal case, to take a sequence $a_{t}$ constantly equal to $1$, which automatically satisfies Condition 3. By Condition 5, we must have $\alpha_{t}=\gamma/L$. We note that Conditions 2 and 4 respectively provide upper and lower bounds on $\beta_{t}$. In view of satisfying Condition 6 with the smallest possible value for $e_{t}$, we aim to pick $\beta_{t}$ to be as large as possible. We therefore saturate the upper bound on $\beta_{t}$ in Condition 2 and take
$\beta_{t}=\alpha_{t}-\frac{\rho}{2L}=\frac{\gamma}{L}-\frac{2\gamma(1-\varepsilon)\eta}{2L}=\frac{\gamma}{L}\left[1-\eta+\eta\varepsilon\right].$
This choice is compatible with Condition 4 as
$\gamma^{2}\leq\alpha_{t}+\beta_{t}\iff\gamma L-1\leq 1-\eta+\eta\varepsilon,$
which is true since
$\gamma L-1\leq(\gamma L-1)_{+}=1-\eta\leq 1-\eta+\eta\varepsilon.$
Finally, the choice of $e_{t}$ is such that Condition 6 is an equality, and hence also satisfied. We note that Condition 1 is verified, as $\beta_{t}\geq 0$ since $\rho\leq 2\gamma$, and the other parameters are naturally nonnegative.
To derive the simplified bound it suffices to compute
$\frac{\bar{e}_{T}}{\rho}=\frac{\gamma^{2}}{\rho}\frac{2-\eta(1-\varepsilon)}{2-\gamma L-\eta(1-\varepsilon)}=\frac{\gamma}{2(1-\varepsilon)\eta}\frac{2-\eta(1-\varepsilon)}{2-\gamma L-\eta+\eta\varepsilon}.$
We may simplify the second fraction of this variance term, by considering two cases;
  • $\bullet$
    if $\gamma L\geq 1$, then $\eta=2-\gamma L$ and
    $\frac{2-\eta(1-\varepsilon)}{2-\gamma L-\eta+\eta\varepsilon}=\frac{2-(2-\gamma L)(1-\varepsilon)}{\eta\varepsilon}=\frac{\gamma L+(2-\gamma L)\varepsilon}{\eta\varepsilon}=\frac{\gamma L+\eta\varepsilon}{\eta\varepsilon}.$
  • $\bullet$
    if $\gamma L\leq 1$, then $\eta=1$ and
    $\frac{2-\eta(1-\varepsilon)}{2-\gamma L-\eta+\eta\varepsilon}=\frac{1+\varepsilon}{1-\gamma L+\varepsilon}.$
    The above fraction can be seen as a convex function $\phi(\gamma L)$ of $\gamma L\in(0,1)$. As such we can write that $\phi(\gamma L)\leq\phi(0)+\gamma L(\phi(1)-\phi(0))$ where $\phi(0)=1$ and $\phi(1)=\tfrac{1+\varepsilon}{\varepsilon}$, which in our case implies that
    $\frac{1+\varepsilon}{1-\gamma L+\varepsilon}\leq\frac{\gamma L+\varepsilon}{\varepsilon}=\frac{\gamma L+\varepsilon\eta}{\varepsilon\eta}.$
In both cases the second fraction of our variance term is bounded by $\tfrac{\gamma L+\varepsilon\eta}{\varepsilon\eta}$, which gives the desired bound.
Theorems 4.2 and 4.4 now follow by combining Theorem 3.1 and Lemma 7, and noting that $\eta=1$ and $\eta=2-\gamma L$ when $\gamma L=1$ and $\gamma L\geq 1$, respectively.
A.2 Strongly Convex Setting
As in the previous section, we derive parameters that satisfy the sufficient conditions of Theorem 3.1 in order to derive upper bounds for SGD, this time in the strongly convex setting. We again note that all propositions could simply be verified symbolically, yet we include the proofs to provide intuition on how these parameters were derived.
Because it seems impossible to obtain a bound with tight bias term for critical step-sizes (see Sections 7), we provide here a more relaxed bound, where the bias term is $\varepsilon$ away from being tight, and subsequently the variance term remains bounded for every step-size $\gamma L\in(0,2)$.
Proposition 8(Lyapunov parameters. Strongly convex case, any step-size, $L>\mu$)
Let $\gamma L\in(0,2)$, with $L>\mu$. Let $\varepsilon\geq 0$ and $\gamma_{\operatorname{crit}}=\tfrac{2}{\mu+L}$. Assume that either $\varepsilon>0$ or $\gamma\neq\gamma_{\operatorname{crit}}$. Consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\bullet$
    $\rho=0$,
  • $\bullet$
    $a_{t}=\phi^{2(T-t)}$, where $\phi^{2}=\phi_{{\operatorname{opt}}}^{2}+\varepsilon$, with
    $\phi_{{\operatorname{opt}}}=\max\{1-\gamma\mu;\gamma L-1\}\quad\text{ and }\quad\varepsilon\in[0,1-\phi^{2}_{{\operatorname{opt}}}),$
  • $\bullet$
    $\alpha_{t}=\beta_{t}=\alpha a_{t+1}$, where $\alpha=\omega+\omega_{\varepsilon}$ and
    $\omega=\frac{\gamma\phi_{{\operatorname{opt}}}}{L-\mu}\quad\text{ and }\quad\omega_{\varepsilon}=\frac{\varepsilon+\sqrt{\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta}}{(L-\mu)^{2}}\quad\text{ with }\quad\delta=|\gamma-\gamma_{\operatorname{crit}}|,$
  • $\bullet$
    $e_{t}=ea_{t+1}$, where
    $e=\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)}{2\omega_{\varepsilon}(L-\mu)+({L+\mu})\gamma\delta}\right).$
Then the sufficient Lyapunov conditions of Theorem 3.1 are verified. Moreover, we have
$e_{T}^{sum}\coloneqq\sum_{t=0}^{T-1}e_{t}=e\frac{1-\phi^{2T}}{1-\phi^{2}}\leq\frac{e}{1-\phi^{2}},$
and
$e\leq\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)^{2}}{4\varepsilon+(L+\mu)(L-\mu)\gamma\delta}\right).$
Proof
We start the proof by exploiting some insights from our numerical analysis. First, we observe that we can impose $\rho=0$ and $\alpha_{t}=\beta_{t}$ without changing anything to the empirical best bias term. We moreover infer that ${a_{t}}/{a_{t+1}}\in(0,1)$ is constant, so let us denote this ratio by $\phi^{2}$. Since $\phi_{{\operatorname{opt}}}^{2}$ is the optimal rate for the deterministic gradient descent in the strongly convex smooth case, we anticipate that $\phi^{2}\geq\phi_{\operatorname{opt}}^{2}$. Without loss of generality, we impose $a_{T}=1$, from which we deduce that $a_{t}=\phi^{2(T-t)}$. We were also able to deduce that $e_{t}=ea_{t+1}$ and that $\alpha_{t}=\alpha a_{t+1}$ for some constants $e,\alpha>0$, yet to be determined.
It remains to study the conditions of Theorem 3.1 to find feasible values for $\alpha$ and $e$.
Condition 6 yields
$a_{t+1}\gamma^{2}(2\alpha a_{t+1})\leq(2\alpha a_{t+1}-a_{t+1}\gamma^{2})ea_{t+1},$
which holds true if
$2\gamma^{2}\alpha\leq(2\alpha-\gamma^{2})e.$
As we want to minimize $e$, we set
$e=\frac{2\gamma^{2}\alpha}{2\alpha-\gamma^{2}},$
and impose that the denominator $2\alpha-\gamma^{2}$ is positive. Now it remains to find $\alpha$. The above expression is decreasing with $\alpha$, so we aim to find the largest $\alpha$ possible such that all conditions are satisfied.
Turning to Condition 5 will determine $\alpha$. With our current notations and restrictions on the parameters, it can be written as
$(a_{t+1}\gamma-\alpha a_{t+1}(L+\mu))^{2}\leq(2\mu L\alpha a_{t+1}+a_{t}-a_{t+1})(2\alpha a_{t+1}-a_{t+1}\gamma^{2}),$
which is true if
$(\gamma-\alpha(L+\mu))^{2}-(2\mu L\alpha-(1-\phi^{2}))(2\alpha-\gamma^{2})\leq 0.$
This is a polynomial of degree at most $2$ in $\alpha$, which must be nonpositive. Expanding the terms to find the coefficients of this polynomial $P(\alpha)$, we obtain
$P(\alpha)$ $=$ $\gamma^{2}+\alpha^{2}(L+\mu)^{2}-2\alpha(L+\mu)\gamma-4\mu L\alpha^{2}+2\mu L\alpha\gamma^{2}+2\alpha(1-\phi^{2})-\gamma^{2}(1-\phi^{2})$
$=$ $\alpha^{2}\left[(L+\mu)^{2}-4\mu L\right]-\alpha\left[2\gamma(L+\mu)-2\mu L\gamma^{2}-2(1-\phi^{2})\right]+\left[\gamma^{2}-\gamma^{2}(1-\phi^{2})\right]$
$=$ $a\alpha^{2}-2b\alpha+c,$
where
$\begin{cases}a&=(L+\mu)^{2}-4\mu L=(L-\mu)^{2},\\ b&=\gamma(L+\mu)-\mu L\gamma^{2}-(1-\phi^{2}),\\ c&=\gamma^{2}\phi^{2}=\gamma^{2}(\phi_{{\operatorname{opt}}}^{2}+\varepsilon).\end{cases}$ (7)
It is a simple exercise to verify that $\phi_{{\operatorname{opt}}}=1-\mu\gamma$ if $\gamma\leq\gamma_{\operatorname{crit}}$ and $\phi_{{\operatorname{opt}}}=\gamma L-1$ if $\gamma\geq\gamma_{\operatorname{crit}}$, which leads us to compute the value of $b$ through a case distinction:
  • $\bullet$
    if $\gamma\leq\gamma_{\operatorname{crit}}$ then $1-\phi_{{\operatorname{opt}}}^{2}=\gamma\mu(2-\gamma\mu)$ and
    $b$ $=$ $\gamma(L+\mu)-\mu L\gamma^{2}-\gamma\mu(2-\gamma\mu)+\varepsilon$
    $=$ $\gamma\left(L+\mu-\mu L\gamma-2\mu+\gamma\mu^{2}\right)+\varepsilon$
    $=$ $\gamma(1-\gamma\mu)(L-\mu)+\varepsilon$
    $=$ $\gamma\phi_{{\operatorname{opt}}}(L-\mu)+\varepsilon.$
  • $\bullet$
    if $\gamma\geq\gamma_{\operatorname{crit}}$ then $1-\phi_{{\operatorname{opt}}}^{2}=\gamma L(2-\gamma L)$ and
    $b$ $=$ $\gamma(L+\mu)-\mu L\gamma^{2}-\gamma L(2-\gamma L)+\varepsilon$
    $=$ $\gamma\left(L+\mu-\mu L\gamma-2L+\gamma L^{2}\right)+\varepsilon$
    $=$ $\gamma(\gamma L-1)(L-\mu)+\varepsilon$
    $=$ $\gamma\phi_{{\operatorname{opt}}}(L-\mu)+\varepsilon.$
In both cases we obtain $b=\gamma\phi_{{\operatorname{opt}}}(L-\mu)+\varepsilon\geq 0$.
Since $L>\mu$, we have that $a>0$, so the largest feasible $\alpha$ is the largest root of $P$, if existent. The discriminant $\Delta$ of $P$ is given by
$\Delta$ $=$ $(2b)^{2}-4ac$
$=$ $4\left(\gamma\phi_{{\operatorname{opt}}}(L-\mu)+\varepsilon\right)^{2}-4(L-\mu)^{2}\gamma^{2}(\phi^{2}_{{\operatorname{opt}}}+\varepsilon)$
$=$ $4\gamma^{2}\phi_{{\operatorname{opt}}}^{2}(L-\mu)^{2}+4\varepsilon^{2}+8\varepsilon\gamma\phi_{{\operatorname{opt}}}(L-\mu)-4(L-\mu)^{2}\gamma^{2}\phi^{2}_{{\operatorname{opt}}}-4(L-\mu)^{2}\gamma^{2}\varepsilon$
$=$ $4\left[\varepsilon^{2}+\varepsilon\left(2\gamma\phi_{{\operatorname{opt}}}(L-\mu)-(L-\mu)^{2}\gamma^{2}\right)\right]$
$=$ $4\left[\varepsilon^{2}+\varepsilon\gamma(L-\mu)\left(2\phi_{{\operatorname{opt}}}-(L-\mu)\gamma\right)\right].$
A simple calculation shows that
$2\phi_{{\operatorname{opt}}}-(L-\mu)\gamma=\begin{cases}2-\gamma(L+\mu)&\text{ if }\gamma\leq\gamma_{\operatorname{crit}}\\ -2+\gamma(L+\mu)&\text{ if }\gamma\geq\gamma_{\operatorname{crit}}\end{cases}=|2-\gamma(L+\mu)|=(L+\mu)|\gamma-\gamma_{\operatorname{crit}}|.$
If we introduce $\delta\coloneq|\gamma-\gamma_{\operatorname{crit}}|$ then we have the simpler and nonnegative expression
$\Delta=4\left[\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta\right]\geq 0.$
As such, $P$ has real roots. We thus define $\alpha$ to be the largest of the two roots
$\alpha$ $=$ $\frac{2b+\sqrt{\Delta}}{2a}$
$=$ $\frac{\gamma\phi_{{\operatorname{opt}}}(L-\mu)+\varepsilon+\sqrt{\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta}}{(L-\mu)^{2}}$
$=$ $\omega+\omega_{\varepsilon},$
where we noted
$\omega=\frac{\gamma\phi_{{\operatorname{opt}}}}{L-\mu}\quad\text{ and }\quad\omega_{\varepsilon}=\frac{\varepsilon+\sqrt{\varepsilon^{2}+\varepsilon\gamma(L-\mu)(L+\mu)\delta}}{(L-\mu)^{2}}.$
To conclude the proof, we are now going to verify that this choice of $\alpha$ is feasible, in the sense that it satisfies all the conditions from Theorem 3.1. The nonnegativity Condition 1 is satisfied by construction. Condition 2 is satisfied as we set $\rho=0$ and $\alpha_{t}=\beta_{t}$. Condition 3 is equivalent to
$1\leq 2\alpha\mu L+\frac{a_{t}}{a_{t+1}}=2\alpha\mu L+\phi^{2}.$
We have $\alpha=\omega+\omega_{\varepsilon}\geq\omega$ and $\phi^{2}\geq\phi^{2}_{{\operatorname{opt}}}$, and the above is thus true if
$1\leq\mu L2\omega+\phi^{2}_{{\operatorname{opt}}}$
$\iff$ $1\leq\frac{2\mu L\gamma\phi_{{\operatorname{opt}}}}{L-\mu}+\phi^{2}_{{\operatorname{opt}}}$
$\iff$ $0\leq 2\mu L\gamma\phi_{{\operatorname{opt}}}-(1-\phi^{2}_{{\operatorname{opt}}})(L-\mu).$
To assess whether this inequality is true, we are going to consider two cases:
  • $\bullet$
    if $\gamma\leq\gamma_{\operatorname{crit}}$, then $\phi_{{\operatorname{opt}}}=1-\gamma\mu$, so the inequality becomes
    $0$ $\leq$ $2\mu L\gamma(1-\gamma\mu)-\gamma\mu(2-\gamma\mu)(L-\mu)$
    $=$ $\mu\gamma(-\mu\gamma L-\gamma\mu^{2}+2\mu)$
    $=$ $\mu^{2}\gamma(2-\gamma(L+\mu))$
    $=$ $\mu^{2}\gamma(\mu+L)\delta.$
  • $\bullet$
    if $\gamma\geq\gamma_{\operatorname{crit}}$, then $\phi_{{\operatorname{opt}}}=\gamma L-1$, so the inequality becomes
    $0$ $\leq$ $2\mu L\gamma(\gamma L-1)-\gamma L(2-\gamma L)(L-\mu)$
    $=$ $L\gamma(\mu\gamma L+\gamma L^{2}-2L)$
    $=$ $-L^{2}\gamma(2-\gamma(L+\mu))$
    $=$ $L^{2}\gamma(\mu+L)\delta.$
In both cases the inequality is verified, therefore Condition 3 holds true. Let us now turn to Condition 4, which, in our context, is equivalent to
$0\leq 2\alpha-\gamma^{2},$
which holds true as
$2\alpha-\gamma^{2}$ $=$ $2\omega_{\varepsilon}+2\omega-\gamma^{2}=2\omega_{\varepsilon}+\frac{\gamma}{L-\mu}\left(2\phi_{{\operatorname{opt}}}-\gamma(L-\mu)\right)$
$=$ $2\omega_{\varepsilon}+\frac{\gamma}{L-\mu}(L+\mu)\delta.$
Condition 5 is true by the choice of $\alpha$. Finally, Condition 6 holds true provided $2\alpha-\gamma^{2}>0$, which is verified through the above computation provided $\varepsilon>0$ or $\delta\neq 0$. In this case,
$e$ $=$ $\gamma^{2}\frac{2\omega+2\omega_{\varepsilon}}{2\omega_{\varepsilon}+\frac{L+\mu}{L-\mu}\gamma\delta}$
$=$ $\gamma^{2}\frac{2\gamma\phi_{{\operatorname{opt}}}+2(L-\mu)\omega_{\varepsilon}}{2\omega_{\varepsilon}(L-\mu)+({L+\mu})\gamma\delta}$
$=$ $\gamma^{2}\frac{\gamma((L+\mu)\delta+\gamma(L-\mu))+2(L-\mu)\omega_{\varepsilon}}{2\omega_{\varepsilon}(L-\mu)+({L+\mu})\gamma\delta}$
$=$ $\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)}{2\omega_{\varepsilon}(L-\mu)+({L+\mu})\gamma\delta}\right).$
The bound on $e_{T}^{sum}$ is a direct consequence of the fact that $a_{t}$ is a geometric sequence. The upper bound for $e$ comes from the lower bound $\omega_{\varepsilon}\geq\frac{2\varepsilon}{(L-\mu)^{2}}$.
We note the above proposition only holds true for $\mu<L$, thus not covering the case $L=\mu$. We cover this case in a separate proposition.
Proposition 9(Lyapunov parameters. Strongly convex case, any step-size, $L=\mu$)
Let $\gamma L\in(0,2)$, with $L=\mu$. Let $\varepsilon>0$. Consider the parameters $\rho,a_{t},e_{t},\alpha_{t},\beta_{t}$ defined by
  • $\bullet$
    $\rho=0$,
  • $\bullet$
    $a_{t}=\phi^{2(T-t)}$, where $\phi^{2}=\phi_{{\operatorname{opt}}}^{2}+\varepsilon$, with
    $\phi_{{\operatorname{opt}}}=\max\{1-\gamma\mu;\gamma L-1\}\quad\text{ and }\quad\varepsilon\in[0,1-\phi^{2}_{{\operatorname{opt}}}),$
  • $\bullet$
    $\alpha_{t}=\beta_{t}$ and $\alpha_{t}=\alpha a_{t+1}$ where
    $\alpha=\max\left\{\frac{\gamma^{2}}{2}\left(1+\frac{\phi_{{\operatorname{opt}}}^{2}}{\varepsilon}\right),\frac{1-\phi^{2}_{{\operatorname{opt}}}-\varepsilon}{2L^{2}}\right\},$
  • $\bullet$
    $e_{t}=ea_{t+1}$ where
    $e=\frac{2\gamma^{2}\alpha}{2\alpha-\gamma^{2}}.$
Then the sufficient Lyapunov conditions of Theorem 3.1 are verified. Moreover, we have
$\phi\xrightarrow{\varepsilon\to 0}\phi_{{\operatorname{opt}}},\quad\alpha\xrightarrow{\varepsilon\to 0}+\infty,\quad e\xrightarrow{\varepsilon\to 0}\gamma^{2}.$
Proof
The arguments in this proof are mostly recycled from Proposition 8. In particular, we shall use the same values for $\rho$ and $a_{t}$. Regarding the value of $\alpha$, we impose $a\alpha^{2}-2b\alpha+c\leq 0$ where $a,b,c$ are defined in Equation (7), which we recall for convenience:
$\begin{cases}a&=(L-\mu)^{2},\\ b&=\gamma(L+\mu)-\mu L\gamma^{2}-(1-\phi^{2}),\\ c&=\gamma^{2}(\phi_{{\operatorname{opt}}}^{2}+\varepsilon).\end{cases}$
Because $L=\mu$, it holds that $a=0$, and that
$b=2\gamma L-L^{2}\gamma^{2}-(1-\phi_{{\operatorname{opt}}}^{2})+\varepsilon=2\gamma L-L^{2}\gamma^{2}-\gamma L(2-\gamma L)+\varepsilon=\varepsilon.$
Therefore, $\alpha$ must satisfy $c\leq 2b\alpha$, that is
$\alpha\geq\frac{c}{2b}=\frac{\gamma^{2}(\phi_{{\operatorname{opt}}}^{2}+\varepsilon)}{2\varepsilon}=\frac{\gamma^{2}}{2}\left(1+\frac{\phi_{{\operatorname{opt}}}^{2}}{\varepsilon}\right).$ (8)
It then remains to verify the conditions of Theorem 3.1. Conditions 1 and 2 are readily satisfied. Condition 3 can be rewritten as
$\alpha\geq\frac{1-\phi^{2}_{{\operatorname{opt}}}-\varepsilon}{2L^{2}},$
thus justifying our definition of $\alpha$. Condition 4 requires that $\alpha\geq\tfrac{\gamma^{2}}{2}$, which is already true because of Equation (8). Condition 5 is satisfied under Equation (8). Finally, Condition 6 is satisfied with
$e=\frac{2\gamma^{2}\alpha}{2\alpha-\gamma^{2}},$
where the denominator is strictly positive because of Equation (8).
Theorem A.2(Strongly convex case, general step-size)
Let Assumptions 1 and 2 hold, with $\mu>0$. Let $\varepsilon\geq 0$ and $\gamma_{\operatorname{crit}}=\tfrac{2}{\mu+L}$. Let $x_{t}$ be generated by (SGD), with $\gamma L\in(0,2)$. Assume that either $\varepsilon>0$ or $\gamma\neq\gamma_{\operatorname{crit}}$. Then, for every $T\geq 1$,
$\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\phi^{2T}\cdot\|x_{0}-x_{*}\|^{2}+\frac{1-\phi^{2T}}{1-\phi^{2}}e\sigma_{*}^{2},$
where $\phi^{2}=\varepsilon+(\max\{1-\gamma\mu;\gamma L-1\})^{2}\in[0,1)$, and $e$ is defined in Proposition 8 and verifies
$e\leq\gamma^{2}\left(1+\frac{\gamma^{2}(L-\mu)^{2}}{4\varepsilon+(L+\mu)(L-\mu)\gamma|\gamma-\gamma_{\operatorname{crit}}|}\right)=\gamma^{2}\frac{\gamma^{2}_{\operatorname{crit}}\varepsilon+\gamma\gamma_{\operatorname{crit}}\phi_{\gamma}\phi_{\operatorname{crit}}}{\gamma^{2}_{\operatorname{crit}}\varepsilon+\gamma\phi_{\operatorname{crit}}|\gamma-\gamma_{\operatorname{crit}}|}.$
In particular, the bounds of Theorems 5.1 and 5.2 hold true.
Proof
We separate the proof into two cases, namely the case $L>\mu$ and the case $L=\mu$.
  • $\bullet$
    Consider $L>\mu$. Proposition 8, combined with Theorem 3.1, provides parameters ensuring that the Lyapunov energy decreases with, in particular, $a_{T}=1$, $a_{0}=\phi^{2T}$. We deduce the wanted bound from Lemma 8.
  • $\bullet$
    Consider $L=\mu$. Let us consider first $\varepsilon>0$. Combining Lemma 8, Theorem 3.1 and Proposition 9 we obtain
    $\mathbb{E}\mathopen{}\left[{\|x_{T}-x_{*}\|^{2}}\right]\mathclose{}\leq\phi^{2T}\cdot\|x_{0}-x_{*}\|^{2}+\frac{1-\phi^{2T}}{1-\phi^{2}}e_{\varepsilon}\sigma_{*}^{2},$
    where $e_{\varepsilon}$ converges to $\gamma^{2}$ and $\phi\to\phi_{\operatorname{opt}}$ when $\varepsilon\to 0$, thus yielding the wanted bound when $\varepsilon=0$. If we wish to prove the claim for $\varepsilon>0$, it is enough to realize that $\phi_{{\operatorname{opt}}}\leq\phi$, together with the fact that $\tfrac{1-\phi^{2T}}{1-\phi^{2}}$ is nondecreasing with $\phi$.
The bounds of Theorems 5.1 and 5.2 now follow by either setting $\varepsilon=0$ or using that $\gamma=\gamma_{\operatorname{crit}}$.
We end by showing the bias-optimality claim of Theorem 5.1.
Proposition 10(Bias-optimality. Strongly convex case)
The bias obtained in the bound of Theorem 5.1 is optimal, in the sense that there exists a convex and smooth problem for which $m=1$ and the inequality is satisfied with equality.
Proof
It suffices to consider a one-dimensional quadratic function given by $f(x)=\frac{\mu}{2}\|x\|^{2}$ if $\gamma\leq\frac{2}{L+\mu}$, and by $f(x)=\frac{L}{2}\|x\|^{2}$ otherwise, and any $x_{0}\in\mathbb{R}^{d}$. The iterates $(x_{t})$ of gradient descent are given by $x_{t}=(1-\gamma\mu)^{t}x_{0}$ or $x_{t}=(1-\gamma L)^{t}x_{0}$. Noting that $x_{*}=0$ in either case, we obtain that
$\|x_{T}-x_{*}\|^{2}=\begin{cases}(1-\gamma\mu)^{2T}\cdot\|x_{0}-x_{*}\|^{2}&\quad\text{if $\gamma\leq\frac{2}{L+\mu}$}\\ (1-\gamma L)^{2T}\cdot\|x_{0}-x_{*}\|^{2}&\quad\text{if $\gamma\geq\frac{2}{L+\mu}$}\end{cases}=\max(1-\gamma\mu,\gamma L-1)^{2T}\cdot\|x_{0}-x_{*}\|^{2},$
matching the bias term obtained in Theorem 5.1.
Appendix B Performance Estimation Problem Methodology
In this section, we describe the Performance Estimation Problem (PEP) methodology in more detail, and explain how we made use of it in this scenario. The methodology was initially introduced by drori_performance_2014 , further improved by taylor_smooth_2017 , and adapted by taylor_stochastic_2019 in the context of stochastic algorithms, that we will follow closely.
Throughout this section, we define $\mathcal{F}_{\mu,L}(\mathbb{R}^{d})$ to be the set of $L$-smooth $\mu$-strongly convex functions. We shall often drop the dependency in $\mathbb{R}^{d}$ when the space is clear from context.
B.1 Problem Reformulation
As we aim to obtain a computationally tractable problem, we do restrict ourselves to the finite-sum setting in this section. We do however note that the number of functions $m$ does not influence the final result, and can hence theoretically be taken to be infinite, thus recovering the results of Sections A.1 and A.2. In the same spirit, we here assume $\mathcal{H}$ to be a finite-dimensional euclidean space, although the proofs of our results directly extend to any real Hilbert space.
As explained by Lemmas 7 and 8, our goal is to derive nonnegative parameters $\rho,(a_{t},e_{t})$ such that $\mathbb{E}\mathopen{}\left[{E_{t+1}}\right]\mathclose{}\leq\mathbb{E}\mathopen{}\left[{E_{t}}\right]\mathclose{}$ for all $t=0,\ldots,T-1$, where
$E_{t}=a_{t}\|x_{t}-x_{*}\|^{2}+\rho\cdot\sum_{s=0}^{t-1}[f(x_{s})-\min f]-\sum_{s=0}^{t-1}e_{s}\cdot\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}\quad\text{for $t=0,\ldots,T-1$},$
and $x_{*}\in\operatorname{argmin}f$. We replaced $\sigma_{*}^{2}$ by its equivalent value $\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$ for simplicity, to avoid the additional variable $\sigma_{*}^{2}$.
We call parameters $\rho,(a_{t}),(e_{t})$ Lyapunov parameters if $\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}\leq 0$ holds for all $t=0,\ldots,T-1$, for all $d\geq 1$, for all functions $f_{1},\dots,f_{m}\in\mathcal{F}_{\mu,L}(\mathbb{R}^{d})$, all $x_{0}\in\mathbb{R}^{d}$, and all $x_{*}\in\operatorname{argmin}f$. Note that being Lyapunov parameters depends on the number of functions $m$ and the time horizon $T$. We denote the set of all Lyapunov parameters by $\mathcal{V}_{T}$, where we omit the dependence on $m$, as our results happen to be independent of $m$.
To unify the presentation for the convex and the strongly convex setting, we introduce some notation.
  • $\bullet$
    In the convex setting, Lemma 7 tells us that we are interested in maximizing $\rho$, or equivalently minimizing $\tfrac{1}{\rho}$, where we fix $a_{0}=1$. In this setting we denote the bias and normalization by
    $\text{Bias}=\frac{1}{\rho}\quad\text{and}\quad\mathcal{N}_{T}=\{\rho,(a_{t},e_{t})\in\mathcal{V}_{T}\colon a_{0}=1\}.$
  • $\bullet$
    In the strongly convex setting, Lemma 8 tells us we are interested in minimizing $a_{0}$, where we fix $a_{T}=1$. In this setting we denote the bias and normalization by
    $\text{Bias}=a_{0}\quad\text{and}\quad\mathcal{N}_{T}=\{\rho,(a_{t},e_{t})\in\mathcal{V}_{T}\colon a_{T}=1\}.$
With this notation in mind, we may formally state our problem as
$\text{Bias}_{\text{opt}}=\text{Bias}_{\text{opt}}(T)\coloneq\inf_{\rho,(a_{t},e_{t})\in\mathcal{N}_{T}}\left\{\text{Bias}\right\}.$
The definition of being a Lyapunov parameter is such that it generates a decrease in energy at each step. Specifically, it holds true that
$\text{Bias}_{\text{opt}}$ $=\inf_{\mathcal{N}_{T}}\left\{\text{Bias}\colon\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}\leq 0~\forall t=0,\ldots,T-1,\forall x_{0}\in\mathbb{R}^{d},\forall f_{i}\in\mathcal{F}_{\mu,L},\forall d\in\mathbb{Z}_{\geq 1}\right\}.$
By permuting the quantifiers, this may then be reformulated as a bilevel program, namely:
$\text{Bias}_{\text{opt}}$ $=\inf_{\mathcal{N}_{T}}\left\{\text{Bias}\colon B_{t}\leq 0\quad\forall t=0,\ldots,T-1\right\},$
where
$B_{t}=B_{t}(\rho,(a_{t},e_{t}))\coloneq\quad\quad\sup_{d\in\mathbb{Z}_{\geq 1}}\quad$ $\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}$
[…]
$\nabla f(x_{*})=0,\quad x_{0},x_{*}\in\mathbb{R}^{d},\quad f_{i}\in\mathcal{F}_{\mu,L}.$
Note that the initial point $x_{0}$ is a decision variable of the inner (follower) problem. As a result, the constraint that $x_{t}$ must be generated by running SGD for $t$ steps from $x_{0}$ is not restrictive. Indeed, for any target point $x\in\mathbb{R}^{d}$, one can construct an initial point $x_{0}\in\mathbb{R}^{d}$ and functions $f_{1},\dots,f_{m}\in\mathcal{F}_{\mu,L}$ such that the SGD iterates satisfy $x_{t}=x$. Consequently, allowing $x_{0}$ to vary freely effectively makes the SGD-generation constraint slack. Using this, and recalling the definition of $E_{t}$, we may write that
$B_{t}=\quad\quad\sup_{d\in\mathbb{Z}_{\geq 1}}\quad$ $a_{t+1}\mathbb{E}\mathopen{}\left[{\|x_{t+1}-x_{*}\|^{2}}\right]\mathclose{}-a_{t}\|x_{t}-x_{*}\|^{2}+\rho(f(x_{t})-f(x_{*}))-e_{t}\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}$
subject to $\sum_{i=1}^{m}\nabla f_{i}(x_{*})=0,\quad x_{t},x_{*}\in\mathbb{R}^{d},\quad f_{i}\in\mathcal{F}_{\mu,L}.$
By introducing the notation $f_{j}^{(i)}=f_{i}(x_{j})$ and $g_{j}^{(i)}=\nabla f_{i}(x_{j})$ for $j\in\{t,*\}$ and $i=1,\ldots,m$, and by rewriting the expectations as finite sums, we may rewrite the problem equivalently as
$B_{t}=\quad\quad\sup_{d\in\mathbb{Z}_{\geq 1}}\quad$ $\frac{a_{t+1}}{m}\sum_{i=1}^{m}{\|x_{t}-\gamma g_{t}^{(i)}-x_{*}\|^{2}}-a_{t}\|x_{t}-x_{*}\|^{2}+\frac{\rho}{m}\sum_{i=1}^{m}(f_{t}^{(i)}-f_{*}^{(i)})-\frac{e_{t}}{m}\sum_{i=1}^{m}\|g_{*}^{(i)}\|^{2}$
subject to $\sum_{i=1}^{m}g_{*}^{(i)}=0,$
there exist $f_{i}\in\mathcal{F}_{\mu,L}$ such that $f_{i}(x_{j})=f_{j}^{(i)}$ and $\nabla f_{i}(x_{j})=g_{j}^{(i)}$
$\qquad\qquad\qquad\qquad\qquad\qquad\text{for $j\in\{t,*\}$ and $i=1,\ldots,m$},$
$x_{t},x_{*}\in\mathbb{R}^{d},\quad g_{t}^{(i)},g_{*}^{(i)}\in\mathbb{R}^{d},\quad f_{t}^{(i)},f_{*}^{(i)}\in\mathbb{R}\quad\text{for all }i=1,\ldots,m.$
The existence condition may seem like a complicated reformulation, but it allows us to use (taylor_smooth_2017, , Theorem 4), a cornerstone of the performance estimation framework:
Theorem B.1
Let $L\geq\mu\geq 0$ and let $\left\{(x_{i},g_{i},f_{i})_{i\in\mathcal{I}}\right\}\subset\mathbb{R}^{d}\times\mathbb{R}^{d}\times\mathbb{R}$ be a finite set of triplets. There exists a function $f\in\mathcal{F}_{\mu,L}$ such that
$f(x_{i})=f_{i}\quad\text{and}\quad\nabla f(x_{i})=g_{i}\quad\text{for all $i\in\mathcal{I}$},$
if, and only if, for every pair of indices $(i,j)\in\mathcal{I}^{2}$, it holds that
$f_{i}-f_{j}-\langle g_{j},x_{i}-x_{j}\rangle\geq\frac{1}{2(1-\frac{\mu}{L})}\left(\frac{1}{L}\|g_{i}-g_{j}\|^{2}+\mu\|x_{i}-x_{j}\|^{2}-2\frac{\mu}{L}\langle g_{i}-g_{j},x_{i}-x_{j}\rangle\right).$ (9)
Using this theorem, we may rewrite our program equivalently as
$B_{t}=\quad\quad\sup_{d\in\mathbb{Z}_{\geq 1}}\quad$ $\frac{a_{t+1}}{m}\sum_{i=1}^{m}{\|x_{t}-\gamma g_{t}^{(i)}-x_{*}\|^{2}}-a_{t}\|x_{t}-x_{*}\|^{2}+\frac{\rho}{m}\sum_{i=1}^{m}(f_{t}^{(i)}-f_{*}^{(i)})-\frac{e_{t}}{m}\sum_{i=1}^{m}\|g_{*}^{(i)}\|^{2}$
subject to $\left\|\sum_{i=1}^{m}g_{*}^{(i)}\right\|^{2}=0,$
[…]
$x_{t},x_{*}\in\mathbb{R}^{d},\quad g_{t}^{(i)},g_{*}^{(i)}\in\mathbb{R}^{d},\quad f_{t}^{(i)},f_{*}^{(i)}\in\mathbb{R}\quad\text{for all }i=1,\ldots,m.$
The variable $g_{*}^{(m)}$ is redundant, as it is given by $g_{*}^{(m)}=-\sum_{i=1}^{m-1}g_{*}^{(i)}$. To simplify notation, we introduce
$P_{t}$ $=\left(g_{t}^{(1)},\ldots,g_{t}^{(m)},g_{*}^{(1)},\ldots,g_{*}^{(m-1)},x_{t}-x_{*}\right)\in\mathbb{R}^{d\times 2m},$ (10a)
$F_{t}$ $=\left(f_{t}^{(1)},\ldots,f_{t}^{(m)},f_{*}^{(1)},\ldots,f_{*}^{(m)}\right)\in\mathbb{R}^{2m},$ (10b)
and define the vectors $\bm{p}_{i}\in\mathbb{R}^{2m}$ for $i=1,\ldots,2m+1$, such that
$\begin{cases}P_{t}\bm{p}_{i}=g_{t}^{(i)}&\text{for $i=1,\ldots,m$},\\ P_{t}\bm{p}_{m+i}=g_{*}^{(i)}&\text{for $i=1,\ldots,m-1$},\\ P_{t}\bm{p}_{2m}=-\sum_{i=1}^{m-1}g_{*}^{(i)},&\\ P_{t}\bm{p}_{2m+1}=x_{t}-x_{*},\end{cases}$
and $\bm{f}_{i}\in\mathbb{R}^{2m}$ for $i=1,\ldots,2m$ such that, for $i=1,\ldots,m$,
$F_{t}\bm{f}_{i}=f_{t}^{(i)},\quad F_{t}\bm{f}_{m+i}=f_{*}^{(i)}.$
With this, we may write our program as
$B_{t}=\quad\quad\quad\sup_{d\in\mathbb{Z}_{\geq 1}}$ $P_{t}^{T}\Delta_{t}P_{t}+F_{t}\tilde{\Delta}$
subject to $F_{t}\bm{f}_{m+i}-F_{t}\bm{f}_{i}+P_{t}^{T}A_{t,*}^{(i)}P_{t}\leq 0\text{ for all $i=1,\ldots,m$},$
$F_{t}\bm{f}_{i}-F_{t}\bm{f}_{m+i}+P_{t}^{T}A_{*,t}^{(i)}P_{t}\leq 0\text{ for all $i=1,\ldots,m$},$
$P_{t}\in\mathbb{R}^{d\times 2m},\quad F_{t}\in\mathbb{R}^{2m},$
where
[…]
This formulation is a non-convex quadratic program. A standard technique to render this solvable is to cast it into a semi-definite program. We do so by introducing $G_{t}=P_{t}^{T}P_{t}\succeq 0$, and obtain
$B_{t}=\sup_{d\in\mathbb{Z}_{\geq 1}}\quad\quad\quad\sup_{G_{t}\in\mathcal{S}^{2m},F_{t}\in\mathbb{R}^{2m}}\quad$ $\operatorname{Tr}(\Delta_{t}G_{t})+F_{t}\tilde{\Delta}$
subject to $F_{t}\bm{f}_{m+i}-F_{t}\bm{f}_{i}+\operatorname{Tr}(A_{t,*}^{(i)}G_{t})\leq 0\text{ for all $i=1,\ldots,m$}$
$F_{t}\bm{f}_{i}-F_{t}\bm{f}_{m+i}+\operatorname{Tr}(A_{*,t}^{(i)}G_{t})\leq 0\text{ for all $i=1,\ldots,m$}$
$\text{rank}(G_{t})\leq d,$
where $\mathcal{S}^{k}$ denotes the set of symmetric positive definite matrices of dimension $k\times k$. This formulation is equivalent to the previous one due to the rank condition: given any matrix $G_{t}\in\mathcal{S}^{k}$ of rank lesser than $d$, a Cholesky factorization $G_{t}=P_{t}^{T}P_{t}$ would yield a $P_{t}\in\mathbb{R}^{d\times 2m}$, which is feasible to the previous formulation. Now, observe that the non-convex non-continuous rank constraint may be removed due to the exterior supremum over $d\geq 1$. Once this is done, we may as well remove the supremum over $d$ because it does not appear anymore in the inner supremum. This allows us to write
$B_{t}=\quad\quad\quad\sup_{G_{t}\in\mathcal{S}^{2m},F_{t}\in\mathbb{R}^{2m}}\quad$ $\operatorname{Tr}(\Delta_{t}G_{t})+F_{t}\tilde{\Delta}$ (Primal)
subject to $F_{t}\bm{f}_{m+i}-F_{t}\bm{f}_{i}+\operatorname{Tr}(A_{t,*}^{(i)}G_{t})\leq 0\text{ for all $i=1,\ldots,m$},$
$F_{t}\bm{f}_{i}-F_{t}\bm{f}_{m+i}+\operatorname{Tr}(A_{*,t}^{(i)}G_{t})\leq 0\text{ for all $i=1,\ldots,m$}.$
We have now expressed $B_{t}$ as the optimal value of a semi-definite program. It is a simple exercise to compute the dual of this problem, which happens to be a feasibility problem:
$\tilde{B}_{t}=\quad\quad\quad\sup_{\Lambda_{t},\lambda_{t,*}^{(i)},\lambda_{*,t}}\quad$ $0$ (Dual)
subject to $-\Delta_{t}+\sum_{i=1}^{m}\left(\lambda_{t,*}^{(i)}A_{t,*}^{(i)}+\lambda_{*,t}^{(i)}A_{*,t}^{(i)}\right)=\Lambda_{t},$
$-\tilde{\Delta}+\sum_{i=1}^{m}\lambda_{t,*}^{(i)}(\bm{f}_{m+i}-\bm{f}_{i})+\sum_{i=1}^{m}\lambda_{*,t}^{(i)}(\bm{f}_{i}-\bm{f}_{m+i})=0,$
$\Lambda_{t}\in\mathcal{S}^{2m},\quad\lambda_{t,*}^{(i)},\lambda_{*,t}^{(i)}\in\mathbb{R}_{\geq 0}\quad\text{for all $i=1,\ldots,m$}.$
We prove in Section B.3 that this is indeed the dual problem of Problem (Primal). We further prove that strong duality holds, meaning that $\tilde{B}_{t}=-B_{t}$. As a consequence,
[…]
where the last equivalence follows from the structure of Problem (Dual). In conclusion, we can rewrite the problem of minimizing Bias over the Lyapunov parameters into
$\text{Bias}_{\text{opt}}=\quad\quad\inf_{\mathcal{N}_{T}}\quad$ Bias
subject to $-\Delta_{t}+\sum_{i=1}^{m}\left(\lambda_{t,*}^{(i)}A_{t,*}^{(i)}+\lambda_{*,t}^{(i)}A_{*,t}^{(i)}\right)=\Lambda_{t},$
$-\tilde{\Delta}+\sum_{i=1}^{m}\lambda_{t,*}^{(i)}(\bm{f}_{m+i}-\bm{f}_{i})+\sum_{i=1}^{m}\lambda_{*,t}^{(i)}(\bm{f}_{i}-\bm{f}_{m+i})=0,$
$\Lambda_{t}\in\mathcal{S}^{2m},\quad\lambda_{t,*}^{(i)},\lambda_{*,t}^{(i)}\in\mathbb{R}_{\geq 0}\quad\text{for all $i=1,\ldots,m$ and $t=0,\ldots,T-1$}.$
This formulation is a finite-dimensional semi-definite program that can be solved numerically.
B.2 Obtaining Mathematical Proofs
The Lagrangian attached to Problem (Primal) is given by
$L(G_{t},F_{t},\Lambda_{t},\lambda_{t,*}^{(i)},\lambda_{*,t}^{(i)})=$ $-\operatorname{Tr}(\Delta_{t}G_{t})-F_{t}\tilde{\Delta}-\operatorname{Tr}(\Lambda_{t}G_{t})$
$+\sum_{i=1}^{m}\lambda_{t,*}^{(i)}\left(F_{t}\bm{f}_{m+i}-F_{t}\bm{f}_{i}+\operatorname{Tr}(A_{t,*}^{(i)}G_{t})\right)$
$+\sum_{i=1}^{m}\lambda_{*,t}^{(i)}\left(F_{t}\bm{f}_{i}-F_{t}\bm{f}_{m+i}+\operatorname{Tr}(A_{*,t}^{(i)}G_{t})\right).$
If we have a primal-dual optimal solution $(\bar{G}_{t},\bar{F}_{t},\bar{\Lambda}_{t},(\bar{\lambda}_{t,*}^{(i)}),(\bar{\lambda}_{*,t}^{(i)}))$, it holds that, for all $G_{t}\in\mathcal{S}^{2m}$ and $F_{t}\in\mathbb{R}^{2m}$,
$L(\bar{G}_{t},\bar{F}_{t},\bar{\Lambda}_{t},(\bar{\lambda}_{t,*}^{(i)}),(\bar{\lambda}_{*,t}^{(i)}))\leq L(G_{t},F_{t},\bar{\Lambda}_{t},(\bar{\lambda}_{t,*}^{(i)}),(\bar{\lambda}_{*,t}^{(i)})).$
Due to complementary slackness, it holds that,
$L(\bar{G}_{t},\bar{F}_{t},\bar{\Lambda}_{t},(\bar{\lambda}_{t,*}^{(i)}),(\bar{\lambda}_{*,t}^{(i)}))=-\operatorname{Tr}(\Delta_{t}\bar{G}_{t})-\bar{F}_{t}\tilde{\Delta}=B_{t}.$
By strong duality (see Section B.3), we know that the existence of a primal-dual solution implies that $B_{t}=0$. As such, it holds that, for all $G_{t}\in\mathcal{S}^{2m}$ and $F_{t}\in\mathbb{R}^{2m}$,
$0\leq L(G_{t},F_{t},\bar{\Lambda}_{t},(\bar{\lambda}_{t,*}^{(i)}),(\bar{\lambda}_{*,t}^{(i)})),$
or in other terms that
$\operatorname{Tr}(\Delta_{t}G_{t})+F_{t}\tilde{\Delta}+\operatorname{Tr}(\bar{\Lambda}_{t}G_{t})$ $\leq\sum_{i=1}^{m}\bar{\lambda}_{t,*}^{(i)}\left(F_{t}\bm{f}_{m+i}-F_{t}\bm{f}_{i}+\operatorname{Tr}(A_{t,*}^{(i)}G_{t})\right)+\sum_{i=1}^{m}\bar{\lambda}_{*,t}^{(i)}\left(F_{t}\bm{f}_{i}-F_{t}\bm{f}_{m+i}+\operatorname{Tr}(A_{*,t}^{(i)}G_{t})\right).$ (11)
As $(\bar{G}_{t},\bar{F}_{t},\bar{\Lambda}_{t},(\bar{\lambda}_{t,*}^{(i)}),(\bar{\lambda}_{*,t}^{(i)}))$ is primal-dual optimal, in particular it is primal-dual feasible, and hence $\bar{\Lambda}_{t}\succeq 0$ and $\bar{\lambda}_{t,*}^{(i)},\bar{\lambda}_{*,t}^{(i)}\geq 0$ for all $i=1,\ldots,m$. In particular, if $G_{t}$ and $F_{t}$ are generated by functions $f_{i}\in\mathcal{F}_{\mu,L}$ through Equations (10), it holds that the right-hand side of Equation (11) is nonpositive by Theorem B.1. By recalling that $\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}=\operatorname{Tr}(\Delta_{t}G_{t})+F_{t}\tilde{\Delta}$, we write
$\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}\leq-\operatorname{Tr}(\bar{\Lambda}_{t}G_{t})\leq 0,$
where the latter follows as $\bar{\Lambda}_{t},G_{t}\succeq 0$. Specifically, one can prove that $\mathbb{E}\mathopen{}\left[{E_{t+1}-E_{t}}\right]\mathclose{}\leq 0$ by rewriting it in the form of (11), using the obtained dual optimal solution, and writing $\operatorname{Tr}(\bar{\Lambda}_{t}G_{t})$ as an appropriate sum of squares based on the Cholesky factorization of $\bar{\Lambda}_{t}$.
B.3 Strong Duality
Theorem B.2
Problem (Dual) is the dual of Problem (Primal). Moreover, strong duality holds.
Proof
To see that Problem (Dual) is in fact the dual of Problem (Primal), we consider the Lagrangian, which is given by
$L(G_{t},F_{t},\Lambda_{t},\lambda_{t,*}^{(i)},\lambda_{*,t}^{(i)})=$ $\operatorname{Tr}\left(\left(-\Delta_{t}+\sum_{i=1}^{m}\lambda_{t,*}^{i}A_{t,*}^{(i)}+\sum_{i=1}^{m}\lambda_{*,t}^{i}A_{*,t}^{(i)}-\Lambda_{t}\right)G_{t}\right)$
$+\left(-\tilde{\Delta}+\sum_{i=1}^{m}\lambda_{*,t}^{i}(\bm{f}_{m+i}-\bm{f}_{i})+\sum_{i=1}^{m}\lambda_{*,t}^{i}(\bm{f}_{i}-\bm{f}_{m+i})\right)F_{t},$
from which we readily recover Problem (Dual) as the dual. In order to show that strong duality holds, we employ a standard Slater argument (boyd_convex_2023, ) and construct a feasible point $(G_{t},F_{t})$ to Problem (Primal) such that $G_{t}\succ 0$.
To do so, we follow the same ideas as in (taylor_smooth_2017, , Theorem 6). Specifically, leveraging the discussion in Section B.1, we shall construct $m$ functions $f_{i}\in\mathcal{F}_{\mu,L}$ and points $x_{t},x_{*}\in\mathbb{R}^{d}$ such that the matrix $P_{t}$ given by Equation (10a) is upper triangular with positive entries on its diagonal. As such, assuming $d\geq 2m$, it will hold that $G_{t}=P_{t}^{T}P_{t}\succ 0$. We assume without loss of generality that $x_{*}=0$, and set $d=2m$.
We shall start by proving the result for $\mu=1$ and $L=6$. This is without loss of generality, as will be shown later.
We introduce the following functions $f_{i}$ and their gradients $\nabla f_{i}$ for $i=1,\ldots,m$,
$f_{i}\colon\mathbb{R}^{d}\to\mathbb{R},x\mapsto\frac{1}{2}x^{T}Q_{i}x+b_{i}x\quad\text{and}\quad\nabla f_{i}\colon\mathbb{R}^{d}\to\mathbb{R}^{d},x\mapsto Q_{i}x+b_{i}.$
We first define $f_{i}$ for $i=1,\ldots,m-1$. We let $Q_{i}$ be a block diagonal matrix of the form
$Q_{i}=\begin{pmatrix}Q_{i}^{1}&0\\ 0&Q\end{pmatrix}-\bm{e}_{m+i}\bm{e}_{m+i}^{T},$
where $Q_{i}^{1}\in\mathbb{R}^{(m-1)\times(m-1)}$ is a diagonal matrix with entries $2$ everywhere expect an entry of $5$ on the $i$th component, and $Q\in\mathbb{R}^{(m+1)\times(m+1)}$ is a tridiagonal matrix whose diagonal elements are $4$ and whose off diagonal elements are 1-1. We note that , $\sigma(Q_{i})\subset(1,6)$ for all $i=1,\ldots,m-1$, by the Greshgorian Theorem (quarteroni_numerical_2007, ).
We define $x_{t}=x=[y,z]$ with $y\in\mathbb{R}^{m-1}$ and $z\in\mathbb{R}^{m+1}$. Then, for $i=1,\ldots,m-1$,
$\nabla f_{i}(x_{t})=g_{t}^{(i)}=2\sum_{j=1,j\neq i}^{m-1}\bm{e}_{j}x_{j}+5\bm{e}_{i}x_{i}+Qz-\bm{e}_{m+i}x_{m+i}+b_{i}.$
As such, we define $b_{i}$ as
$b_{i}=-2\sum_{j=1,j\neq i}^{m-1}\bm{e}_{j}x_{j}+\bm{e}_{m+i}x_{m+i}-\bm{e}_{m},$
and we impose that $x>0$ and that $Qz=[1,0,\ldots,0]\in\mathbb{R}^{m+1}$ (may be shown possible by a simple induction argument). This thus gives us that $g_{t}^{(i)}=5\bm{e}_{i}x_{i}$. Moreover, as $x_{*}=0$, we have that $g_{*}^{(i)}=b_{i}$. Note that $g_{t}^{(i)}$ has a positive entry in its $i$th component and 0s everywhere else. Moreover, $g_{*}^{(i)}$ has a positive component in its $m+i$th component and 0 in all subsequent components. The vector $x_{t}-x_{*}$ also has a positive entry in component $2m$. In order to show that $P_{t}$ is upper triangular with positive diagonal entries, we thus only need to construct $g_{t}^{(m)}$.
By construction we require $\nabla f(x_{*})=0$, namely that
$b_{m}=-\sum_{i=1}^{m-1}b_{i}=\sum_{i=1}^{m-1}(2m-1)\bm{e}_{i}x_{i}-\sum_{i=1}^{m-1}\bm{e}_{m+i}x_{m+i}+(m-1)\bm{e}_{m}.$
As such, we may define $Q_{m}$ to be the identity matrix (which satisfies $\sigma(Q_{m})=\{1\}$), such that
$\nabla f_{m}(x_{t})=g_{t}^{(m)}=\sum_{i=1}^{m-1}2m\bm{e}_{i}x_{i}+m\bm{e}_{m}.$
As such, $g_{t}^{(m)}$ has a positive entry in its $m$th component and zero entries in all subsequent components. As such, $P_{t}$ is upper triangular with positive diagonal entries, as wanted.
As such, if $(\mu,L)=(1,6)$, the result holds. For arbitrary $(\mu,L)$, we consider the following transformed functions
$\tilde{f}_{i}(x)=\frac{L-\mu}{5}\left(f_{i}(x)-\frac{1}{2}\|x\|^{2}\right)+\frac{\mu}{2}\|x\|^{2}.$
As $f_{i}\in\mathcal{F}_{1,6}$, we have $\tilde{f}_{i}\in\mathcal{F}_{\mu,L}$. Moreover, this transformation preserves the property on $P_{t}$, hence constructing a Slater point for arbitrary $(\mu,L)$.
Appendix C SGD with Non-Uniform Sampling and Mini-Batching
All the main results in our paper can be applied to variants of (SGD), such as mini-batching or nonuniform sampling. This is because these variants can be seen as instances of (SGD) applied to different but equivalent problems. Our results can be applied as a black-box, and the results follow by computing the main constants of the equivalent problem, namely $L,\mu,\sigma_{*}^{2}$. This strategy was described in gower_sgd_2019 . For the sake of completeness, we include some details for the two aforementioned instances below.
C.1 Non-Uniform Sampling
Consider a finite family of functions $f_{1},\dots,f_{n}$ which we assume to be $\mu_{i}$-strongly convex and $L_{i}$-smooth, for $\mu_{i}\geq 0$ and $L_{i}>0$. We want to study the non-uniform SGD algorithm, which computes
$x_{t+1}=x_{t}-\frac{\gamma}{np_{i_{t}}}\nabla f_{i_{t}}(x_{t}),$ (SGDp)
where $i_{t}\in\{1,\dots,n\}$ is sampled according to $\mathbb{P}(i_{t}=i)=p_{i}$, with $p_{i}>0$ and $\sum_{i=1}^{n}p_{i}=1$. Note that we are using here a specific renormalization of the step-size, following the ideas from gower_sgd_2019 . It is immediate to see that this algorithm is exactly (SGD) applied to the problem
$\min_{x}\ f(x)=\mathbb{E}_{{\mathcal{P}}}\mathopen{}\left[{\hat{f}_{i}(x)}\right]\mathclose{},\text{ where }\hat{f}_{i}(x)\coloneqq\frac{1}{np_{i}}f_{i}(x),$
where the expectation is taken with respect $\mathcal{P}$, the distribution whose density is given by the probabilities $p_{i}$. Note that the function $f$ remains unchanged, meaning that the set of minimizers is the same.
We note that Assumption 1 holds for $\mu=\min_{i}\tfrac{\mu_{i}}{np_{i}}$ and $L=\max_{i}\tfrac{L_{i}}{np_{i}}$. Moreover, one can compute that
$\hat{\sigma}_{*}^{2}=\mathbb{E}\mathopen{}\left[{\|\nabla\hat{f}_{i}(x_{*})\|^{2}}\right]\mathclose{}=\mathbb{E}\mathopen{}\left[{\frac{1}{n^{2}p_{i}^{2}}\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}\leq\frac{\mathbb{E}\mathopen{}\left[{\|\nabla f_{i}(x_{*})\|^{2}}\right]\mathclose{}}{n^{2}\min_{i}p_{i}^{2}}=\frac{\sigma_{*}^{2}}{n^{2}\min_{i}p_{i}^{2}},$
which is finite if Assumption 2 holds for the original problem. If we now impose that $\gamma L\in(0,2)$, we can apply all our results and obtain bounds on the iterates of (SGDp) depending on the above constants.
C.2 Mini-Batching
Consider a finite family of functions $f_{1},\dots,f_{n}$ which we assume to be $\mu_{i}$-strongly convex and $L_{i}$-smooth, for $\mu_{i}\geq 0$ and $L_{i}>0$. We want to study the mini-batch SGD algorithm, which computes at each iteration a mini-batch of fixed size $1\leq b\leq n$, that is
$x_{t+1}=x_{t}-\frac{\gamma}{b}\sum_{i\in B_{t}}\nabla f_{i}(x_{t}),$ (SGDb)
where $B_{t}$ is sampled i.i.d. and uniformly among the subsets of size $b$ of $\{1,\dots,n\}$. It is a simple exercise to see that this algorithm is exactly (SGD) applied to the problem
$\min_{x}\ f(x)=\mathbb{E}_{{\mathcal{B}}}\mathopen{}\left[{\hat{f}_{B}(x)}\right]\mathclose{},\text{ where }\hat{f}_{B}(x)\coloneqq\frac{1}{b}\sum_{i\in B}\nabla f_{i}(x),$
and where the expectation is taken with respect to $\mathcal{B}$ which is the uniform law over
$\text{batch}_{b}\coloneqq\{B\subset\{1,\dots,n\}\ :\ |B|=b\}.$
We note that Assumption 1 holds for $\mu=\min_{i}\mu_{i}$ and $L=\max_{i}L_{i}$. Moreover, one can compute that
$\sigma_{*}^{2}$ $=\mathbb{E}_{{\mathcal{B}}}\mathopen{}\left[{\|\nabla f_{B}(x_{*})\|^{2}}\right]\mathclose{}=\frac{1}{|\mathcal{B}|}\sum_{B\in\mathcal{B}}\left\|\frac{1}{b}\sum_{i\in B}\nabla f_{i}(x_{*})\right\|^{2}=\frac{1}{\binom{n}{b}\cdot b^{2}}\sum_{B\in\mathcal{B}}\sum_{i,j\in B}\langle\nabla f_{i}(x_{*}),\nabla f_{j}(x_{*})\rangle$
$=\frac{1}{\binom{n}{b}\cdot b^{2}}\left[\binom{n-2}{b-2}\sum_{i,j=1}^{n}\langle\nabla f_{i}(x_{*}),\nabla f_{j}(x_{*})\rangle+\left[\binom{n-1}{b-1}-\binom{n-2}{b-2}\right]\sum_{i=1}^{n}\|\nabla f_{i}(x_{*})\|^{2}\right]=\frac{n-b}{nb(n-1)}\sum_{i=1}^{n}\|\nabla f_{i}(x_{*})\|^{2},$
which is finite if Assumption 2 holds for the original problem.